### v1

# Tighter Lower Bounds for Single Source Personalized PageRank

Abstract
We investigate the lower bounds for approximating the well-known Single Source Personalized PageRank (SSPPR) query, which quantifies the probabilities of an $\alpha$-decay random walk originating from node $s$ and ending at all other nodes. Considering two standard error settings—absolute error (SSPPR-A) and relative error (SSPPR-R)—our goal is, to with constant probability, return an estimate $\widehat{\pi}_{s}(t)$ satisfying $|\widehat{\pi}_{s}(t)-\pi_{s}(t)|\leq\varepsilon$ for all nodes $t$ with accuracy parameter $\varepsilon>0$, and $|\widehat{\pi}_{s}(t)-\pi_{s}(t)|\leq c\cdot\pi_{s}(t)$ for nodes $t$ with $\pi_{s}(t)\geq\delta$, where $c\in(0,1)$ is a constant and $\delta>0$ is the threshold parameter (e.g., $\delta=1/n$, reflecting the average PPR score over the graph).
The pursuit of tighter upper bounds for SSPPR query computation has been extensively explored, with the best-known complexities given by $\operatorname{\mathcal{O}}\left(\min\left(\frac{\log(1/\delta)}{\delta},\sqrt{\frac{m\log(n)}{\delta}},m\log\left(\frac{\log(n)}{\delta m}\right)\right)\right)$ for the SSPPR-R setting, and $\operatorname{\mathcal{O}}\left(\min\left(\frac{\log(1/\varepsilon)}{\varepsilon^{2}},\frac{\sqrt{m\log(n)}}{\varepsilon},m\log\frac{1}{\varepsilon}\right)\right)$ for the SSPPR-A setting. However, the corresponding lower bounds remain significantly looser—only $\Omega\left(\min(m,1/\delta)\right)$ and $\Omega\left(\min(n,1/\varepsilon)\right)$, respectively. This results in substantial gaps between the upper and lower bounds. For instance, in the SSPPR-R case, there is still a logarithmic gap ranging from $\log(1/\delta)$ up to $\log m$, which highlights a crucial margin yet to be closed in the theoretical understanding of the problem’s complexity.
To bridge this gap, we achieve tighter lower bounds for this important problem as follows:
  • For SSPPR-R, we prove an improved lower bound of $\Omega\left(\min(m,\frac{\log(\frac{1}{\delta})}{\delta})\right)$ for any $\delta\in(0,1)$.
  • For SSPPR-A, we establish an overall lower bound of $\Omega\left(\min(m,\frac{\log(\frac{1}{\varepsilon})}{\varepsilon})\right)$ for any $\varepsilon\in(0,1)$ and $m\in\operatorname{\mathcal{O}}(n^{2-\beta})$ for any arbitrary small constant $\beta\in(0,1)$.
In summary, we improve the previous best-known lower bound results by a logarithmic factor $\log(\frac{1}{\delta})$ (SSPPR-R) or $\log(\frac{1}{\varepsilon})$ (SSPPR-A), where we achieve matching lower and upper bounds on SSPPR-R query when $\delta$ is around $1/n.$
[…]
1 Introduction
The Personalized PageRank (PPR) algorithm, first proposed by Google [34], has gradually become a fundamental graph query, attracting widespread interest from both academia and industry. Given a graph $G$, the PPR value $\pi_{s}(t)$ with respect to a source node $s$ and a target node $t$ is defined as the probability of an $\alpha$-decay random walk***with probability $\alpha$ to stop at the current node and probability $1-\alpha$ to randomly transition to neighbors starting from $s$ and terminating at $t$. This definition naturally captures the importance of node $t$ relative to $s$, making the Single Source Personalized PageRank (SSPPR) query—which computes the PPR values of all other nodes relative to $s$—widely applicable in various domains including web search [6; 42], recommendation systems [9; 49; 32], graph sparsification [36; 51], graph neural networks [17; 4; 40], and others. Exact computation of SSPPR is computationally expensive, requiring matrix inversion with $\Omega(n^{2}\log n)$ time complexity for a graph with $n$ nodes [35], which is prohibitive for large real-world graphs. Consequently, significant efforts have been dedicated to developing efficient approximation algorithms for SSPPR, aiming to achieve a favorable trade-off between efficiency and accuracy. In recent decades, a series of practically effective algorithms with theoretical guarantees have emerged [11; 48; 38; 30; 50; 3; 24; 23; 26]. Building upon classical methods such as the Monte Carlo (MC) and Forward Push (FP) algorithms, recent state-of-the-art approaches combine these techniques to improve time complexity. However, despite continual practical improvements, fundamental breakthroughs on the computational upper bounds appear to have stalled since the introduction of two representative methods, FORA [48] and SpeedPPR [39], both of which remain based on MC and FP combinations. This raises a natural question: to what extent can current upper bounds be further improved?
However, non-trivial and broadly applicable lower bounds for this important query remain scarce, as constructing hard instances that compel any algorithm to perform substantial computation for SSPPR is challenging. The best known lower bounds are $\Omega\big{(}\min(n,1/\varepsilon)\big{)}$ for SSPPR with absolute error, established by [47], and $\Omega\big{(}\min(m,1/\delta)\big{)}$ for SSPPR with relative error, recently presented by [5]. The latter, a concurrent work, primarily addresses single pair lower bounds, which naturally extend to single source queries but do not fully exploit the single source property involving multiple target nodes—a gap we further address. Specifically, the former bound arises from considering the worst-case output size of the PPR vector, while the latter is derived by extending results from the simpler Single Pair PPR query, which examines only one node pair and thus represents a weaker form of the SSPPR query. In this paper, we make significant progress by establishing tighter lower bounds improved by a substantial logarithmic factor. Our results considerably narrow the gap between previously known upper and lower bounds, underscoring the optimality of the Monte Carlo method under relevant parameter regimes.
The remainder of this section formally states the problem, reviews existing bounds, and outlines our motivations and contributions
1.1 Problem Formulation
Among the comprehensive approximation measurements including the $\ell_{1}$, $\ell_{2}$, absolute, relative and degree-normalized absolute error, we focus on the research problem of SSPPR query with relative and additive error guarantee unifiedly. These two settings are widely utilized and studied in mainstreaming PPR-related researches [48; 38; 11; 45] which requires the estimator to be accurate within chosen error requirement. We give their formal definition below:
Definition 1.
Single Source Personalized PageRank Query with Relative Error (SSPPR-R) guarantee. Given a graph $G$, a decay factor $\alpha$, a source node $s$, a relative error parameter $c\in(0,0.5)$, an overall failure probability parameter $p_{f}$ and choosing a targeting PPR threshold $\delta\in(0,1)$, the SSPPR-R query requires an estimated PPR vector $\widehat{\pi}_{s}(\cdot)$ such that with at least $1-p_{f}$ probability,
$|\widehat{\pi}_{s}(t)-\pi_{s}(t)|\leq c\cdot\pi_{s}(t)$
holds for all node $t$ with its PPR value $\pi_{s}(t)\geq\delta$. Note that the threshold $\delta$ can be chosen arbitrary close to 0 or 1, which correspond to the shallow or large targeting PPR values.
Definition 2.
Single Source Personalized PageRank Query with Additive Error (SSPPR-A) guarantee. Given a graph $G$, a decay factor $\alpha$, a source node $s$, an additive error parameter $\varepsilon\in(0,1)$ and an overall failure probability parameter $p_{f}$ the SSPPR-A query requires an estimated PPR vector $\widehat{\pi}_{s}(\cdot)$ such that with at least $1-p_{f}$ probability,
$|\widehat{\pi}_{s}(t)-\pi_{s}(t)|\leq\varepsilon$
holds for all node $t$. Following prior work [16; 18; 46; 22], we assume the parameters $\alpha,c$ and $p_{f}$ as constants by default.
Table 1: Best-Known results of SSPPR-A and SSPPR-R query and our improvements. Note that [5] is a recent concurrent work to ours, and we remark its results with underlining. We also highlight our improvement when $\varepsilon,\delta$ are sufficiently small. For ease of simplicity, we here assume $m\in\operatorname{\mathcal{O}}(n^{2-\beta})$ for any arbitrary small constant $\beta\in(0,1)$ in our Tighter SSPPR-A lower bound.
Best-Known Lower Bounds Tighter Lower Bounds (Ours) Improvements
SSPPR-A $\Omega\left(\min(n,\frac{1}{\varepsilon})\right)$ [47] $\Omega\left(\min(m,\frac{\log(\frac{1}{\varepsilon})}{\varepsilon})\right)$ $\log(\frac{1}{\varepsilon})$ or ${\frac{m}{n}}$.
SSPPR-R $\Omega\left(\min(m,\frac{1}{\delta})\right)$ [5] $\Omega\left(\min(m,\frac{\log(\frac{1}{\delta})}{\delta})\right)$ $\log(\frac{1}{\delta}),$ up to ${\log(m)}$
Best-Known Upper Bounds
[…]
1.2 Existing Results
From a theoretical perspective, we summarize the state-of-the-art upper and lower bounds for both SSPPR-R and SSPPR-A queries. An overview of these results is presented in Table 1.
1.3 Upper Bounds
We briefly summarize the best-known computational upper bounds for answering Single-Source Personalized PageRank (SSPPR) queries under both absolute (SSPPR-A) and relative (SSPPR-R) error guarantees. These bounds are primarily derived from three foundational algorithmic paradigms—Monte Carlo (MC), Forward Push (FP), and Backward Push (BP)—and have been further refined through various state-of-the-art (SOTA) techniques. A comprehensive review of these algorithms, along with their complexity analyses, is provided in Appendix A.1.
Basic Techniques.
The MC method provides a sample-based approximation to the PPR vector using $\alpha$-decay random walks, and achieves high-probability relative error guarantees with expected time complexity $\operatorname{\mathcal{O}}\left(\log(1/\delta)/\delta\right)$ for SSPPR-R, or $\operatorname{\mathcal{O}}\left(\log(1/\varepsilon)/\varepsilon^{2}\right)$ for SSPPR-A. The FP method is a local-push algorithm that simulates random walks deterministically via graph traversal. To guarantee an additive error of $\varepsilon$, it requires setting a residue threshold $r_{\max}=\varepsilon/m$, resulting in a worst-case time complexity of $\operatorname{\mathcal{O}}(m/\varepsilon)$. The BP method is similarly structured but performs the push operation in reverse, and when applied naively to SSPPR-A for all target nodes, also yields $\operatorname{\mathcal{O}}(m/\varepsilon)$ complexity.
State-of-the-Art Methods.
To address the inefficiencies of the basic methods, hybrid algorithms combining MC with local-push approaches have been developed. FORA [48] improves upon MC by first performing deterministic FP and then refining the result via targeted MC walks. By optimally balancing both phases, it achieves an expected complexity of $\operatorname{\mathcal{O}}(\sqrt{m\log(n)/\delta})$ for SSPPR-R. SpeedPPR [38] restructures FP to enforce a more regular push schedule, reducing the sum of residues more efficiently. Its enhanced analysis enables a bound of $\operatorname{\mathcal{O}}(m\log(1/\varepsilon))$ for SSPPR-A and similar efficiency for SSPPR-R when combined with MC. More recently, ABP [47] combines MC with an adaptive variant of BP. It constructs a candidate set through rough MC estimation, then uses selective backward pushes and additional walks to refine the result. This yields a sublinear expected time complexity of $\operatorname{\mathcal{O}}(\sqrt{m\log(n)}/\varepsilon)$ for SSPPR-A.
These methods offer near-optimal performance under practical settings and showcase how combining sampling with localized deterministic updates can significantly improve scalability for large-scale graph PPR queries.
1.4 Lower Bounds
Unlike the extensive body of work focused on improving the upper bounds for SSPPR query computation, research on lower bounds remains largely heuristic, offering limited rigorous insight into the fundamental computational complexity. The first lower bound relevant to Personalized PageRank (PPR) computation was introduced in FastPPR [20], establishing a lower bound of $\Omega(\sqrt{1/\delta})$ for computing Single-Pair PPR queries with relative error (SSPPR-R). More recently, concurrent to our work, [5] improved this lower bound to $\Omega\big{(}\min(m,1/\delta)\big{)}$ by constructing more sophisticated hard instances based on edge-level analysis. To the best of our knowledge, no other non-trivial direct lower bounds exist for the SSPPR-R query. While the SSPPR-A lower bound can be naturally extended to SSPPR-R, it remains somewhat trivial—estimating PPR values above $\delta$ inherently implies that there are at most $\operatorname{\mathcal{O}}(1/\delta)$ nodes $t$ such that $\pi(s,t)\geq\delta$, leading to a straightforward lower bound of $\Omega(1/\delta)$. This leaves a notable gap between the best-known lower and upper bounds, motivating the need for tighter analysis.
Regarding the SSPPR-A query, [47] presents a similar lower bound of $\Omega(\min(n,1/\varepsilon))$, considering that algorithms must return nonzero estimates for nodes $t$ with $\pi_{s}(t)\geq\varepsilon$, of which there may be $\Omega(\min(n,1/\varepsilon))$ nodes. While this setting is conceptually simple, it underestimates the difficulty of accurately estimating all required PPR values within satisfactory error margins. To the best of our knowledge, no further studies have provided deeper insights into SSPPR-A lower bounds.
Meanwhile, significant progress has been made on a related but distinct problem: estimating nodal PageRank (PR) values. It has been shown that the simple BiPPR algorithm achieves optimal computational complexity for directed graphs [46], and a Monte Carlo variant, BACKMC [37], attains matching upper and lower bounds for undirected graphs. The proof techniques in these works are notably simple yet elegant, strongly motivating the pursuit of tighter lower bounds for SSPPR queries. Inspired with these insights, we carefully design a hard instance tailored specifically for SSPPR queries. Its structure and analytical approach differ substantially from those in [46; 37], involving more intricate complexity analysis and leading to new novel results and conclusions.
1.5 Our Results and Discussion
Our lower bounds are mainly based on the arc-centric model [12; 14], which will be formally defined in Section 3.
Theorem 3 (Lower bound of SSPPR-R query).
Choose any decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, and error bound $c\in(0,\frac{1}{2}]$, and any functions $\delta(n)\in(0,1),m(n)\in\Omega(n)\cap\operatorname{\mathcal{O}}(n^{2})$. Consider any oracle (possibly randomized) algorithm within the arc-centric model that, for arbitrary graph $G$ with $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes and any source node $s$, with probability at least $1-p$ gives an estimation $\widehat{\pi}_{s}(v)$ for PPR vector $\pi_{s}(v)$, satisfying $|\widehat{\pi}_{s}(t)-\pi_{s}(t)|\leq c\cdot\pi_{s}(t)$ for eacg node $v$ at $\pi_{s}(v)\geq\delta$. Then there exists a graph $\mathrm{U}$ and a source node $s$ such that:
  1. (i).
    $\mathrm{U}$ contains $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes;
  2. (ii).
    the algorithm requires $\Omega\left(\min(m,\frac{\log(\frac{1}{\delta})}{\delta})\right)$ queries to succeed in expectation.
Theorem 4 (Lower bound of SSPPR-A query).
Choose any decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, and error bound $c\in(0,\frac{1}{2}]$, and any functions $\varepsilon(n)\in(0,1),m(n)\in\Omega(n)\cap\operatorname{\mathcal{O}}(n^{2-\beta})$, where $\beta\in(0,1)$ is an arbitrary small constant. Consider any oracle (possibly randomized) algorithm within the arc-centric model that, for arbitrary graph $G$ with $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes and any source node $s$, with probability at least $1-p$ gives an estimation $\widehat{\pi}_{s}(v)$ for PPR vector $\pi_{s}(v)$, satisfying $|\widehat{\pi}_{s}(t)-\pi_{s}(t)|\leq\varepsilon$ for each node $v$. Then there exists a graph $\mathrm{U}$ and a sources $s$ such that:
  1. (i).
    $\mathrm{U}$ contains $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes;
  2. (ii).
    the algorithm requires $\Omega\left(\min(m,\frac{\log(\frac{1}{\varepsilon})}{\varepsilon})\right)$ queries to succeed in expectation.
Remark.
We improve the previous best lower bounds by a logarithmic factor, as summarized in Table 1, and achieve matching upper and lower bounds when $\delta$ is on the order of $\frac{1}{n}$. Note that $\varepsilon$ and $\delta$ are tunable parameters in the range $(0,1)$ and can be set arbitrarily small. These parameters are often related to the graph scale parameters $n$ and $m$. For instance, $\varepsilon$ is frequently chosen to be as small as $\frac{1}{m}$, and $\delta=\operatorname{\mathcal{O}}(\frac{1}{n})$ represents one of the most widely adopted settings in SSPPR query computations [48; 39; 22; 18]. This highlights the practical significance of our results and marks a milestone in the theoretical understanding of SSPPR algorithms.
2 Other Related Works
Over the past decades, Personalized PageRank (PPR) computation has been extensively studied, with the majority of research focusing on accelerating and scaling up single-source [11; 48; 38; 30; 50; 3; 24; 23; 26], single-target [21; 19; 43; 45; 44], and top-$k$ [10; 19; 20; 48; 41] PPR queries on general graphs. Among these, [3; 11; 26; 10] use large-scale Monte Carlo (MC) simulations to estimate PPR values, while [21; 2] apply deterministic graph traversal to compute exact PPR vectors. Subsequent works [48; 38; 30; 41; 19; 43] improve scalability and efficiency by combining deterministic traversal and MC in separate phases. More recent studies [45; 44; 13] investigate executing deterministic graph traversal and MC steps in an integrated fashion, known as Randomized Push (RS). Although RS can be used to solve the SSPPR-R query, it requires $\operatorname{\mathcal{O}}\left(\log^{3}(n)/\delta\right)$ total computation time [22], which is worse than the basic MC method and thus omitted in earlier discussions. Other contemporary efforts address PPR computation in dynamic graph settings [33; 31; 15; 29; 52], and its integration with graph neural networks [27; 28; 25]. In summary, despite the substantial body of work on PPR computation, theoretical understanding—particularly improving lower bounds for SSPPR-A and SSPPR-R queries—remains limited. This is largely due to the challenges in constructing hard instances tailored to single-source computations. Our work thus provides new theoretical insights and opens avenues for further research into lower bounds for PPR algorithms.
3 Notations and Tools
Notations.
[…]
Definition 5 (Multi-graph and edge multiplicity).
A multi-graph is a graph that may contain parallel edges (multiple edges between the same pair of nodes); a simple graph contains no such parallel edges. For a multi-graph $G=(V,E)$, the edge multiplicity is defined as
$\mathrm{MUL}(G)\;=\;\max_{u,v\in V}\bigl{|}\{\,e\in E\mid e=(u,v)\}\bigr{|},$
representing the maximum number of parallel edges between any node pair $u,v\in V$.
Definition 6 (Maximum posterior estimation).
Let $\Theta$ be a prior distribution over a finite set of possible values, and let $\mathbf{x}$ denote the observed data. Then, there exists a maximum a posteriori (MAP) estimate $\widehat{\theta}_{\mathrm{MAP}}\in\Theta$ such that the success probability of any algorithm that predicts $\theta\in\Theta$ based on $\mathbf{x}$ is at most $\operatorname{{\bf Pr}}[\widehat{\theta}_{\mathrm{MAP}}\mid\mathbf{x}]$. Furthermore, replacing $\mathbf{x}$ with any function of $\mathbf{x}$ (i.e., using only partial or derived information) cannot increase the prediction success probability.
Arc–Centric Graph Query Model [12; 14] and Query Complexity.
For our results, we focus on the computational complexity under the standard RAM model. To formalize graph interactions, we allow algorithms to access the underlying graph $G$ through a graph oracle. Specifically, we adopt the standard arc-centric graph-access model [12; 14], where the oracle supports the following queries in unit time:
  • JUMP(): returns a uniformly random node $v\in V$ and marks $v$ as covered;
  • INDEG$(v)$/OUTDEG$(v)$: for any covered node $v\in V$, returns its in-degree $d_{in}(v)$ or out-degree $d_{out}(v)$, respectively;
  • ADJout$(v,i)$/ADJin$(v,i)$: for any covered node $v\in V$ and integer $1\leq i\leq d_{out}(v)/d_{in}(v)$, returns the $i^{th}$ out-/in-neighbor of $v$.
This model has been widely used in prior work [20; 18; 46] to establish lower bound complexities. In our analysis, we strengthen the adjacency query ADJout$(v,i)$ (analogously for ADJin) to not only return the out-neighbor node $w$ of $v$ but also provide the label of the in-edge $(v,w)$ for node $w$ denoted as $k$, such that edge $(v,w)$ is the $i$-th out edge for $v$ and $k$-th in-edge for $w$. Note that this stronger model reveals more information than the original, ensuring that any lower bounds we derive remain valid for the standard arc-centric model. For simplicity and clarity in subsequent sections, we adopt this enhanced model by default. In the context of graph algorithms, query complexity refers to the number of queries the algorithm invokes to the graph oracle. Hence, query complexity serves as a lower bound to computational complexity.
Definition 7 (Graph with edge and node indices).
Following the notation in [12], we define a graph with edge and node indices as a graph $G=(V,E)$ together with the following labeling:
  • Node labels. Each node $v\in V$ is assigned a unique label from the set $\{1,\ldots,|V|\}$;
  • Edge labels. For each node $v\in V$, its in-edges and out-edges are distinctly labeled with integers from $\{1,2,\ldots,d_{\mathrm{in}}(v)\}$ and $\{1,2,\ldots,d_{\mathrm{out}}(v)\}$, respectively.
4 Proofs
High-level Ideas.
We begin by constructing a carefully designed graph instance family, denoted by $\mathcal{U}(n,D,d)$, parameterized by integers $n$, $D$, and $d$ to be specified later. Each instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ contains a distinguished source node $s$ and $2n$ target nodes, evenly partitioned into two disjoint sets: $X_{1}$ and $X_{2}$, as illustrated in Figure 1(a). By carefully engineering the edge structure, we ensure that every node $x_{1}\in X_{1}$ has a higher PPR value from $s$ than any node $x_{2}\in X_{2}$, satisfying $\pi_{s}(x_{1})-\pi_{s}(x_{2})\geq 2\varepsilon$ for SSPPR-A, or $\pi_{s}(x_{1})/\pi_{s}(x_{2})\geq 2/(1-c)$ for SSPPR-R, as shown in Lemmas 10 and 11, respectively. That is, nodes in $X_{1}$ and $X_{2}$ differ in their PPR values by at least the required additive or relative error threshold. To succeed in this setting, any algorithm $\operatorname{\mathcal{A}}$ must detect structural differences between nodes in $X_{1}$ and $X_{2}$, which are intentionally subtle and primarily reflected in degree statistics. To model this formally, we define a uniform distribution $\Sigma(n,D,d)$ over the legal graph instances, as described in Lemma 12. This distribution is constructed by considering all valid permutations of node and edge labels and all balanced splits of the $X$ node set. Under this formulation, we can quantify the success probability of any algorithm $\operatorname{\mathcal{A}}$ operating under a query budget $T$. A subtle technical challenge arises from the potential presence of parallel edges, which may occur in our multi-graph construction. Lemma 14 guarantees that the maximum multiplicity of parallel edges is bounded by a constant $L$ with high probability. Combined with Theorem 15, this ensures that any algorithm $\operatorname{\mathcal{A}}$ designed for simple graphs can be adapted to work on multi-graphs without incurring significant additional query complexity.
[…]
(a) Instance $\mathcal{U}$
(b) $r$-padded instance $\mathcal{U}(r)$
Figure 1: Graph instance family $\mathcal{U}(n,D,d)=\mathcal{U}(3,2,1)$. (a) denotes our base graph instance family $\mathcal{U}(n,D,d)$ and (b) represents the $r$-padded instance $\mathcal{U}(r)$ based on $\mathcal{U}$.
4.1 Hard Instance Family — $\mathcal{U}(n,D,d)$
We define the graph family $\mathcal{U}(n,D,d)$ such that, for every sufficiently large $n$, it serves as a collection of hard instances for estimating SSPPR queries. This family is constructed based on a simple yet insightful core structure $\mathbf{U}$, illustrated in Figure 1(a). The graph consists of five disjoint node sets: $\{s\}\cup Y_{1}\cup Y_{2}\cup X_{1}\cup X_{2}$. For brevity, we let $Y=Y_{1}\cup Y_{2}$ and $X=X_{1}\cup X_{2}$. The node $s$ serves as the designated source node, while all nodes in $X$ act as target nodes whose PPR values with respect to $s$ are of interest. We ensure that each of the sets $X_{1}$, $X_{2}$, $Y_{1}$, and $Y_{2}$ contains $\Theta(n)$ nodes. Regarding edge construction, each node $y_{1}\in Y_{1}$ receives an incoming edge from the source node $s$. For each node $x_{1}\in X_{1}$, we construct $D$ incoming edges from nodes in $Y_{1}$ (i.e., $\deg_{Y_{1}}(x_{1})=D$) and $D-d$ edges from $Y_{2}$ (i.e., $\deg_{Y_{2}}(x_{1})=D-d$). Similarly, each $x_{2}\in X_{2}$ receives $D$ edges from $Y_{2}$ and $D-d$ from $Y_{1}$. The critical difference is that nodes in $X_{1}$ are indirectly reachable from $s$ via a useful path through $Y_{1}$, whereas nodes in $X_{2}$ are only connected through useless paths via $Y_{2}$ that do not connect to $s$. This design ensures a measurable PPR gap between $X_{1}$ and $X_{2}$. Additionally, we impose that every node $y\in Y$ has out-degree $d_{\text{out}}(y)=2D-d$ to ensure uniformity and support our lower-bound construction. With appropriately chosen parameters, we show that nodes in $X_{1}$ and $X_{2}$ exhibit a clear and quantifiable difference in their PPR values with respect to $s$. This distinction is formally captured in Lemma 10 for the SSPPR-A query and in Lemma 11 for the SSPPR-R query. To facilitate analysis in the SSPPR-R setting, we next introduce a modified structure, referred to as the $r$-padded instance, which better accommodates the relative-error guarantees.
Definition 8 ($r$-padded instance).
$\forall\mathrm{U}\in\mathcal{U}(n,D,d)$, we define the $r$-padded instance of $\mathrm{U}$, denoted as $\mathrm{U}(r)=(V_{U(r)},E_{U(r)})$ in the below by involving two independent node sets $Z_{X}$ and $Z_{Y}$:
  1. 1.
    $V_{U(r)}=V_{U}\cup Z_{X}\cup Z_{Y}$, where $Z_{X},Z_{Y}$ are two additional node sets containing exactly $r$ nodes;
  2. 2.
    $E_{U(r)}=E_{U}\cup\{(z_{x},x)\;\mid\forall x\in X,z_{x}\in Z_{X}\}\cup\{(y,z_{y})\;\mid\forall y\in Y,z_{y}\in Z_{Y}\}$, which incorporates extra edges coming from node set $Z_{x}$ to $X$ and $Y$ to $Z_{Y}$.
An illustrative example is provided in Figure 1(b). The $r$-padded instance $\mathrm{U}(r)$ is specifically designed to facilitate the analysis of SSPPR-R queries. Intuitively, when $r$ is sufficiently large relative to $D$, each ADJ query originating from a node in $X$ or $Y$ may be misdirected to the auxiliary node sets $Z_{X}$ or $Z_{Y}$ with non-negligible probability. This increases the expected query cost for algorithms attempting to explore useful edges. With this construction in place, we now proceed to compute the PPR values from the source node $s$ to nodes in the sets $X_{1}$ and $X_{2}$, as formalized in the following lemma:
Lemma 9 (PPR Calculation).
For any instance $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following properties hold:
  1. (i).
    The graph $\mathrm{U}$ contains $14n+$ 1 nodes and $2n(2D-d)+n$ edges;
  2. (ii).
    For each node $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$, the Personalized PageRank values with respect to the source node $s$ are $\pi_{s}(x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d)n}$ and $\pi_{s}(x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d)n}$.
Moreover, for the $r$-padded instance $\mathrm{U}(r)$ with $r\geq D$, we have:
  1. (i).
    The graph $\mathrm{U}(r)$ contains $14n+2r+$ 1 nodes and $2n(2D-d+2r)+n$ edges;
  2. (ii).
    For each node $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$, the PPR values are $\pi_{s}(x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d+r)n}$ and $\pi_{s}(x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d+r)n}$, and in particular, $\pi_{s}(x_{2})=\frac{D-d}{D}\cdot\pi_{s}(x_{1})$.
Proof.
The number of nodes and edges in both instances $\mathrm{U}$ and $\mathrm{U}(r)$ follows directly from their definitions. We now compute the PPR values from the source node $s$ on the instance $\mathrm{U}$. First, since each node $y\in Y_{1}$ receives one incoming edge from $s$, and there are $n$ such nodes, it holds that $\pi_{s}(y)=\frac{1-\alpha}{n}$. Next, consider nodes $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Each node in $X_{1}$ receives $D$ in-edges from $Y_{1}$, and each node in $X_{2}$ receives $D-d$ in-edges from $Y_{1}$. Since every $y\in Y_{1}$ has out-degree $2D-d$, the probability of transitioning from a given $y$ to any of its children is $\frac{1}{2D-d}$. By aggregating over all paths from $s$ to $x_{1}$ via $Y_{1}$, we obtain $\pi_{s}(x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d)n}$ and $\pi_{s}(x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d)n}$. The difference is then $\pi_{s}(x_{1})-\pi_{s}(x_{2})=(1-\alpha)^{2}\cdot\frac{d}{(2D-d)n}$. For the padded instance $\mathrm{U}(r)$, the analysis is similar. Each node in $Y$ now has out-degree $2D-d+r$ due to the extra outgoing edges to $Z_{Y}$. Thus, the transition probabilities are updated accordingly, and we obtain $\pi_{s}(x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d+r)n}$ and $\pi_{s}(x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d+r)n}$. In particular, this yields the ratio $\pi_{s}(x_{2})=\frac{D-d}{D}\cdot\pi_{s}(x_{1})$. This completes the proof. ∎
Equipped with the above lemma, we now carefully select the graph parameters $n$, $D$, and $d$, along with the error tolerances $\varepsilon$ and $\delta$, to ensure that the PPR values of nodes in $X_{1}$ and $X_{2}$ exhibit a controlled but noticeable difference.
Lemma 10.
(SSPPR-A) Choose any small constant $\beta\in(0,1)$, decay factor $\alpha$ and any functions $\varepsilon_{0}(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n^{2-\beta})$. For sufficiently large $n_{0}$, there exists graph parameters $n,D,d$ (as function of $n_{0}$) s.t. for all graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following conditions hold:
  1. (i).
    The node count of $\mathrm{U}$ is in $\operatorname{\mathcal{O}}(n_{0})$, and the edge count $\mathrm{U}$ is in $\operatorname{\mathcal{O}}(m_{0})$;
  2. (ii).
    $(n)^{1-\beta}\geq D\geq 2d$ and $d\leq\log(n)$;
  3. (iii).
    $Dn\ \text{is }\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})\right)$;
  4. (iv).
    For all node $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$, their PPR values satisfy
    $\pi_{s}(x_{1})-\pi_{s}(x_{2})>2\varepsilon_{0}.$
Proof.
We manipulate the parameters $n$ and $D$ to fulfill the conclusion. Let $C$ be a constant to be determined later. We set $nD=C\beta^{2}\min\left(n_{0}^{2-\beta},m_{0},\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}}\right)$, $n\geq D$, and $d=\log(n)$. Note that $d\geq\frac{1}{2}\log(nD)$. From Lemma 9, we have:
$\pi_{s}(x_{1})-\pi_{s}(x_{2})$ $=\frac{d}{2D-d}\cdot\frac{(1-\alpha)^{2}}{n}\geq\frac{\log(nD)}{8Dn}\cdot(1-\alpha)^{2}\geq\frac{\log(C\beta^{2}\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})}{8C\beta^{2}\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}}}\cdot(1-\alpha)^{2}$
$\geq\frac{(1-\alpha)^{2}}{8C\beta^{2}}\frac{\log(\frac{1}{\varepsilon_{0}})+log(C\beta^{2})+\log\log(\frac{1}{\varepsilon_{0}})}{\log(\frac{1}{\varepsilon_{0}})}\varepsilon_{0}\geq\frac{(1-\alpha)^{2}}{16C\beta^{2}}\varepsilon_{0}$
The last inequality requires $\log(C\beta)+\log(\frac{1}{\varepsilon_{0}})>\frac{1}{2}\log(\frac{1}{\varepsilon_{0}})$, which further needs $\varepsilon_{0}<(C\beta)^{2}$. Setting $C=\Theta((1-\alpha)^{2})$ directly accomplishes the property $(iv)$. Next, we need to fit the conditions $(i),(ii)$ and $(iii)$ to accomplish the proof. We give an overall setting of
$n=\left(C\beta^{2}\min(n_{0}^{2-\beta},m_{0},\log(\frac{1}{\varepsilon_{0}})/{\varepsilon_{0}})\right)^{\frac{1}{2-\beta}}\ \text{and}\ D=\left(C\beta^{2}\min(n_{0}^{2-\beta},m_{0},\log(\frac{1}{\varepsilon_{0}})/{\varepsilon_{0}})\right)^{\frac{1-\beta}{2-\beta}}.$
Condition $(i)$ holds since the total number of edges is at most $2n(2D-d)+n\leq 4nD=\Theta(\beta m_{0})$, and the number of nodes is $4n+1=\Theta(\beta n_{0})$. For condition (ii), it is clear that $n^{1-\beta}\geq D$, and since $2d=2\log n\leq\log\left((C\beta^{2})^{2/(2-\beta)}n_{0}\right)$, we have $2d\leq(C\beta^{2}n_{0})^{(1-\beta)/(2-\beta)}\leq D$ for sufficiently large $n$. Finally, condition $(iii)$ holds directly by the definition of $nD$. ∎
Lemma 11 (SSPPR-R).
Choose any constant $c\in(0,\frac{1}{2}]$ and any functions $\varepsilon_{0}(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n^{2})$. For sufficiently large $n_{0}$, there exists graph parameters $n,D,d,r$ (as function of $n_{0}$), such that for $r$-padded instance $\mathrm{U}(r)$ based on all graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following conditions hold:
  1. (i).
    The node count of $\mathrm{U}(r)$ is in $\operatorname{\mathcal{O}}(n_{0})$, and the edge count $\mathrm{U}(r)$ is in $\operatorname{\mathcal{O}}(m_{0})$;
  2. (ii).
    $(n)^{\frac{1}{2}}>D\geq\frac{3}{2}d$ and $d\leq\log(n)$;
  3. (iii).
    $n(r+D)=\Omega\left(\min(m,\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$;
  4. (iv).
    For all $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ with PPR values satisfying $\pi_{s}(x_{1})\geq\delta_{0}$ and $\pi_{s}(x_{2})\geq\delta_{0}$, we have
    $\pi_{s}(x_{1})\geq\frac{1}{(1-c)^{2}}\cdot\pi_{s}(x_{2}).$
Proof.
We consider two cases with an overall setting of $\delta$ such that $\frac{\log\frac{1}{\delta}}{\delta}=\min(cm_{0},\frac{\log\frac{1}{\delta_{0}}}{\delta_{0}}).$
Case 1:
$n_{0}\leq\frac{1}{32}(1-\alpha)^{2}\frac{1}{\delta}.$ And we here set $n=n_{0},\quad d=\log(n_{0}),\quad D=\left(1+\frac{1}{4c}\right)d$ and $r=\frac{(1-\alpha)^{2}d}{16cn_{0}\delta}>0.$ Under these parameters, conditions (i), (ii), and (iii) naturally hold when $n$ is sufficiently large. Moreover for condition (iv), we have
$nr=\Omega\left((1-\alpha)^{2}d/\delta c\right)=\Omega\left(\min(m_{0},\frac{\log\frac{1}{\delta_{0}}}{\delta_{0}})\right).$
By Lemma 9, the PPR value for any $x_{2}\in X_{2}$ satisfies
$\pi_{s}(x_{2})=(1-\alpha)^{2}\frac{D-d}{(2D-d+r)n_{0}}=\frac{(1-\alpha)^{2}}{(4c+2)n_{0}+\frac{(1-\alpha)^{2}}{4\delta}}\geq\delta\geq\delta_{0}.$
Case 2:
$n_{0}>\frac{1}{32}(1-\alpha)^{2}\frac{1}{\delta}.$ Under this setting, we set $r=0$ since no padded instance is required here. Further, we give $n=\frac{1}{32}(1-\alpha)^{2}\frac{1}{\delta}$ and $D=\left(1+\frac{1}{4c}\right)d.$ Similarly, for sufficiently large $n$, conditions (i), (ii), and (iii) hold. Additionally, it holds that $nD=nd\left(1+\frac{1}{4c}\right)=\Omega\left(\frac{\log(\frac{1}{\delta})}{\delta}\right)=\Omega\left(\min(m_{0},\frac{\log\frac{1}{\delta_{0}}}{\delta_{0}})\right)$ and $\pi_{s}(x_{2})=\frac{(1-\alpha)^{2}}{n(4c+2)}\geq\delta\geq\delta_{0}$, which ensures condition (iv) and thus we complete the proof. ∎
Lemma 10 shows that the node sets $X_{1}$ and $X_{2}$ differ in their PPR values by more than twice the error parameter $\varepsilon_{0}$. Consequently, any algorithm $\operatorname{\mathcal{A}}$ must distinguish between these two sets with constant success probability, otherwise we argue that it is impossible to distinguish the two node sets with expected query cost at $o(\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}})$ in later sections.
Notice that under the assumption $\min(m_{0},\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}})\leq n_{0}^{2-\beta}$, the instance family we construct contains $nD=\Theta\left(\min(m_{0},\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}})\right)$ edges, which is proportional to the total query complexity stated in Theorem 4. To prove this main theorem, it suffices to show that there exists an absolute constant $C_{A}$ such that any algorithm attempting to solve the SSPPR-A problem with expected query cost at most $C_{A}\cdot(1-p)nD$ must fail on our constructed graph family with probability at least $p$. By Markov’s inequality, we can enforce that the algorithm halts and raises an error if it performs more than $C_{A}\cdot\frac{nD}{\beta}$ queries, and subsequently demonstrate the existence of a graph instance on which the failure probability is at least $\frac{p}{2}$.
A similar argument applies in the SSPPR-R setting, where the expected query budget is bounded by $C_{R}\cdot nD$ for some constant $C_{R}$. To derive these conclusions, we first formalize the execution of a specific query $Q$ and its corresponding response $R$ as an event on an instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ with assigned node and edge labels.
Node and Edge Labeling in $\mathcal{U}(n,D,d)$.
In accordance with the model specified in Definition 7, our graph instances are equipped with both node and edge labels. While our overarching goal is to preserve in-distinguishability among instances within the family, we partially fix the labeling scheme to ensure clarity and consistency throughout the construction process described below:
  1. (i).
    The source node $s$ is assigned label $1$;
  2. (ii).
    Nodes in $X$ are labeled consecutively from $2$ through $12n+$ 1;
  3. (iii).
    Nodes in $Y_{1}$ are labeled from $22n+$ 2 to $13n+$ 1, and nodes in $Y_{2}$ from $23n+$ 2 to $14n+$ 1;
  4. (iv).
    The $i^{\text{th}}$ out-edge of the source node $s$ points to the node labeled $2n+1+i$.
For ease of reference, we adopt the following conventions in place of explicit numeric labels:
  1. (i).
    $s$ denotes the node labeled $1$, which has in-degree zero and out-degree $n$;
  2. (ii).
    $x_{1},x_{2},\ldots,x_{2n}$ refer to the nodes labeled $2$ through $12n+$ 1, each with in-degree $2D-d$ and out-degree zero;
  3. (iii).
    $y_{1},y_{2},\ldots,y_{2n}$ correspond to the nodes labeled $22n+$ 2 through $14n+$ 1, each with in-degree one and out-degree $2D-d$;
  4. (iv).
    We use $y^{(1)}_{1},\ldots,y^{(1)}_{n}$ to refer to nodes $y_{1},\ldots,y_{n}\in Y_{1}$, and $y^{(2)}_{1},\ldots,y^{(2)}_{n}$ for $y_{n+1},\ldots,y_{2n}\in Y_{2}$. For edge references, we denote $(y^{(1)}_{i},k)$ as the $k^{\text{th}}$ out-edge from node $y^{(1)}_{i}\in Y_{1}$, and analogously for nodes in $Y_{2}$.
The above notations serve interchangeably to denote both the node label and the corresponding node in the graph. Note that the labeling scheme clearly identifies whether a node belongs to $Y_{1}$, $Y_{2}$, or $X$, but it does not distinguish between the subsets $X_{1}$ and $X_{2}$. Additionally, once a node’s label is known, its in-degree and out-degree are assumed to be fully determined.
We now illustrate how queries operate on a graph $\mathrm{U}\in\mathcal{U}(n,D,d)$ under this labeling convention. Each query-response pair $(Q,R)$ corresponds to an observable property of $\mathrm{U}$. For example, consider the node labeled $x_{1}\in X$ and the node labeled $y^{(1)}_{1}\in Y_{1}$. Suppose the $3^{\text{rd}}$ in-edge of $x_{1}$ corresponds to the $2^{\text{nd}}$ out-edge of $y^{(1)}_{1}$. Then the query $Q={\textsf{ADJ}}_{\text{in}}(x_{1},3)$ yields the response $R=(y^{(1)}_{1},2)$. Conversely, for any graph $\mathrm{U}_{0}\in\mathcal{U}(n,D,d)$, if executing the query $Q={\textsf{ADJ}}_{\text{in}}(x_{1},3)$ returns the response $R=(y^{(1)}_{1},2)$, then $\mathrm{U}_{0}$ must exhibit exactly this structural property. Under this procedure, any query-response pair $(Q,R)$ is referred to as an event with respect to a graph instance distribution, i.e., $\Sigma(n,D,d)$ (defined later in Section 4.2). For any $\sigma\sim\Sigma(n,D,d)$, it corresponds to a graph $U_{\sigma}$ with fixed edges and node labels. For brevity, we write such an event as $Q=R$ when the context is clear.
4.2 Distribution $\Sigma(n,D,d)$
Figure 2: An example for generating graph distribution $\Sigma(n,D,d)=\Sigma(2,2,1)$.
We now describe the procedure for generating an instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ with assigned node and edge labels, which together define the distribution $\Sigma(n,D,d)$ over all possible label permutations. An illustrative example is provided in Figure 2. Intuitively, $\Sigma(n,D,d)$ is constructed by randomly splitting the node set $X$, thereby determining which node labels in $X$ belong to $X_{1}$ or $X_{2}$, and then independently sampling every possible edge permutation that is consistent with the hard-instance template. The complete construction procedure is detailed as follows:
  1. (i).
    Split of $X$. We first choose a split $\mathrm{SP}$ of the index set $\{1,\dots,2n\}$ into two disjoint subsets of size $n$, which determines the split of the node set $X$ into $X_{1}$ and $X_{2}$. Specifically, the nodes labeled $x_{1},\dots,x_{2n}$ are assigned to $X_{1}$ or $X_{2}$ according to the selected split. We denote the set of all possible such splits by $\mathrm{SP}\in\binom{X}{n}$. For example, in Figure 2, the last column lists nodes $x_{1},\dots,x_{4}$ (from top to bottom), and the chosen split $\mathrm{SP}=\{1,3\}$ assigns nodes labeled $x_{1}$ and $x_{3}$ to $X_{1}$, while nodes $x_{2}$ and $x_{4}$ belong to $X_{2}$;
  2. (ii).
[…]
  3. (iii).
    Permuting out-edges of nodes in $Y_{1}$ and $Y_{2}$. We apply two independent permutations $\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}\in\mathrm{S}^{n(2D-d)}$ over the set $\{1,\dots,n(2D-d)\}$ to determine the ordering of the out-edges from nodes in $Y_{1}$ and $Y_{2}$, respectively. These permutations govern whether the $k^{\text{th}}$ out-edge of a node $y^{(1)}_{i}\in Y_{1}$ (or $y^{(2)}_{i}\in Y_{2}$) connects to a node in $X_{1}$ or $X_{2}$. Specifically, the $k^{\text{th}}$ out-edge of $y^{(1)}_{i}$ points to a node in $X_{1}$ if $\mathrm{P}_{Y_{1}}^{n(2D-d)}[(2D-d)(i-1)+k]\leq nD$; otherwise, it connects to a node in $X_{2}$. An analogous rule applies for nodes in $Y_{2}$ via $\mathrm{P}_{Y_{2}}^{n(2D-d)}$. We denote the collections of such permutations as $\mathrm{S}^{n(2D-d)}(Y_{1})$ and $\mathrm{S}^{n(2D-d)}(Y_{2})$, respectively. This step corresponds to determining connections between virtual nodes in the second and third columns of Figure 2. For example, if we choose $\mathrm{P}_{Y_{1}}^{n(2D-d)}=(1,3,2,4,5,6)$, then the first four virtual nodes in the second column represent out-edges connecting to nodes in $X_{1}$, as determined by the corresponding permutation values;
  4. (iv).
[…]
    From previous steps, the $nD$ selected out-edges from $Y_{1}$ are encoded as a set of tuples $\{(y^{(1)}_{i_{j}},k_{j})\}_{j=1}^{nD}$, where each tuple represents the $k_{j}^{\text{th}}$ out-edge of node $y^{(1)}_{i_{j}}$. Likewise, the $nD$ selected in-edges to $X_{1}$ are represented as $\{(x^{(1)}_{i_{j}^{\prime}},k_{j}^{\prime})\}_{j=1}^{nD}$, with each tuple representing the $k_{j}^{\prime\text{th}}$ in-edge of node $x^{(1)}_{i_{j}^{\prime}}$. Both edge sets are naturally ordered; for instance, each out-edge $(y^{(1)}_{i_{j}},k_{j})$ can be uniquely indexed by the expression $(2D-d)i_{j}+k_{j}$. The bijection $\mathrm{S}(Y_{1},X_{1})$ then defines the matching: the $k_{j}^{\text{th}}$ out-edge of node $y^{(1)}_{i_{j}}$ is connected to the $(k_{j}^{\prime})^{\text{th}}$ in-edge of node $x^{(1)}_{i_{j}^{\prime}}$. In Figure 2, the mapping $\mathrm{S}(Y_{1},X_{1})=(4,3,2,1)$ is visually indicated by edges between the virtual nodes (shown as black-filled) in the third and fourth columns.
By the above construction, we can then derive the following lemmas and definitions, which recover a graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ accordingly.
Lemma 12.
For all possible $\mathrm{SP},\mathrm{P}_{i}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)},\mathrm{P}\in\binom{X}{n}\times\mathrm{S}^{2D-d}(X)\times\mathrm{S}^{n(2D-d)}(Y_{1})\times\mathrm{S}^{n(2D-d)}(Y_{2})\times\Pi_{Y_{a}\in\{Y_{1},Y_{2}\},X_{b}\in\{X_{1},X_{2}\}}S(Y_{a},X_{b})$: the above procedure links a legal graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$.
Definition 13.
$\Sigma(n,D,d)$ is defined as the uniform distribution over $\binom{X}{n}\times\mathrm{S}^{2D-d}(X)\times\mathrm{S}^{n(2D-d)}(Y_{1})\times\mathrm{S}^{n(2D-d)}(Y_{2})\times\Pi_{Y_{i}\in\{Y_{1},Y_{2}\},X_{i}\in\{X_{1},X_{2}\}}S(Y_{i},X_{i})$. For $\sigma\sim\Sigma(n,D,d)$, we use $\mathrm{U}_{\sigma}$ to denote the corresponding labeled graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$. For ease of presentation, we write $\mathrm{U}\sim\Sigma(n,D,d)$ to denote a graph instance sampled from this distribution.
Based on the above construction and definitions, we now formulate the probability that a graph instance contains parallel edges in the following lemma:
Lemma 14.
(proof in Appendix A.2) $\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\bigl{[}\text{edge multiplicity of }\mathrm{U}\geq L\bigr{]}\leq 4n^{2}(\frac{4D}{n})^{L}$, for $1\leq L\leq D$.
Lemma 14 demonstrates that, with high probability, the edge multiplicity in our constructed graph instances is bounded by a small constant. In other words, high multiplicity edges rarely appear under our construction. Leveraging this result, we show in the next section that any algorithm designed for simple graphs can be adapted to handle multi-graphs with bounded edge multiplicity. From Lemmas 11 and 10, since $D$ is upper bounded by a sublinear polynomial in $n$ (see condition (ii) in both lemmas), we can choose a constant $L$ such that
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\bigl{[}\text{edge multiplicity of }\mathrm{U}>L\bigr{]}\leq\frac{1}{n}.$
In particular, we set $L=\frac{8}{\beta}$ for the SSPPR-A setting and $L=16$ for SSPPR-R.
4.3 Establishing Main Results
To prove our main results, such as Theorem 3, we assume the existence of an algorithm $\operatorname{\mathcal{A}}$ that successfully answers the SSPPR-R problem with lower query complexity than our target lower bound. We then derive a contradiction by showing that the overall success probability over the graph distribution becomes $\operatorname{\mathcal{O}}(1/n)$, which violates the requirement of constant success probability. Note that $\operatorname{\mathcal{A}}$ is designed for simple graphs, whereas instances $\mathrm{U}\sim\Sigma(n,D,d)$ may admit parallel edges. Fortunately, our earlier analysis guarantees that the edge multiplicity is bounded by a constant with high probability. The following theorem provides a mechanism for executing $\operatorname{\mathcal{A}}$ on such multi-graphs.
Lemma 15.
(proof in Appendix A.3) [Reduction from simple to multi-graphs] Choose any constant $L$, decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, constant $\beta\in(0,1)$, and any function $\varepsilon(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2-\beta}),T(n_{0})>0$ (Or, error bound $c\in(0,\frac{1}{2}]$, and any function $\delta(n_{0})\in(0,1),m(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2})$ for SSPPR-R setting). Suppose there exists an algorithm $\operatorname{\mathcal{A}}$ such that, for any simple directed graph with $\Theta(Ln_{0})$ nodes and $\Theta(Ln_{0}+Lm_{0})$ edges, $\operatorname{\mathcal{A}}$ answers the SSPPR-A problem with decay factor $\alpha$, error $\varepsilon$ and failure probability bounded by $p$ (or SSPPR-R problem with threshold $\delta$, error bound $c$), and expected query complexity $T$ in arc-centric model.
Then there exists an algorithm $\operatorname{\mathcal{A}}_{L}$ in arc-centric model such that, for any integers $n\leq n_{0},\;D,\;d,\;r,\;n(D+r)\leq m$, and any $\mathrm{U}\in\mathcal{U}(n,D,d)$ with edge multiplicity bounded by $L$, $\operatorname{\mathcal{A}}_{L}$ solves the SSPPR-A problem with decay factor $1-(1-\alpha)^{3}$ and error parameter $\varepsilon$ (or SSPPR-R problem with threshold $\delta$, error requirement $c$) on $\mathrm{U}(r)$ for target nodes in $X$, with expected query complexity bounded by a function of $n_{0}$ in $\operatorname{\mathcal{O}}(T)$ and with failure probability bounded by $p$. Specifically, $\mathcal{A}_{L}$ executes each query ${\textsf{ADJ}}_{in/out}(v,k)$ by a uniformly randomly selection of an uncovered label $k$ among $v$’s in or out edges.
Combining all the preceding results, we arrive at the following theorem, which encapsulates the current milestone of our analysis:
Theorem 16.
[…]
  1. (i).
    For any $\mathrm{U}\in\mathcal{U}(n,D,d)$, $\mathrm{U}(r)$ contains no more than $\beta n_{0}$ nodes and no more than $\beta m_{0}$ edges;
  2. (ii).
    $d\leq\log(n),n\geq 2D,D\geq\frac{3}{2}d$;
And there exists an algorithm $\mathcal{A}^{{}^{\prime}}$ holding that:
  1. (iii).
    $\mathcal{A}^{{}^{\prime}}$ does at most $T(n_{0})\leq\gamma nD$ queries;
  2. (iv).
    For $\mathrm{U}\sim\Sigma(n,D,d)$, the algorithm $\mathcal{A}^{{}^{\prime}}$ can be applied on any multi-graph instance $\mathrm{U}$, with expected probability at least $\frac{1-p}{4}$, $\mathcal{A}^{{}^{\prime}}$ can output the correct split $X_{1},X_{2}$ of $X$.
Proof.
We begin with the SSPPR-R case. For the graph related parameters, Lemma 11 gives the integers $n,\;D,\;d,\;r$, directly satisfying the condition $(i)$ and $(ii)$ with $n(D+r)=\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\delta})}{\delta_{0}})\right)$. Next, we’d construct algorithm $\operatorname{\mathcal{A}}^{\prime}$ to fit the condition $(iii)$ and $(iv)$. The steps are described as:
  1. (i).
    $\mathcal{A}$ to $\mathcal{A}_{L}$ for executable on multi graph $\mathrm{U}(r)$. As established in our earlier analysis, for $L=16$, a graph sampled from the distribution $\Sigma(n,D,d)$ admits edge multiplicity at most $L$ with high probability, specifically at least $1-\frac{1}{n}$. By invoking Theorem 15, this guarantees the existence of an algorithm $\mathcal{A}_{L}$ that can solve the SSPPR-R problem on such multi-graphs with expected success probability at least $1-p$, with expected query complexity of $o(n(D+r))$.
  2. (ii).
    From PPR estimation to split of $X$. To ensure that $\mathcal{A}_{L}$ produces PPR estimates of nodes in $X$ within the specified relative error tolerance $c$, it must be capable of distinguishing the two subsets $X_{1}$ and $X_{2}$, whose PPR values differ by a clear, quantifiable gap. We therefore append a simple post-processing step to $\mathcal{A}_{L}$ that outputs a split of $X$ based on the estimated PPR values. This split is correct if the estimates produced by $\mathcal{A}_{L}$ satisfy the SSPPR-R accuracy requirements. Specifically, by Lemma 11, for any $x_{i_{1}}\in X_{1}$ and $x_{i_{2}}\in X_{2}$, we have $\pi_{s}(x_{i_{1}})\geq\frac{1}{(1-c)^{2}}\pi_{s}(x_{i_{2}})$. This implies the existence of a threshold $\pi_{0}$ such that for all $x_{i_{1}}\in X_{1}$, $\pi_{s}(x_{i_{1}})(1-c)>\pi_{0}$, and for all $x_{i_{2}}\in X_{2}$, $\pi_{s}(x_{i_{2}})(1+c)<\pi_{0}$. Furthermore, Lemma 11 ensures that $\pi_{s}(x_{i})\geq\delta$ for all $x_{i}\in X$. If $\mathcal{A}_{L}$ succeeds on $\mathrm{U}(r)$, its estimate $\widehat{\pi}_{s}(x_{i})$ must satisfy $|\widehat{\pi}_{s}(x_{i})-\pi_{s}(x_{i})|\leq c\pi_{s}(x_{i})$. As a result, for all $x_{i_{1}}\in X_{1}$, we have $\widehat{\pi}_{s}(x_{i_{1}})\geq(1-c)\pi_{s}(x_{i_{1}})>\pi_{0}$, and for all $x_{i_{2}}\in X_{2}$, we have $\widehat{\pi}_{s}(x_{i_{2}})\leq(1+c)\pi_{s}(x_{i_{2}})<\pi_{0}$. Thus, a simple thresholding rule at $\pi_{0}$ suffices to correctly separate $X$ into $X_{1}$ and $X_{2}$.
  3. (iii).
    From queries on $\mathrm{U}(r)$ to queries on $\mathrm{U}$. In the previous step, we constructed an algorithm that successfully distinguishes between $X_{1}$ and $X_{2}$ using an expected $o(n(D+r))$ queries on the $r$-padded graph $\mathrm{U}(r)$, under the assumption that the edge multiplicity of the underlying graph $\mathrm{U}$ is bounded by $L=16$ with high probability $1-\frac{1}{n}$. We now demonstrate that queries involving the auxiliary padding nodes in $Z_{X}$ and $Z_{Y}$ are not essential. Consequently, the algorithm can be adapted to the original graph $\mathrm{U}$ with an expected query complexity of $o(nD)$.
    Recall that we have not fixed a specific labeling scheme for $\mathrm{U}(r)$ in earlier steps; instead, we simply assigned labels to the padding nodes $Z_{X}$ and $Z_{Y}$ as $14n+2,\dots,4n+2r+$ 1. The out-edges from each node $y_{i}\in Y$ to $Z_{Y}$ are assigned labels $2D-d+1,\dots,2D-d+r$, and the in-edges from $Z_{X}$ to each $x_{i}\in X$ are labeled analogously. Given this fixed labeling, all information regarding $Z_{X}$ and $Z_{Y}$ is known in advance from the perspective of $\mathcal{A}_{L}$. Since $\mathcal{A}_{L}$ operates only on $\mathrm{U}(r)$, these edges can be hard-coded into the algorithm. We initialize a subgraph $G_{0}$ internal to $\mathcal{A}_{L}$ that stores all such predetermined edges in $\mathrm{U}(r)$—specifically, all the nodes $s$, $X$, $Y$, $Z_{X}$, and $Z_{Y}$, along with the edges from $Z_{X}$ to $X$ and from $Y$ to $Z_{Y}$.
    We then define the modified algorithm $\mathcal{A}^{\prime}_{L}$ as follows. The core idea is to simulate $\mathcal{A}_{L}$ on the original graph $\mathrm{U}$ by internally emulating its behavior on the padded graph $\mathrm{U}(r)$. Whenever $\mathcal{A}_{L}$ issues a degree query or an $\mathsf{Adj}$ query involving a node in the padding sets $Z_{X}\cup Z_{Y}$, $\mathcal{A}^{\prime}_{L}$ returns the predetermined response encoded in $G_{0}$, which is consistent across all $\mathrm{U}(r)$ constructed from some $\mathrm{U}\in\mathcal{U}(n,D,d)$. In the case where $\mathcal{A}_{L}$ issues a query of the form $\mathsf{Adj}^{\mathrm{in}}(x_{i},k)$, we proceed as follows: if $k>2D-d$, the query refers to a synthetic edge from $Z_{X}$, so we return the $(k-2D+d+1)$-st node in $Z_{X}$; otherwise, we forward the actual query $\mathsf{Adj}^{\mathrm{in}}(x_{i},k)$ to the graph $\mathrm{U}$, and add the resulting edge to the internal subgraph $G_{0}$. In this way, $\mathcal{A}^{\prime}_{L}$ simulates the interaction between $\mathcal{A}_{L}$ and $\mathrm{U}(r)$ using queries exclusively on $\mathrm{U}$. Therefore, with success probability at least $p$, $\mathcal{A}^{\prime}_{L}$ correctly outputs the split $(X_{1},X_{2})$.
  4. (iv).
    Bounding query complexity of $\mathcal{A}_{L}^{{}^{\prime}}$. We now bound the query complexity of $\mathcal{A}^{\prime}_{L}$. As described in Theorem 15, the algorithm $\mathcal{A}_{L}$ issues queries of the form ${\textsf{ADJ}}_{\mathrm{out}/\mathrm{in}}(v,k)$ by selecting $k$ uniformly at random from the un-queried incident edge labels. For any node $v\in\mathrm{U}$ where fewer than $(D+r)/2$ incident labels have been examined, the probability that a new query returns a real (non-padding) edge is at most $(2D-d)/(2D-d+r-(D+r)/2)$. Consequently, the expected number of genuine adjacency queries issued by $\mathcal{A}^{\prime}_{L}$ can be bounded by the sum of two terms: the first is $(2D-d)\cdot\mathbb{E}[|\{v\in\mathrm{U}\mid\text{at least }(D+r)/2\text{ incident labels examined}\}|]$; the second is $(2D-d)/(2D-d+r-(D+r)/2)\cdot o((D+r)n)$. Since both terms are asymptotically $o(nD)$, the overall expected query complexity of $\mathcal{A}^{\prime}_{L}$ is $o(nD)$.
For notational simplicity, we refer to the final algorithm constructed in the above theorem as $\operatorname{\mathcal{A}}$. Naturally, we decompose $\operatorname{\mathcal{A}}$ into two components: the first part, denoted $\mathcal{A}_{1}$, is an exploration procedure that performs queries on the graph; the second part, $\mathcal{A}_{2}$, processes the information collected and outputs a split of $X$ into subsets $X_{1}$ and $X_{2}$. The algorithm $\operatorname{\mathcal{A}}$ succeeds if and only if $\mathcal{A}_{2}$ outputs a correct split. Since we are working under the RAM model, we can assume that $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ operate independently—that is, $\mathcal{A}_{2}$ is a separate algorithm whose input is the data written into memory by $\mathcal{A}_{1}$, and whose output is the inferred split of $X$. Given that RAM is finite (though arbitrarily large) and that the only inputs to $\mathcal{A}_{1}$ are the query responses $R_{1},\ldots,R_{T}$, we may treat the final RAM state as a function of the query-response sequence $Q_{1},\ldots,Q_{T},R_{1},\ldots,R_{T}$, along with any internal randomness. According to the definition in 6, this tuple $(Q_{1},\ldots,Q_{T},R_{1},\ldots,R_{T})$ constitutes the observation, while the correct split of $X$ represents the prior distribution. The expected success probability of $\operatorname{\mathcal{A}}$ can thus be formulated as:
$\mathbb{E}_{\mathrm{U}\sim\Sigma(n,D,d)}\bigl{[}\operatorname{{\bf Pr}}[\operatorname{\mathcal{A}}\text{ succeeds on }U]\bigr{]}\leq\sum_{Q_{1}=R_{1},\cdots,Q_{T}=R_{T}}\operatorname{{\bf Pr}}_{\operatorname{\mathcal{A}},{\mathrm{U}\sim\Sigma(n,D,d)}}[Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}]$
×$\times\max_{\mathrm{\mathrm{SP}}\in\binom{X}{n}}\operatorname{{\bf Pr}}_{\mathcal{U}\sim\Sigma(n,D,d)}[\mathrm{SP}\text{ is the correct splitof }\mathrm{U}\mid Q_{1}=R_{1},\cdots,Q_{T}=R_{T}].$ (1)
Note that the notion of an event $Q=R$ was defined in the final part of Section 4.1. Our goal now is to evaluate the conditional probability
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\left[\mathrm{SP}\text{ is the correct split of }\mathrm{U}\,\middle|\,Q_{1},\ldots,Q_{T};\,R_{1},\ldots,R_{T}\right],$
and to show that it is bounded by $\frac{1}{n}$. This result is formally captured in the following Lemma 17.
Lemma 17.
(proof in Appendix A.4) There is a constant $\gamma$, such that. For sufficiently large $n$, $d\leq log(n)$, $D>\frac{3d}{2}$, $T\leq\gamma nD$. Given any query and returning sequence $Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}$, and a split of $X$, $\mathrm{SP}$ we have:
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}[\mathrm{SP}\text{ is thecorrect split of }\mathrm{U}\mid Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}]\leq\frac{1}{n}.$
The intuition for why this lemma holds is that, considering two naive ways of exploring $\mathrm{U}$, both require $\Theta(nD)$ queries to get the correct split: The first way is a traversal of the graph. For instance in a breadth-first search, one need expected $\Omega(nD)$ queries to get all the information. The second approach is a random walk. For each node $x\in X$, there are at least $D-d$ edges from $Y_{1}$ to $X$ and at least $D-d$ edges from $Y_{2}$ to $x$, the key for determining whether $x$ is in $X_{1}$ or $X_{2}$ lies in the left $d$ edges. Notice that each random walk of length $1$ from $X$ has probability at most $\frac{d}{D}$ to find one these key edges. Assume we have done $o(D)$ such walks at each point, the probability of finding no key-edges at some node is at least $n(1-\frac{d}{D})^{o(D)}$. Since $d\leq log(n),n(1-\frac{d}{D})^{o(D)}=\Omega(1)$. Details can be found in Appendix A.4.
By Lemma 17, we then conclude that the expected success probability of $\operatorname{\mathcal{A}}$ is less than $\frac{1}{n}$ under the total query limits $T\in o\left(\min(m_{0},\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$ for SSPPR-R problem and $T\in o\left(\min(m_{0},\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})\right)$ for SSPPR-A query. This contradiction illustrates that the assumption in Theorem 16 does not hold. Thus, it requires $\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})\right)$ queries for SSPPR-R and $\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$ for SSPPR-A problems in expectation, which finally leads to our main lower bound results in Theorem 3 and Theorem 4.
5 Conclusion
In this paper, we focus on the well-studied Single Source Personalized PageRank (SSPPR) computation problem under both additive and relative error settings. Compared to the best-known lower bound results, we achieve tighter bounds as follows:
  • For SSPPR-R, we prove an improved lower bound of $\Omega\left(\min(m,\frac{\log(\frac{1}{\delta})}{\delta})\right)$ for any $\delta\in(0,1)$.
  • For SSPPR-A, we establish an overall lower bound of $\Omega\left(\min(m,\frac{\log(\frac{1}{\varepsilon})}{\varepsilon})\right)$ for any $\varepsilon\in(0,1),m\in O(n^{2-\beta})$ for any arbitrary small constant $\beta\in(0,1)$.
To elaborate, we improve the best-known lower bound results by a logarithmic factor $\log\left(\frac{1}{\delta}\right)$ or $\log\left(\frac{1}{\varepsilon}\right)$, and achieve matching lower and upper bounds for the SSPPR-R query when $\delta$ is approximately $\frac{1}{n}$, which corresponds to the most practical regime in SSPPR computations. This advancement provides significant insights and inspiration for the research community to further design efficient algorithms and deepen the theoretical understanding of PPR.
Appendix A Appendix
A.1 Upper Bounds Discussion
Since there lacks a unified summarization regarding the computational upper bounds for SSPPR-A and SSPPR-R queries, we first review the three basic PPR computation components: Monte Carlo (MC), Forward Push (FP), and Backward Push, and formulate the best-known upper bounds of later state-of-the-art (sota) methods.
A.1.1 Basic Components
MC.
The Monte Carlo method (MC) is a widely used technique for estimating desired values via sample-based computations. Leveraging the interpretation of Personalized PageRank (PPR) as $\alpha$-decay random walks, MC effectively approximates the SSPPR vector. Despite its apparent simplicity, MC provides unbiased and accurate PPR estimates with high probability, potentially surpassing theoretical lower bounds established for deterministic algorithms. The core idea of MC is to simulate $\alpha$-decay random walks starting from node $s$ and use the empirical distribution of their termination nodes as an approximation for $\pi_{s}(\cdot)$. This approach offers a straightforward and intuitive solution to the SSPPR query. Assuming $W$ independent random walks, the expected computational time is $\Theta(W/\alpha)$. By applying Chernoff bounds [8] and union bounds over all potential nodes, setting the number of random walks to $W=\Theta\left(\frac{\log(n/p_{f})}{c^{2}\delta}\right)$ suffices to answer the SSPPR-R query with relative error guarantees. Note that the $\log(n)$ term arises from the union bound over all $n$ nodes. Since only nodes with PPR value larger than $\delta$ are of interest, their number is at most $\frac{1}{\delta}$. This results in an expected computation time of $\operatorname{\mathcal{O}}\left(\frac{\log\left(1/\delta p_{f}\right)}{\alpha c^{2}\delta}\right)=\operatorname{\mathcal{O}}\left(\frac{\log\left(1/\delta\right)}{\delta}\right)$. Similarly, MC can be adapted to answer the SSPPR-A query with high probability in time $\operatorname{\mathcal{O}}\left(\frac{\log\left(1/\varepsilon\right)}{\varepsilon^{2}}\right)$.
Forward Push (FP).
FP [1] is a celebrated “local-push” algorithm for the SSPPR query, which takes as input a parameter $r_{\max}$ and runs in $O(1/r_{\max})$ time. As detailed in Algorithm 1, it simulates random walks deterministically by transport graph traversal probability and obtains an estimation $\widehat{\pi}_{s}(\cdot)$ which satisfies an overall approximation quality with $||\widehat{\pi}_{s}(\cdot)-\pi_{s}(\cdot)||_{1}=r_{sum}=\sum_{u\in V}r_{s}(u)\leq m\cdot r_{max}$. To meet the error requirement of SSPPR-A, $r_{max}$ is required to set to $\varepsilon/m$ and results a pessimistic bound of $\operatorname{\mathcal{O}}(m/\varepsilon)$.
Backward Push (BP).
BP [20] is also a “local-push” algorithm, while for approximating $\pi(v,t)$ from all $v\in V$ to a given target node $t\in V$, known as the Single-Target PPR (STPPR) query. It works similarly as FP while from a inverse direction as described in Algorithm 2 and BP can also utilized to answer the SSPPR-A query by performing it with $r_{\max}=\varepsilon$ for each $t\in V$, resulting in a complexity of $O(m/\varepsilon)$, too. This bound is inferior but latter methods will modify it to achieve better results.
Input: graph $G$, source node $s$, decay factor $\alpha$ and error threshold $r_{max}$.
Output: estimation $\widehat{\pi}_{s}(\cdot)$ and resulted residue $r_{s}(\cdot)$.
$r_{s}(s)\leftarrow 1;r_{s}(u)\leftarrow 0$ for all $u\neq s$;
$\widehat{\pi}_{s}(u)\leftarrow 0$ for all $u\in V$;
while $\exists u\in V\ s.t.\ r_{s}(u)\geq r_{max}/d_{out}(u)$ do
       for each $v\in N_{out}(u)$ do
             $r_{s}(v)\leftarrow r_{s}(v)+(1-\alpha)\cdot r_{s}(u)/d_{out}(u)$;
       end for
      $\widehat{\pi}_{s}(u)\leftarrow\widehat{\pi}_{s}(u)+\alpha\cdot r_{s}(u);$
       $r_{s}(u)\leftarrow 0;$
end while
return $\widehat{\pi}_{s}(\cdot)$ and $r_{s}(\cdot)$
Algorithm 1 Forward Push [1] (FP)
Input: graph $G$, target node $t$, decay factor $\alpha$ and error threshold $r_{max}$.
Output: estimation $\widehat{\pi}_{t}(\cdot)$ and resulted residue $r_{t}(\cdot)$.
$r_{t}(t)\leftarrow 1;r_{t}(u)\leftarrow 0$ for all $u\neq t$;
$\widehat{\pi}_{t}(u)\leftarrow 0$ for all $u\in V$;
while $\exists u\in V\ s.t.\ r_{t}(u)\geq r_{max}$ do
       for each $v\in N_{in}(u)$ do
             $r_{t}(v)\leftarrow r_{t}(v)+(1-\alpha)\cdot r_{t}(u)/d_{out}(v)$;
       end for
      $\widehat{\pi}_{t}(u)\leftarrow\widehat{\pi}_{t}(u)+\alpha\cdot r_{t}(u);$
       $r_{t}(u)\leftarrow 0;$
end while
return $\widehat{\pi}_{t}(\cdot)$ and $r_{t}(\cdot)$
Algorithm 2 Backward Push [20] (BP)
A.1.2 SOTA Methods
FORA [48].
The FORA method is a representative algorithm for answering the SSPPR-R query with sub-linear time complexity. It adopts a two-phase framework by combining the FP and MC components. Generally speaking, it conducts random walks deterministically through the FP method before doing so randomly through MC, aiming to reduce the number of samples needed in MC. In detail, it first uses the FP method obtain an initial estimation. To further eliminating the error for each specific node, it performs MC on nodes with non-zero residues. Specifically, for each node $u$ with $r_{s}(u)>0$, it generates $W_{u}=\lceil r_{s}(u)\cdot W\rceil$ $\alpha$-decay random walks. Denoted $w(u,v)$ the number of such walks stops at $v$, it will add on the estimator $\widehat{\pi}_{s}(v)$ by $\frac{r_{s}(u)\cdot w(u,v)}{W_{v}}$ and the final estimation $\widehat{\pi}^{\prime}_{s}(v)$ is calculated as
[…]
The FP part requires $\operatorname{\mathcal{O}}(\frac{1}{r_{max}})$ running costs and the MC phase needs to generate at most $\operatorname{\mathcal{O}}(m\cdot W\cdot r_{max})$ random walks. Thus the overall expected time complexity can be minimized by setting $r_{max}=\frac{1}{\sqrt{m\cdot W}}$, which results in the $\operatorname{\mathcal{O}}\left(\frac{1}{c}\sqrt{\frac{m\log\left(n/p_{f}\right)}{\delta}}\right)=\operatorname{\mathcal{O}}\left(\sqrt{\frac{m\log\left(n\right)}{\delta}}\right)$ bound.
SpeedPPR [38].
The SpeedPPR method is another sota method which can answer the SSPPR-A query efficiently. Similar to the concept of FORA, it also incorporates the phases of push and random walk. While it discovers that the push order in FP will greatly influence the decreasing rate of remaining residues. The arbitrary push operation ordering of FP makes the analysis of the error bound during the algorithm very challenging. Instead, it conducts the the push operations and makes the entire process more structured to bound the decrease rate of the residues’ summation $r_{sum}$ in an exponential manner feasibly. This results in a different complexity of $\operatorname{\mathcal{O}}(m\cdot\log(\frac{1}{m\cdot r_{max}})+m)$ in the push phase. By setting $r_{max}=m/\varepsilon$, a superior $\operatorname{\mathcal{O}}\left(m\log(\frac{1}{\varepsilon})\right)$ time bound is achieved for SSPPR-A query. Furthermore by combination with MC where for each node $u$ with non-zero residue $r_{s}(u)>0$, perform $\lceil r_{s}(u)\cdot W\rceil\leq\lceil d_{u}\cdot r_{max}\cdot W\rceil$ random walks, the overall complexity
$\operatorname{\mathcal{O}}\left(m\cdot\log\left(\frac{1}{m\cdot r_{max}}+m+m\cdot r_{max}\cdot W\right)\right)$
is minimized by setting $r_{max}=1/W$, resulting in $\operatorname{\mathcal{O}}\left(m\log(\frac{1}{\varepsilon})\right)$ time cost for SSPPR-R query.
ABP[47].
The ABP method is a recent sota method derived to answer the SSPPR-A query in sublinear time. Notably that it combines MC with an adaptive version of Backward Push by conducting push operation through a candidate node sets. This candidate node set is constructed by first running a $\operatorname{\mathcal{O}}(\log(n)/\varepsilon)$ random walks to obtain a rough estimation of $\pi_{s}^{\prime}(\cdot)$. Then, the adaptive backward push will iteratively halves it $r_{max}$ parameter until the cost exceeding a pre-defined threshold. Finally, it runs extra random walks to refine its estimation to meet the additive error requirement and achieves an overall expected time cost of $\operatorname{\mathcal{O}}\left(\frac{\sqrt{m\log(n)}}{\varepsilon}\right)$ for SSPPR-A query.
A.2 Proof of Lemma 14
Lemma.
$\operatorname{{\bf Pr}}_{\mathrm{U}(r)\sim\Sigma(n,D,d)}[\mathrm{U}(r)\text{'sedge multiplicity is at least $L$}]\leq 4n^{2}(\frac{4D}{n})^{L}$, for $L\leq D$.
Proof.
Fix an arbitrary realization of $\left(\mathrm{SP},\,\mathrm{P}^{2D-d}_{1},\dots,\mathrm{P}^{2D-d}_{2n},\,\mathrm{P}^{n(2D-d)}_{Y_{1}},\,\mathrm{P}^{n(2D-d)}_{Y_{2}}\right)$, i.e., the split of $X$ and all in-/out-edge permutations chosen in Steps (i)–(iii) as described in Section 4.2. These choices determine the following graph information:
  • The set of labels for out-edges starting from nodes in $Y_{1}$ and pointing to nodes in $X_{1}$;
  • For every $x_{i}\in X_{1}$, the set of labels of its in-edges originating from $Y_{1}$.
The only remaining randomness that may affect parallel edge multiplicity arises from the uniform bijection $\mathrm{S}(Y_{1},X_{1})$. As described in Step (iv), this bijection maps the set of all out-edge labels from nodes in $Y_{1}$ directed toward $X_{1}$ to the set of all in-edge labels of nodes in $X_{1}$ originating from $Y_{1}$, thereby determining the actual edge connections. Fix a pair of nodes $x_{i}\in X_{1}$ and $y_{j}\in Y_{1}$ (an analogous argument applies to the cases $(X_{1},Y_{2})$, $(X_{2},Y_{1})$, and $(X_{2},Y_{2})$). Let $\mathrm{LABEL}_{\text{out}}(y_{j},X_{1})$ denote the set of at most $2D-d$ out-edge labels of $y_{j}$ that are directed toward $X_{1}$, and let $\mathrm{LABEL}_{\text{in}}(Y_{1},x_{i})$ denote the set of $D$ in-edge labels of $x_{i}$ originating from nodes in $Y_{1}$. We also define $\mathrm{LABEL}_{\text{out}}(Y_{1},X_{1})$ as the complete set of out-edge labels from all nodes in $Y_{1}$ targeting $X_{1}$, where each label includes both the source node $y_{j^{\prime}}^{(1)}$ and its local out-edge label. Since $\mathrm{S}(Y_{1},X_{1})$ is a uniformly random bijection over two sets of size $nD$, for any given $L$-subset of labels in $\mathrm{LABEL}_{\text{out}}(y_{j},X_{1})$ (i.e., any collection of $L$ out-edges from node $y_{j}$), its image under $\mathrm{S}^{-1}(Y_{1},X_{1})$ is distributed uniformly over all possible $L$-subsets of $\mathrm{LABEL}_{\text{out}}(Y_{1},X_{1})$. Therefore, for any particular $L$-subset $F\subset\mathrm{LABEL}_{in}(Y_{1},x_{i})$, we have that
$\operatorname{{\bf Pr}}\bigl{[}\mathrm{S}^{-1}(Y_{1},X_{1})(F)\subseteq\mathrm{LABEL}_{out}(y_{j},X_{1})\bigr{]}\;=\;\frac{\binom{D}{L}}{\binom{nD}{L}}.$
There are at most $\binom{2D-d}{L}$ ways to choose such an $F$. A union bound therefore gives
[…]
When $n\geq 4$ (so that $nD\geq 4D\geq 2L$), we have the bounds $\binom{D}{L}\leq\frac{D^{L}}{L!}$, $\binom{2D}{L}\leq\frac{(2D)^{L}}{L!}$, and $\binom{nD}{L}\geq\frac{(nD/2)^{L}}{L!}$. These imply $\binom{2D}{L}\cdot\frac{\binom{D}{L}}{\binom{nD}{L}}\leq\left(\frac{2D}{n}\right)^{L}$. The same inequality holds for all other combinations of $(x_{i},y_{j})$, including $x_{i}\in X_{2}$ or $y_{j}\in Y_{2}$. Since there are only $2n$ nodes in $X$ and $2n$ nodes in $Y$, we conclude that
$\operatorname{{\bf Pr}}_{\mathrm{U}(r)\sim\Sigma(n,D,d)}\left[\text{edge multiplicity of }\mathrm{U}\text{ is at least }L\right]\leq 4n^{2}\left(\frac{4D}{n}\right)^{L},$
which completes the proof.
A.3 Proof of Lemma 15
Lemma (Reduction from simple to multi-graphs).
Choose any constant $L$, decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, error bound $c\in(0,\frac{1}{2}]$, and any function $\delta(n_{0})\in(0,1),m(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2})$. Suppose there exists an algorithm $\operatorname{\mathcal{A}}$ such that, for any simple directed graph with $\Theta(Ln_{0})$ nodes and $\Theta(Ln_{0}+Lm_{0})$ edges, $\operatorname{\mathcal{A}}$ answers the SSPPR-R problem with threshold $\delta$, error bound $c$, and expected query complexity $T$ in arc-centric model.
Then, there exists an algorithm $\mathcal{A}_{L}$ in arc-centric model such that for any integers $n\leq n_{0},\;D,\;d,\;r,\;n(D+r)\leq m$, and any $r$-padded instance of $\mathrm{U}\in\mathcal{U}(n,D,d)$ with edge multiplicity bounded by $L$:
  1. (i).
    Expected query complexity of $\mathcal{A}_{L}$ is bounded by a function of $n_{0}$ in $\operatorname{\mathcal{O}}(T)$ , and the failure probability of $\mathcal{A}_{L}$ is also bounded by $p$;
  2. (ii).
    $\mathcal{A}_{L}$ solves the SSPPR-R problem with decay factor $1-(1-\alpha)^{3}$, error requirement $c$, threshold $\delta$ for target nodes in set $X$ on $\mathrm{U}(r)$;
  3. (iii).
    $\mathcal{A}_{L}$ executes each query ${\textsf{ADJ}}_{in/out}(v,k)$ by a uniformly random selection of an uncovered label $k$ among $v$’s in or out edges.
An analogous conclusion holds for SSPPR-A case.
Our proof begins by considering how an algorithm $\mathcal{A}$ can operate on a multi-graph $\mathrm{U}(r)$ whose edge multiplicity is bounded by $L$. Intuitively, we construct an $L$-simple lift of $\mathrm{U}(r)$ by creating $L$ parallel copies of every node and edge, thereby transforming $\mathrm{U}(r)$ into a simple graph. This lifted graph allows $\mathcal{A}$ to run as if it were operating on a simple graph, while its queries can be simulated through queries on the original multi-graph $\mathrm{U}(r)$. However, this construction presents subtle challenges: since we are restricted to local queries on $\mathrm{U}(r)$, we might not even know the number of parallel edges between two already explored nodes. To address this, we introduce independent randomness at each local structure to probabilistically simulate a consistent global lift.
Lift of graph.
Figure 3: Illustration of a $2$-Lift.
We now formalize this idea by describing the $L$-lift of a multi-graph, which encompasses the family of lifted simple graphs on which $\mathcal{A}$ will be executed, and which can be generated based on local information. Let $\mathrm{U}(r)=(V_{\mathrm{U}(r)},E_{\mathrm{U}(r)})$ be a multi-graph with $|V_{\mathrm{U}(r)}|\leq n$, $|E_{\mathrm{U}(r)}|\leq m$, and edge multiplicity $\mathrm{MUL}(V_{\mathrm{U}(r)})\leq L$, We then describe a $L$-lift of $\mathrm{U}(r),\mathcal{L}_{\rho}(\mathrm{U}(r))=(\widehat{V}_{U},\widehat{E}_{U})$ with new node set $\widehat{V}_{U}$, edge set $\widehat{E}_{U})$ and $\rho=(\rho_{(w,v)})_{(w,v)}$ is a sequence of attributes, which we will define later. Concerning $\widehat{V}_{U}$, we incorporate $2L$ extra nodes for every $v\in V_{\mathrm{U}(r)}$, : $v^{\mathrm{in}}_{1},\dots,v^{\mathrm{in}}_{L}$ , $v^{\mathrm{out}}_{1},\dots,v^{\mathrm{out}}_{L}$, which results in $\widehat{V}_{\mathrm{U}(r)}\;=\;V\;\cup\;\{v^{\mathrm{in}}_{1},\dots,v^{\mathrm{in}}_{L},v^{\mathrm{out}}_{1},\dots,v^{\mathrm{out}}_{L}\mid\forall v\in V_{\mathrm{U}(r)}\}.$ For $\widehat{E}_{U}$, We further construct it by two parts:
  • New edges. Add the edge $(v,\,v^{\mathrm{out}}_{i})$ (if any $v$’s out-degree is 00, we do not insert such edges) and $(v^{\mathrm{in}}_{i},\,v)$ (if any $v$’s in-degree is 00, we do not insert such edges) and $(v^{\mathrm{in}}_{i},\,v)$ for every $v\in V$ and $i\in\{1,\dots,L\}$.
[…]
    $\bigl{(}w^{\mathrm{out}}_{i},\;v^{\mathrm{in}}_{(i+\rho_{(w,v)}[e]-1)\bmod L}\bigr{)}.$
[…]
Construction of the algorithm.
Now that we have defined $L$-lift of a graph, then we can dive into the details of $\mathcal{A}_{L}$. Our goal is to construct an algorithm that queries on the original graph $\mathrm{U}(r)$ as if the original algorithm $\mathcal{A}$ were querying on a $L$-lift of $\mathrm{U}(r)$. Because $\mathcal{A}_{L}$ learns only from the original multi-graph $\mathrm{U}(r)$, it must maintain both a recording of all queries that it have done, and an auxiliary table about the lift parameter $\rho$. Specifically, $\mathcal{A}_{L}$ has an inner simulated algorithm $\mathcal{A}$ and internal state :$<\mathrm{U}_{0},\rho^{0}>$. We next describe the two components $\mathrm{U}_{0}$ and $\rho^{0}$.
  • $\mathrm{U}_{0}$ is the covered subgraph of $\mathrm{U}(r)$ by the queries it has executed on $\mathrm{U}(r)$; It consists of all nodes in $\mathrm{U}(r)$, all covered edges with original labels. For example, if the first query ${\textsf{Adj}}_{in}(x_{i},k)$ returns $(y_{j},k^{{}^{\prime}})$, we put an edge in $U_{0}$ between $x_{i},y_{j}$, and label it with the $k^{th}$ in-edge of $x_{i}$ as well as the $k^{\prime th}$ out edge of $y_{j}$, though there is only one edge to $x_{i}$ in $\mathrm{U}_{0}$.
  • $\rho^{0}=(\rho^{0}_{(w,v)})_{(w,v)}\in\mathrm{INJ}_{L}(\mathrm{U}_{0})$. Specifically for each pair of nodes $v,w$ in $\mathrm{U}(r)$, an injection that maps from {covered edges from w to $\{\text{covered edges from $w$ to $v$}\}$ to $\{1,\cdots,L\}$.
To simplify notation, we claim that an $\mathrm{U},\rho\in\mathrm{INJ}(\mathrm{U}(r))$ is over $<\mathrm{U}_{0},\rho^{0}>$ if and only if $\mathcal{L}_{\rho_{0}}(\mathrm{U_{0}})\subset\mathcal{L}_{\rho}(\mathrm{U}(r))$; Equivalently, this means $\mathrm{U}_{0}\subset\mathrm{U}(r)$ up to their labeling , and $\rho_{(w,v)}(e)=\rho^{0}_{(w,v)}(e)$, for arbitrary node pairs $w,v$ in $\mathrm{U}(r)$ and $e$ from $w$ to $v$ in $\mathrm{U}_{0}$. Intuitively, this means $\mathcal{A}_{L}$ knows that it is possibly executed on $\mathrm{U}(r)$, and the inner $\mathcal{A}$ knows it is possible on $\mathcal{L}_{\rho}(\mathrm{U}(r))$. With this information stored, $\mathcal{A}_{L}$ is now able to simulate $\mathcal{A}$. Specifically, when $\mathcal{A}$ tries to execute a query $Q$, $\mathcal{A}_{L}$ processes it according to the strategies outlined below:
  1. (i).
    $\mathcal{A}_{L}$ needs not to execute any query, and can gives $\mathcal{A}$ with an inferred return, because the return value of $Q$ is identical on $\mathcal{L}_{\sigma}(\mathcal{U}(r))$ for all $\mathrm{U},\rho$ over $<\mathrm{U}_{0},\rho^{0}>$. This means no matter how we choose $\mathrm{U}$ and how we modify $\rho^{0}$ later, the returning value of that query is the same, and then $\mathcal{A}_{L}$ need not do actual queries. The following queries fall into this case:
    • $Q={\textsf{INDEG}}$ or OUTDEG. This is because whatever $\mathrm{U}\in\mathrm{U}(r)$, the degree is fixed for each node, and in every $L$-lift of $\mathrm{U}(r)$, the copied nodes inherits those degrees as in definition;
    • $Q$ is an ${\textsf{ADJ}}_{in/out}$ query with an index $k$ exceeding the corresponding degree; its answer is likewise predetermined;
    • $Q={\textsf{ADJ}}_{in/out}(v,k)$. The answer is simply the node $v^{in/out}_{k}$ or ”None”;
    • $Q={\textsf{ADJ}}_{out}(w_{i}^{out},k)$ (the case $Q={\textsf{ADJ}}_{in}(w_{i}^{in},k)$ is analogous), where the label ”$k^{th}$ out-edge” of $w$ (or $k^{th}$ in-edge of $w$ ) corresponds to an edge $e$ in $\mathrm{U}_{0}$. Denote the other node of that edge $e$ is $v$ and $e$ is $v^{\prime}$s $k^{{}^{\prime}th}$ in-edge. Following the construction of simple lift, we have the target label $j=((i-1+\rho^{0}_{w,v}(e))\bmod L)+1$, then the inferred answer is $v^{in}_{j}$.
  2. (ii).
    $Q={\textsf{JUMP}}$(): just execute JUMP() on $\mathrm{U}(r)$ and return $w$. Then we choose an answer for $\mathcal{A}$ uniformly random in $\{w,w^{in}_{1},\cdots,w^{in}_{L},w^{out}_{1},w^{out}_{L}\}$.
  3. (iii).
    $Q={\textsf{ADJ}}_{out}(w_{i}^{out},k)$ or ${\textsf{ADJ}}_{in}(w_{i}^{in},k)$, where the label ”$k^{th}$ out-edge” of $w$ (or $k^{th}$ in-edge of $w$ ) has not appear in $\mathrm{U}_{0}$. In this case, we can execute query ${\textsf{ADJ}}_{out}(w,k)$ and then it turns to Case $(i)$ in the above.
Up to now, we finish the construction of $\mathcal{A}_{L}$ and then we will reveal next its properties.
Properties of the algorithm
Now it is time to verify each condition in the lemma for $\mathcal{A}_{L}$. Let us begin with the Condition $(iii)$ in the lemma. We first pre-process on $\mathcal{A}$ to make it satisfy this condition. Since $\mathcal{A}$ works on general graphs (without random or arbitrary labeling), and a uniform bound for complexity or failure probability, we can assume that whenever $\mathcal{A}$ tries to execute a ADJ query at a node $v$, it selects an uncovered in (or out) edge of $v$ uniformly at random. Switching to these random selection strategy will make $\mathcal{A}$’s complexity and failure probability to be average on each labeling case, for all possible edge labeling, while earlier works [7; 46], it’s by default that the algorithm executes this random strategy on unlabeled graphs.
Now we address Conditions $(i)$ and $(iii)$. Before proceeding, we compute the PPR values on the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$. We demonstrate that for any $\mathrm{U}(r)$, where $\mathrm{U}\in\mathcal{U}(n,D,d)$, the PPR values of the target nodes in $X$ remain unchanged under the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$. As shown in Lemma 9, when the decay factor is $1-(1-\alpha)^{3}$, the PPR value satisfies $\pi_{U(r)}(s,x_{i})=\frac{D}{n(2D-d+r)}(1-\alpha)^{6}$. For the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$, we compute the PPR values step by step. For any $y_{j}\in Y_{1}$, we have $\pi_{\mathcal{L}_{\rho}(U(r))}(s,y_{j})=\frac{1}{n}\alpha(1-\alpha)$. Since each $x_{i}\in X_{1}$ has no out-edges, we have
$\pi_{\mathcal{L}_{\rho}(\mathrm{U}(r))}(s,x_{i})=\frac{\alpha(1-\alpha)^{6}D}{n(2D-d+r)}\cdot\sum_{t=0}^{\infty}(1-\alpha)^{t}=\frac{(1-\alpha)^{6}D}{n(2D-d+r)}=\pi_{\mathrm{U}(r)}(s,x_{i}).$
An analogous argument holds for any $x_{i^{\prime}}\in X_{2}$, for which $\pi_{\mathcal{L}_{\rho}(U(r))}(s,x_{i^{\prime}})=\frac{D-d}{n(2D-d+r)}(1-\alpha)^{6}=\pi_{U(r)}(s,x_{i^{\prime}})$. Thus, Condition $(ii)$ is satisfied when running $\mathcal{A}$ on $\mathcal{L}_{\rho}(\mathrm{U}(r))$. In conclusion, for any choice of $\rho$, the PPR values on the node set $X$ remain identical between $\mathrm{U}(r)$ and its lift $\mathcal{L}_{\rho}(\mathrm{U}(r))$. This validates the construction of $\mathcal{A}_{L}$, since it simulates the behavior of $\mathcal{A}$ on some lifted version $\mathcal{L}_{\rho}(\mathrm{U}(r))$. By definition, $\mathcal{A}$ is capable of solving the SSPPR problem on any lifted graph $\mathcal{L}_{\sigma}(\mathrm{U}(r))$, for arbitrary $\sigma$, with failure probability at most $p$ and query complexity $\mathcal{O}(T)$. Therefore, if we sample $\rho\sim\mathrm{Uniform}(\mathrm{INJ}(\mathrm{U}(r)))$, and execute $\mathcal{A}$ on $\mathcal{L}_{\rho}(\mathrm{U}(r))$, we obtain expected query complexity $\mathcal{O}(T)$ and failure probability bounded by $p$, successfully solving the SSPPR problem on set $X$.
Then for Condition $(ii)$ and $(iii)$, we need to show that the query-and-answer sequence which $\mathcal{A}_{L}$ receives and sends to its inner simulated $\mathcal{A}$, has the same distribution as the distribution of query-and-answer sequence which $\mathcal{A}$ generates when executed on $\mathcal{L}_{\rho}(\mathrm{U})$ for a uniformly random sample $\rho\sim Uniform(\mathrm{INJ}_{L}(\mathrm{U}(r)))$. That is, for all $Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}$
$|\mathrm{INJ}(\mathrm{U}(r))|^{-1}\sum_{\rho}\operatorname{{\bf Pr}}[\mathcal{A}\text{ generates }Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}\text{ on}\mathcal{L}_{\rho}(\mathrm{U}(r))]$
$=\operatorname{{\bf Pr}}[\mathcal{A}_{L}\text{ generates }Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}].$
Before we use induction on $T$ to prove this result, we will first evaluate each term of
[…]
and then showing that it is almost a constant by another induction.
Now we begin this induction with the recursive formula, by showing it is almost a constant. Since $\mathcal{A}$ can only learn information through the sequence $Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}$, the distribution of next query: $P(Q_{t+1}|Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t})$ is independent of $\rho$. Also notice that the distribution of returning value is independent of $\mathcal{A}$, then we have
$\frac{\operatorname{{\bf Pr}}[\mathcal{A}\text{ generates }Q_{1},\cdots,Q_{t+1};R_{1},\cdots,R_{t+1}\text{ on }\mathcal{L}_{\rho}(\mathrm{U}(r))]}{\operatorname{{\bf Pr}}[\mathcal{A}\text{ generates }Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}\text{ on }\mathcal{L}_{\rho}(\mathrm{U}(r)]}$
$=\operatorname{{\bf Pr}}_{\mathcal{A}}[Q_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}]\operatorname{{\bf Pr}}_{\rho}[Q_{t+1}=R_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}].$
In addition, $\operatorname{{\bf Pr}}_{\rho}[Q_{t+1}=R_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}]$ is either $1$ or 00 (If the query has a deterministic answer on each $\mathcal{L}_{\rho}(\mathrm{U}(r))$), or a constant not related to $\rho$ (JUMP), then the induction is completed, and we have a deterministic function $f$ such that:
$\operatorname{{\bf Pr}}_{\rho}[Q_{t+1}=R_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}]$
=$=\begin{cases}f(Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}),\text{ event }Q_{1}=R_{1},\cdots,Q_{t}=R_{t}\text{ happens on }{L}_{\rho}(\mathrm{U})\\ 0\end{cases}$
Then we can go back to our first induction that we need to proof
|$\frac{|\{\rho,\text{ event }Q_{1}=R_{1},\cdots,Q_{t+1}=R_{t+1}\text{ happens on }\mathcal{L}_{\rho}(\mathrm{U}(r))(U)\}|}{|\{\rho,\text{ event }Q_{1}=R_{1},\cdots,Q_{t}=R_{t}\text{ happens on }\mathcal{L}_{\rho}(\mathrm{U}(r))(U)\}|}$
$=\operatorname{{\bf Pr}}[Q_{t+1}=R_{t+1}\text{ for }\mathcal{A}_{L}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}].$
When $\mathcal{A}_{L}$ can infer the result of $Q_{t+1}$(Case $(i)$which means for all possible $\rho,Q_{1}=R_{1},\cdots,Q_{t+1}=R_{t+1}\text{ happens on }\mathcal{L}_{\rho}(\mathrm{U}(r))$ they share the same answer), both sides of the equation are either $1$ or 00, and the JUMP case is similar. So the only non-trivial case is Case (iii) in the construction of $\mathcal{A}_{L}$. Without loss of generality, we assume $Q_{t+1}={\textsf{ADJ}}_{in}(v^{in}_{i},k),R_{t+1}={\textsf{ADJ}}_{out}(w^{out}_{j},k^{{}^{\prime}})$. In these cases we randomly choose a new edge’s embedding, therefore both side of the equation are
$(L-|\{\text{the already decided embedding in $Q_{1},..,Q_{t};R_{1},\cdots,R_{t}$ of edges between $v,w$}\}|)^{-1}.$
A.4 Proof of Lemma 17
Lemma.
There is a constant $\gamma$, such that. For $n$ sufficiently large, $d\leq log(n)$, $D>\frac{3d}{2}$, $T\leq\gamma nD$. Given any query and returning sequence $Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}$, and a split of $X$, $\mathrm{SP}$ we have:
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}[\mathrm{SP}\text{ is thecorrect split of }\mathrm{U}\mid Q_{1},..,Q_{T};R_{1},\cdots,R_{T}]\leq\frac{1}{n}.$
Proof.
Our goal is to bound
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\!\Bigl{[}\,\mathrm{SP}\text{ is the true split}\;\Bigm{|}\;Q_{1}=R_{1},\dots,Q_{T}=R_{T}\Bigr{]}.$
The distribution $\Sigma(n,D,d)$ is generated in two independent steps:
  1. (i).
    Choose the split.  Pick $\mathrm{SP}\in\binom{X}{n}$ uniformly at random.
  2. (ii).
    Choose all permutations. We choose all $\bigl{(}\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\,\mathrm{P}_{Y_{1}}^{n(2D-d)},\,\mathrm{P}_{Y_{2}}^{n(2D-d)},\,\mathrm{S}(Y_{a},X_{b})\bigr{)}$, where $\mathrm{S}(Y_{a},X_{b})\in\{\mathrm{S}(Y_{1},X_{1}),\mathrm{S}(Y_{1},X_{2}),\mathrm{S}(Y_{2},X_{1}),\mathrm{S}(Y_{2},X_{2})\}$, randomly and independently.
By Bayes’ rule, the posterior probability mass on the event: SP is the true $\,\mathrm{SP}\text{ is the true split}\,$, equals to
$\frac{\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}}{\displaystyle\sum_{\mathrm{SP}^{{}^{\prime}}\in\binom{X}{n}}\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}},$ (2)
where the sum runs over any $\sigma=(\mathrm{SP},\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)},\mathrm{S}(Y_{a},X_{b})),\forall a,b\in\{1,2\}$. Because all splits are a-prior equiprobable, the posterior on any particular $\mathrm{SP}$ is proportional to the likelihood of the observed transcript. Now we will give a systematic evaluation of each term in this fraction. Since queries from $s$ or to $s$ is the same on all the $\mathrm{U}(r)\in\mathcal{U}(n,D,d)$, we assume all the queries $Q_{t},R_{t}$ is only between $X$ and $Y$. We first utilize a toy case for illustration of our proof.
Warm-up.
[…]
  1. (i).
    The split $\mathrm{SP}$ determines whether $x_{i}$ is in $X_{1}$ or $X_{2}$. In our assumption, it is in $X_{1}$.
  2. (ii).
    The permutation $\mathrm{P}^{2D-d}_{i}$ decides whether the $k_{Q}^{th}$ edge of $x_{i}$ is sourced from $Y_{1}$. This occurs with probability $\frac{D}{2D-d}$.
  3. (iii).
    The permutation $\mathrm{P}_{Y_{1}}^{n(2D-d)}$ determines whether the $k_{R}^{th}$ edge of $x_{i}$ points to $X_{1}$. This occurs with probability $\frac{nD}{n(2D-d)}$.
  4. (iv).
    The permutation $\mathrm{S}(Y_{1},X_{1})$ of $nD$ decides the particular out-edge index from the reverse of the $k_{Q}^{th}$ in-edge of $x_{i}$ among $nD$ choices. The probability of finding it is exactly the $k_{R}^{th}$ out-edge of $y_{j}$ is $\frac{1}{nD}$.
  5. (v).
    Other components, such as $\mathrm{P}_{Y_{2}}^{n(2D-d)}$ and $\mathrm{P}_{i^{\prime}}^{2D-d}$ for $i^{\prime}\neq i$, are irrelevant to the query–response pair $(Q,R)$ due to the independence across different parts of the construction.
Therefore, we have $\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)\,|_{SP}}[Q=R]=\frac{D}{D-d}\times\frac{nD}{n(2D-d)}\times\frac{1}{nD}$. Bearing these steps in mind, we now derive Lemma 18 to compute the overall probability. For ease of representation, we include the following additional notations:
  • Query-returning pairs are represented by: $Q_{t}={\textsf{ADJ}}_{in}(x_{i_{t}},k_{Q_{t}}),R_{t}=(y_{j_{t}},k_{R_{t}})$, or $Q_{t}={\textsf{ADJ}}_{out}(y_{j_{t}},k_{Q_{t}}),R_{t}=(x_{i_{t}},k_{R_{t}}),1\leq t\leq T_{0}$.
  • We use $e_{t}$ to represent the edge in $\mathrm{U}(r)$ with the query-returning pair $Q_{t},R_{t}$ goes through. It is an edge from $y_{j_{t}}$ to $x_{i_{t}}$. For node sets (or single node labels) $V_{1},V_{2}$, we use $E_{t}(V_{1}\rightarrow V_{2})$ for the count of $\{1\leq t^{\prime}<t,e_{t^{\prime}}\text{ originates from }V_{1},\text{pointing to }V_{2}\}$, e.g., $E_{t}(Y_{1},x_{i_{t}})$ is the count of edges from $Y_{1}$ to $x_{i_{t}}$ which on of the query-returning pairs $(Q_{1},R_{1}),\cdots,(Q_{t-1},R_{t-1})$ has passed through.
We now formulate the calculation in the following lemma.
Lemma 18.
(Conditional probability of queries) For any $\;x_{i_{T_{0}+1}}\in X_{b},\;y_{j_{T_{0}+1}}\in Y_{a},\;a,b\in\{1,2\},\mathbf{1}_{a=b}=\begin{cases}1,a=b\\ 0,a\neq b\end{cases}$,
$\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\bigl{[}Q_{T_{0}+1}=R_{T_{0}+1}\mid Q_{1}=R_{1},\dots,Q_{T_{0}}=R_{T_{0}}\bigr{]}$
$=\frac{D-(1-\mathbf{1}_{a=b})d-E_{T_{0}+1}(Y_{a}\rightarrow{x_{T_{0}+1}})}{\left(2D-d-E_{T_{0}+1}({Y\rightarrow{x_{T_{0}+1}}})\right)\left(n(2D-d)-E_{T_{0}+1}({Y_{a}\rightarrow X}\right)}).$
We proceed by induction; our induction hypothesis is: In the distribution $\Sigma(n,D,d)|_{SP}$ conditioned on $Q_{1}=R_{1},\dots Q_{t}=R_{t}$, $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ are still independent of each other, and that $\mathrm{S}(Y_{1},X_{1}),\mathrm{S}(Y_{1},X_{2}),\mathrm{S}(Y_{2},X_{1}),\mathrm{S}(Y_{2},X_{2})$ are also independent of each other. Without loss of generality, we prove the case $x_{i_{T_{0}+1}}\in X_{1},y_{j_{T_{0}+1}}\in Y_{1}$. As in the single query:
  1. (i).
    By definition, $\mathrm{SP}$ determines whether $x_{i_{T0+1}}$ is in $X_{1}$ or $X_{2}$. Here we assume that it lies in $X_{1}$.
  2. (ii).
    $\mathrm{P}^{2D-d}_{i_{T_{0}+1}}$ decides whether the $k_{Q_{T_{0}+1}}^{th}$ edge of $x_{i_{T_{0}+1}}$ is sourced from $Y_{1}$. This occurs with probability $\frac{D-E_{T_{0}+1}({Y_{1}\rightarrow{x_{i_{T_{0}+1}}}})}{2D-d-E_{T_{0}+1}({Y\rightarrow{x_{i_{T_{0}+1}}}})}$.
  3. (iii).
    $\mathrm{P}_{Y_{1}}^{n(2D-d)}$ decides whether the $k_{R_{T_{0}+1}}^{th}$ edge of $y_{j_{T_{0}+1}}$ targets to $X_{1}$. This occurs with probability $\frac{nD-E_{T_{0}+1}{(Y_{1}\rightarrow X_{1}})}{n(2D-d)-E_{T_{0}+1}({Y_{1}\rightarrow X})}$.
  4. (iv).
    Since in the composition, the $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ terms are irrelevant to each other by induction hypothesis: For simplification of notations, we call the event for $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ suitable for $R_{t},Q_{t}$: When $x_{i_{t}}\in X_{1},y_{j_{t}}\in Y_{2}$, $Q_{t}={\textsf{ADJ}}_{in}(x_{i_{t}},k_{Q_{t}}),R_{t}=(y_{j_{t}},k_{R_{t}})$, the event holds if and only if the $k_{Q_{t}}$ in-edge of $x_{i_{t}}$, i.e. $(x_{i_{t}},k_{Q_{t}})$ originates from nodes in $Y_{2}$; and the out-edge $(y_{j_{t}},k_{R_{t}})$ points to nodes in $X_{1}$. We analogously define the same event for the other seven cases $x_{i_{t}}\in X_{2}$ or $y_{j_{t}}\in Y_{1}$ or $Q_{t}={\textsf{ADJ}}_{out}(y_{j_{t}},k_{Q_{t}}),R_{t}=(x_{i_{t}},k_{R_{t}})$. Combined with induction hypothesis, the probability of
    $\operatorname{{\bf Pr}}_{\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}}\bigl{[}\text{suitable for }\mathrm{R}_{T_{0}+1},Q_{T_{0}+1}\mid\text{suitable for }R_{t},Q_{t},t=1,\cdots,T_{0}\bigr{]}$
    $=\frac{D-E_{T_{0}+1}({Y_{1}\rightarrow{x_{i_{T_{0}+1}}}})}{2D-d-E_{T_{0}+1}({Y\rightarrow{x_{i_{T_{0}+1}}}})}\times\frac{nD-E_{T_{0}+1}({Y_{1}\rightarrow X_{1}})}{n(2D-d)-E_{T_{0}+1}({Y_{1}\rightarrow X})}.$
  5. (v).
    Conditioned on arbitrary possible $\mathrm{P}^{n(2D-d)}_{i},\mathrm{P}^{n(2D-d)}_{Y_{1}},\mathrm{P}^{n(2D-d)}_{Y_{2}}$, the bijection $\mathrm{S}(Y_{1},X_{1})$ of $nD$ elements decides the particular out-edge index from the reverse of the $k_{Q_{T_{0}+1}}^{th}$ in-edge of $x_{i_{T_{0}+1}}$ among $nD-E_{Y_{1}\rightarrow X_{1}}$ choices, The probability of exactly finding it is the $k_{R_{T_{0}+1}}^{th}$ out-edge of $y_{j_{T_{0}+1}}$ is $\frac{1}{nD-E_{T_{0}+1}({Y_{1}\rightarrow X_{1}})}$.
Consequently, the resulting probability matches the expression in Lemma 18. Since we only set some restriction independently for $\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}^{2D-d}_{i_{T_{0}+1}}$, the $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ are still independent conditioned on $Q_{t+1}=R_{t+1}$, and also for the independence of
$\mathrm{S}(Y_{1},X_{1}),\;\mathrm{S}(Y_{1},\;X_{2}),\;\mathrm{S}(Y_{2},X_{1}),\;\mathrm{S}(Y_{2},X_{2}).$
Thus the induction hypothesis remains valid for $T_{0}+1$.
Combining the analysis above, it remains to show the left hand side expression is smaller than $\frac{1}{n}$ for Lemma 17. When $T\leq\gamma nD$, there are at most $\frac{\gamma}{\gamma_{0}}n$ nodes $i$ such that $E_{T}({Y\rightarrow x_{i}})\geq\gamma_{0}D$. Then under the split $\mathrm{SP}$, there are at least $(1-\frac{\gamma}{\gamma_{0}})\cdot n$ nodes in $X_{1}$, and at least $(1-\frac{\gamma}{\gamma_{0}})\cdot n$ nodes in $X_{2}$. For each pair of $x_{i},x_{i^{{}^{\prime}}}$ such that $E_{T}({Y\rightarrow x_{i}}),E_{T}({Y\rightarrow x_{i^{{}^{\prime}}})}<\gamma_{0}D$ and in $\mathrm{SP}$ $x_{i}\in X_{1},x_{i^{{}^{\prime}}}\in X_{2}$, we have another split $\mathrm{SP}^{{}^{\prime}}$, that $\mathrm{SP}^{{}^{\prime}}$ sends $x_{i}$ to $X_{2}$, $x_{i^{{}^{\prime}}}$ to $X_{1}$, while other nodes are the same as $\mathrm{SP}$; Further, there are at least $(1-\frac{\gamma}{\gamma_{0}})^{2}n^{2}$ such pairs. Intuitively, $\mathrm{SP}^{{}^{\prime}}$ is very close to $\mathrm{SP}$, and the success probability should be very close to that of $\mathrm{SP}$ as well if we have only known insufficient data. Specifically, we calculate the proportion of their success probability as follows:
$\frac{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}}{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP^{{}^{\prime}}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}}$
$=\Pi_{t=0,\cdots,T-1}\frac{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP}}\!\bigl{[}Q_{t+1}=R_{t+1}\mid Q_{1}=R_{1},\dots,Q_{t}=R_{t}\bigr{]}}{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP^{{}^{\prime}}}}\!\bigl{[}Q_{t+1}=R_{t+1}\mid Q_{1}=R_{1},\dots,Q_{t}=R_{t}\bigr{]}}.$
Denote $Q_{t},R_{t}$ by either ${\textsf{ADJ}}_{in}(x_{i_{t}},k_{Q_{t}}),(y_{j_{t}},k_{R_{t}})$ or ${\textsf{ADJ}}_{out}(y_{j_{t}},k_{Q_{t}}),(x_{i_{t}},k_{R_{t}})\ \text{for }t=1,..,T$. By Lemma 18, we have that for any $x_{i_{t+1}}\in X_{b},y_{j_{t+1}}\in Y_{a},a,b\in\{1,2\}$,
$\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\!\bigl{[}Q_{t+1}=R_{t+1}\mid Q_{1}=R_{1},\dots,Q_{t}=R_{t}\bigr{]}$
$=\frac{D-(1-\mathbf{1}_{a=b})d-E_{t+1}(Y_{a}\rightarrow{x_{i_{t+1}}})}{\left(2D-d-E_{t+1}({Y\rightarrow{x_{i_{t+1}}}})\right)\left(n(2D-d)-E_{t+1}({Y_{a}\rightarrow X}\right)}.$
[…]
$1+\frac{d}{2D-d-E_{Y\rightarrow x_{i_{t+1}}}}\leq 1+\frac{d}{2D-d-\gamma_{0}D}\leq 1+\frac{2}{4-3\gamma_{0}}\frac{d}{D}.$
Thus, it holds that
$\frac{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}}{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}}\leq\left(1+\frac{2}{4-3\gamma_{0}}\frac{d}{D}\right)^{\gamma_{0}D}.$
Since $d\leq\log(n)$, this expression is at most $n^{\frac{2\gamma_{0}}{3-4\gamma_{0}}}$. Notice that there are at least $(n-\frac{\gamma}{\gamma_{0}}n)^{2}$ such pairs, which creates $(n-\frac{\gamma}{\gamma_{0}}n)^{2}$ terms of $\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}$ in the lower part of Equation 2, and they are at least $n^{-\frac{2\gamma_{0}}{3-4\gamma_{0}}}\cdot\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl{[}Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr{]}$. Thereby the expression is no more than $(1-\frac{\gamma}{\gamma_{0}})n^{-2+\frac{2\gamma_{0}}{3-4\gamma_{0}}}$. When $\gamma$ is a small constant and $n$ is sufficiently large, the probability is smaller than $\frac{1}{n}$. ∎