### v2

# Tighter Bounds for Personalized PageRank

Abstract
We investigate the fundamental problem of computing Personalized PageRank (PPR), which, for nodes $s$ and $t$ in a graph $G$, measures the probability $\pi(s,t)$ that an $\alpha$-decay random walk starting at $s$ terminates at $t$. We focus on two key PPR queries: (1) Single-Source PPR (SSPPR), which computes $\pi(s,\cdot)$ for a fixed source $s$; (2) Single-Target PPR (STPPR), which computes $\pi(\cdot,t)$ for a fixed target $t$. For SSPPR, two standard error regimes are considered: absolute error (SSPPR-A), requiring $|\widehat{\pi}(s,t)-\pi(s,t)|\leq\varepsilon$ for all $t$, and relative error (SSPPR-R), requiring $|\widehat{\pi}(s,t)-\pi(s,t)|\leq c\cdot\pi(s,t)$ for all $t$ with $\pi(s,t)\geq\delta$. STPPR is typically studied under the relative-error criterion, requiring $|\widehat{\pi}(s,t)-\pi(s,t)|\leq c\cdot\pi(s,t)$ for all $s$ with $\pi(s,t)\geq\delta$. These PPR queries underpin a wide range of applications, including web search, recommendation systems, graph sparsification, and graph neural networks.
While decades of research have steadily improved upper bounds—achieving the current best of $\operatorname{\mathcal{O}}\!\left(\min\!\bigl(\frac{\log(1/\varepsilon)}{\varepsilon^{2}},\frac{\sqrt{m\log n}}{\varepsilon},m\log\frac{1}{\varepsilon}\bigr)\right)$ for SSPPR-A and $\operatorname{\mathcal{O}}\!\left(\min\!\bigl(\frac{\log(1/\delta)}{\delta},\sqrt{\frac{m\log n}{\delta}},m\log\frac{\log n}{\delta m}\bigr)\right)$ for SSPPR-R— the known lower bounds remain far weaker: only $\Omega\!\bigl(\min(n,1/\varepsilon)\bigr)$ and $\Omega\!\bigl(\min(m,1/\delta)\bigr)$ for SSPPR-A and SSPPR-R, and $\Omega\!\bigl(\min(n,1/\delta)\bigr)$ for STPPR. These gaps leave the true complexity of PPR query computation largely unresolved, with no matching upper–lower characterizations.
In this paper we significantly narrow these gaps. We first present a new Monte Carlo–based algorithm that sharpens the SSPPR-A upper bound to $\operatorname{\mathcal{O}}(1/\varepsilon^{2})$, matching our new lower bound in key regimes. Building on a flexible hard-instance construction in the arc-centric model, we then prove: $\Omega\!\left(\min(m,\frac{\log(1/\delta)}{\delta})\right)$ lower bound for SSPPR-R, $\Omega\!\left(\min(m,\frac{1}{\varepsilon^{2}})\right)$ (and intermediate $\Omega\!\left(\min(m,\frac{\log(1/\varepsilon)}{\varepsilon})\right)$) for SSPPR-A, and $\Omega\!\left(\min(m,\frac{n}{\delta}\log(n))\right)$ for STPPR. These results improve prior bounds significantly and for the most relevant settings (e.g., $\delta=\Theta(1/n)$ or $\varepsilon=\Theta(n^{-1/2})$) and many real-world graphs with $m\in\Omega(n\log n)$, our lower bounds match the upper bounds. This establishes, for the first time, the practical optimality of leading algorithms such as Monte Carlo and FORA for SSPPR-R, of our new algorithm for SSPPR-A, and of the best-known method RBS for STPPR, offering a more complete complexity landscape for PPR queries.
[…]
1 Introduction
The Personalized PageRank (PPR) algorithm, first proposed in [15], has become a fundamental graph query, attracting sustained interest from both academia and industry. Given a graph $G$, the PPR value $\pi(s,t)$ with respect to a source node $s$ and a target node $t$ is defined as the probability of an $\alpha$-decay random walk***with probability $\alpha$ to stop at the current node and probability $1-\alpha$ to randomly transition to neighbors starting from $s$ and terminating at $t$. This definition naturally captures the importance of node $t$ relative to $s$, giving rise to two key query types: the Single-Source PPR (SSPPR) query, which computes $\pi(s,\cdot)$—the PPR values of all nodes with respect to a single source ss—and the Single-Target PPR (STPPR) query, which computes $\pi(\cdot,t)$—the PPR values of all nodes with respect to a single target $t$. Both query types are widely used in applications such as web search [6; 42], recommendation systems [9; 49; 33], graph sparsification [36; 51], and graph neural networks [18; 4; 40], among many others.
Exact computation of PPR vectors is expensive, requiring matrix inversion with $\Omega(n^{2}\log n)$ time [35], which is prohibitive for large real-world graphs. Consequently, considerable effort has focused on approximation algorithms for SSPPR and STPPR to balance efficiency and accuracy. Over recent decades, a series of practically effective algorithms with rigorous guarantees has emerged [11; 48; 38; 31; 50; 3; 25; 24; 27]. Building upon classical Monte Carlo (MC) and Forward Push (FP) techniques, state-of-the-art methods such as FORA [48] and SpeedPPR [39] combine these ideas to achieve the best known computational upper bounds. However, fundamental breakthroughs on upper bounds have stalled since these methods, prompting the natural question: to what extent can current upper bounds for SSPPR and STPPR be further improved?
By contrast, the theory of lower bounds for these queries remains far less developed. Constructing hard instances that force any algorithm to incur substantial computation is challenging. The best known lower bounds are $\Omega\!\left(\min(n,1/\varepsilon)\right)$ for SSPPR with absolute error [47] and $\Omega\!\left(\min(n,1/\delta)\right)$ for SSPPR with relative error, both largely derived from worst-case output-size arguments or from the simpler single-pair PPR setting. For STPPR, existing bounds remain similarly loose, with only $\Omega\!\left(\min(n,1/\delta)\right)$ known [45], which does not fully capture the complexity of computing PPR values from all sources to a single target.
In this paper, we significantly narrow these gaps. We present tighter lower bounds with substantial improvements for SSPPR-A, SSPPR-R, and STPPR, and we design a new Monte Carlo–based algorithm that sharpens the upper bound for SSPPR-A. Our results significantly narrow the gap between upper and lower bounds and, in key parameter regimes—such as $\delta=\Theta(1/n)$ for SSPPR-R and STPPR—our lower bounds even match the best known upper bounds on most real-world graphs, thereby revealing the practical optimality of leading algorithms such as MC and FORA.
The remainder of this section formally states the problem, reviews existing bounds, and outlines our motivations and contributions.
1.1 Problem Formulation
Among the comprehensive approximation measurements including the $\ell_{1}$, $\ell_{2}$, absolute, relative and degree-normalized absolute error, we focus on the research problem of SSPPR query with relative and additive error guarantee unifiedly. These two settings are widely utilized and studied in mainstreaming PPR-related researches [48; 38; 11; 45] which requires the estimator to be accurate within chosen error requirement. We give their formal definition below:
Definition 1.
Single Source Personalized PageRank Query with Additive Error Guarantee (SSPPR-A) . Given a graph $G$, a decay factor $\alpha$, a source node $s$, an additive error parameter $\varepsilon\in(0,1)$ and an overall failure probability parameter $p_{f}$, SSPPR-A query requires an estimated SSPPR vector $\widehat{\pi}(s,\cdot)$ such that with at least $1-p_{f}$ probability,
$|\widehat{\pi}(s,t)-\pi(s,t)|\leq\varepsilon$
holds for all node $t$.
Definition 2.
Single Source Personalized PageRank Query with Relative Error Guarantee (SSPPR-R). Given a graph $G$, a decay factor $\alpha$, a source node $s$, a relative error parameter $c\in(0,0.5)$, an overall failure probability parameter $p_{f}$ and choosing a targeting PPR threshold $\delta\in(0,1)$, SSPPR-R query requires an estimated SSPPR vector $\widehat{\pi}(s,\cdot)$ such that with at least $1-p_{f}$ probability,
$|\widehat{\pi}(s,t)-\pi(s,t)|\leq c\cdot\pi(s,t)$
holds for all node $t$ with its PPR value $\pi(s,t)\geq\delta$. Note that the threshold $\delta$ can be chosen arbitrary close to 0 or 1, which correspond to the shallow or large targeting PPR values.
Definition 3.
Single Target Personalized PageRank Query with Relative Error Guarantee (STPPR). Given a graph $G$, a decay factor $\alpha$, a target node $t$, a relative error parameter $c\in(0,0.5)$, an overall failure probability parameter $p_{f}$ and choosing a targeting PPR threshold $\delta\in(0,1)$, STPPR query requires an estimated STPPR vector $\widehat{\pi}(\cdot,t)$ such that with at least $1-p_{f}$ probability,
$|\widehat{\pi}(s,t)-\pi(s,t)|\leq c\cdot\pi(s,t)$
holds for all node $s$ with its PPR value $\pi(s,t)\geq\delta$. Note that the threshold $\delta$ can be chosen arbitrary close to 0 or 1, which correspond to the shallow or large targeting PPR values.
Following conventions in prior work [17; 19; 46; 23], we assume that the parameters $\alpha$, $c$, and $p_{f}$ are treated as constants by default, unless otherwise specified.
Table 1: Best-Known results and our Improvements of SSPPR-A, SSPPR-R and STPPR query. We highlight our new results in red and bold the Improvements when $\varepsilon,\delta$ are sufficiently small. For ease of simplicity, we here assume the number of edges, $m\in\operatorname{\mathcal{O}}(n^{2-\beta})$ for any given arbitrary small constant $\beta\in(0,1)$ in our Tighter SSPPR-A Lower Bounds.
Best-Known Lower Bounds Tighter Lower BoundsA concurrent work [5], developed independently with a different technique, establishes lower bounds $\Omega(\min(m,\tfrac{1}{\delta}))$ for single-node PPR and $\Omega(\min(m,\tfrac{n}{\delta}))$ for STPPR. Our lower bounds are stronger, providing additional $\log(1/\delta)$ and $\log n$ factors, respectively, and are proved solely under the standard arc-centric graph query model, whereas [5] relies on stronger query primitives such as IN-SORTED, which returns sorted in-neighbors. Improvements
SSPPR-A $\Omega\left(\min(n,\frac{1}{\varepsilon})\right)$ [47] $\Omega\left(\min(m,\frac{1}{\varepsilon^{2}})\right)$ $\frac{1}{\varepsilon}$ or $\frac{m}{n}$}
SSPPR-R $\Omega\left(\min(n,\frac{1}{\delta})\right)$ [TRIVIAL] $\Omega\left(\min(m,\frac{\log(\frac{1}{\delta})}{\delta})\right)$ $\log(\frac{1}{\delta})$ or $\frac{m}{n}$}
STPPR $\Omega\left(\min(n,\frac{1}{\delta})\right)$ [45] $\Omega\left(\min(m,\frac{n}{\delta}\log(n))\right)$ $n\log(n)$ or $\frac{m}{n}$
Best-Known (Improved) Upper Bounds
STPPR $\operatorname{\mathcal{O}}\left(\min(m\log(\frac{1}{\delta}),\frac{n}{\delta}\log(n))\right)$
[…]
1.2 Existing Results
From a theoretical perspective, we summarize the best-known upper and lower bounds for SSPPR-A, SSPPR-R and STPPR queries. An overview of these results is presented in Table 1.
1.3 Upper Bounds
In this section, we briefly summarize the best-known computational upper bounds for answering SSPPR-A, SSPPR-R and STPPR queries. These bounds are primarily derived from three foundational algorithmic paradigms—Monte Carlo (MC), Forward Push (FP), and Backward Push (BP)—and have been further refined through various state-of-the-art (SOTA) techniques. A comprehensive review of these algorithms, along with their complexity analyses, is provided in Appendix A.1.
Basic Techniques.
The MC method provides a sample-based approximation to the PPR vector using $\alpha$-decay random walks, and achieves high-probability relative error guarantees with expected time complexity $\operatorname{\mathcal{O}}\left(\log(n)/\delta\right)$ for SSPPR-R, or $\operatorname{\mathcal{O}}\left(\log(1/\varepsilon)/\varepsilon^{2}\right)$ for SSPPR-A. The FP method is a local-push algorithm that simulates random walks deterministically via graph traversal. To guarantee an additive error of $\varepsilon$, it requires setting a residue threshold $r_{\max}=\varepsilon/m$, resulting in a worst-case time complexity of $\operatorname{\mathcal{O}}(m/\varepsilon)$. The BP method is similarly structured but performs the push operation in reverse, tailed to STPPR. With global BP, $\operatorname{\mathcal{O}}(m\log(1/\delta))$ is required for solving STPPR. BP can also be applied naively to SSPPR-A for all target nodes with $\operatorname{\mathcal{O}}(m/\varepsilon)$ complexity.
State-of-the-Art Methods.
To overcome the inefficiencies of basic PPR algorithms, hybrid methods that combine Monte Carlo (MC) sampling with local-push strategies have been developed. FORA [48] improves upon MC by first performing deterministic Forward Push (FP) and then refining the result via targeted MC walks. By optimally balancing these two phases, it achieves an expected complexity of $\operatorname{\mathcal{O}}\!\bigl(\sqrt{(m\log n)/\delta}\bigr)$ for SSPPR-R. SpeedPPR [38] restructures FP to enforce a more regular push schedule, reducing total residual mass more efficiently and yielding $\operatorname{\mathcal{O}}\!\bigl(m\log(1/\varepsilon)\bigr)$ for SSPPR-A and comparable efficiency for SSPPR-R when combined with MC. More recently, ABP [47] integrates MC with an adaptive variant of Backward Push (BP): it first constructs a candidate set through rough MC estimation and then uses selective backward pushes and extra walks for refinement, attaining sublinear expected complexity $\operatorname{\mathcal{O}}\!\bigl(\sqrt{m\log n}/\varepsilon\bigr)$ for SSPPR-A. For the single-target case, RBS [45] injects randomness into BP’s deterministic push, maintaining an unbiased estimator with bounded variance while skipping negligible probability mass, thus answering STPPR in $\operatorname{\mathcal{O}}\!\bigl(n/\delta\log(1/\delta)\bigr)$ time. These methods illustrate how combining sampling with localized deterministic updates leads to near-optimal performance and greatly improves the scalability of PPR computation on large-scale graphs.
1.4 Lower Bounds
Unlike the extensive body of work on improving upper bounds for SSPPR computation, rigorous lower bounds remain sparse and often heuristic, offering limited insight into the fundamental complexity of the problem. The earliest bound relevant to Personalized PageRank (PPR) is due to FastPPR [21], which gives a lower bound of $\Omega(\sqrt{1/\delta})$ for single-pair PPR with relative error (SSPPR-R). However, this is dominated by the trivial output-size bound $\Omega\!\bigl(\min\{n,1/\delta\}\bigr)$ obtained by noting that there can be up to $\min\{n,1/\delta\}$ targets $t$ with $\pi(s,t)\!\geq\!\delta$. To the best of our knowledge, no other nontrivial direct lower bounds are known for SSPPR-R. For the absolute-error variant (SSPPR-A), [47] prove $\Omega\!\bigl(\min\{n,1/\varepsilon\}\bigr)$ by the same output-size reasoning: any algorithm must return nonzero estimates for all $t$ with $\pi(s,t)\!\geq\!\varepsilon$, of which there may be $\Omega\!\bigl(\min\{n,1/\varepsilon\}\bigr)$. While conceptually simple, such arguments understate the difficulty of producing accurate estimates for all entries within prescribed error margins. Likewise, for the single-target PPR (STPPR) problem, [45] show lower bounds of $\Omega(n)$ in the worst case and $\Omega\!\bigl(\min\{n,1/\delta\}\bigr)$ on average, which still leave a substantial gap relative to known upper bounds. In contrast, there has been notable progress on a related but distinct task: estimating nodal PageRank (PR). The BiPPR [19] algorithm attains optimal complexity on directed graphs shown in [46], and a Monte Carlo variant, BACKMC, achieves matching upper and lower bounds on undirected graphs [37]. The proof techniques in these works are simple yet elegant, and they motivate the pursuit of tighter lower bounds for SSPPR. Building on these insights, we construct a hard instance tailored specifically to single-source queries. Its structure and analysis differ substantially from those in [46; 37], requiring more delicate complexity arguments and leading to new lower-bound results for SSPPR.
1.5 Our Results and Discussion
In this section, we present our main results on improved upper and lower bounds for PPR query computation. All our arguments are developed under the standard arc-centric model [12; 14], which will be formally defined in Section 3. We begin with an improved computational upper bound for the SSPPR-A problem, achieved via a new algorithm that refines the classical Monte Carlo (MC) approach.
Theorem 4 (New algorithm and improved upper bound of SSPPR-A).
We present Algorithm 1, which solves the SSPPR-A (or SSPPR-R) problem within $\operatorname{\mathcal{O}}(1/\varepsilon^{2})$ (or $\operatorname{\mathcal{O}}(\log(1/\delta)/\delta)$) queries in expectation under the standard arc-centric model.
Algorithm 1 introduces a refined two-phase estimator compared to naive MC. In the first (discovery) phase, it performs a collection of random walks to identify a candidate set of nodes whose PPR is likely to exceed the target threshold. In the second (estimation) phase, it focuses computation only on this candidate set and estimates their PPR values with higher precision. This design substantially strengthens the classical MC upper bound, achieving, for example, an additional $\log(1/\varepsilon)$ factor of savings for SSPPR-A. To further narrow the remaining gap between upper and lower bounds, we next establish significantly stronger lower bounds.
Theorem 5 (Lower bound of SSPPR-R query).
Choose any decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, error parameter $c\in(0,\frac{1}{2}]$, and any functions $\delta(n)\in(0,1),m(n)\in\Omega(n)\cap\operatorname{\mathcal{O}}(n^{2})$. Consider any oracle (possibly randomized) algorithm within the arc-centric model that, for an arbitrary graph $G$ with $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes and any source node $s$, with probability at least 1p1-p gives an estimation $\widehat{\pi}(s,\cdot)$ for SSPPR vector $\pi(s,\cdot)$, satisfying $|\widehat{\pi}(s,v)-\pi(s,v)|\leq c\cdot\pi(s,v)$ for each node $v$ at $\pi(s,v)\geq\delta$. Then there exists a graph $\mathrm{U}$ and a source node $s$ such that:
  1. (i).
    $\mathrm{U}$ contains $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes;
  2. (ii).
    the algorithm requires $\Omega\left(\min(m,\frac{\log(1/\delta)}{\delta})\right)$ queries to succeed in expectation.
Theorem 5 represents a major advance by improving the best-known trivial lower bound from $\Omega\!\bigl(\min(n,1/\delta)\bigr)$ to $\Omega\!\bigl(\min(m,\log(1/\delta)/\delta)\bigr)$. This improvement is substantial, adding a multiplicative factor of up to $m/n$, which can reach $\Theta(n)$ on dense graphs such as cliques when high approximation precision is required. It further reveals a key insight: when the graph is at least sparse-free ($m=\Omega(n\log n)$) and the error threshold satisfies $\delta=1/n$—the standard setting in PPR query computation [48; 39; 23; 19]—our lower bound matches the known computational upper bounds of MC and FORA. This indicates that MC and FORA are in fact optimal algorithms for computing SSPPR-R queries in practical regimes.
Moreover, the entire proof framework (outlined in Section 5) is readily adaptable. By appropriately reconfiguring the graph parameters in our hard-instance construction or by inverting all edge directions in the SSPPR-R argument, we derive Theorems 6 and 7, which establish even stronger lower bounds for the SSPPR-A and STPPR problems, respectively.
Theorem 6 (Initial lower bound of SSPPR-A query).
Choose any decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, error parameter $c\in(0,\frac{1}{2}]$, and any functions $\varepsilon(n)\in(0,1),m(n)\in\Omega(n)\cap\operatorname{\mathcal{O}}(n^{2-\beta})$, where $\beta\in(0,1)$ is an arbitrary small constant. Consider any oracle (possibly randomized) algorithm within the arc-centric model that, for an arbitrary graph $G$ with $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes and any source node $s$, with probability at least 1p1-p gives an estimation $\widehat{\pi}(s,\cdot)$ for SSPPR vector $\pi(s,\cdot)$, satisfying $|\widehat{\pi}(s,v)-\pi(s,v)|\leq\varepsilon$ for each node $v$. Then there exists a graph $\mathrm{U}$ and a source node $s$ such that:
  1. (i).
    $\mathrm{U}$ contains $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes;
  2. (ii).
    the algorithm requires $\Omega\left(\min(m,\frac{\log(1/\varepsilon)}{\varepsilon})\right)$ queries to succeed in expectation.
Theorem 7 (Lower bound of STPPR query).
Choose any decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, error parameter $c\in(0,\frac{1}{2}]$, and any functions $\delta(n)\in(0,1),m(n)\in\Omega(n)\cap\operatorname{\mathcal{O}}(n^{2})$. Consider any oracle (possibly randomized) algorithm within the arc-centric model that, for an arbitrary graph $G$ with $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes and any target node $t$, with probability at least 1p1-p gives an estimation $\widehat{\pi}(\cdot,t)$ for STPPR vector $\pi(\cdot,t)$, satisfying $|\widehat{\pi}(v,t)-\pi(v,t)|\leq c\cdot\pi(v,t)$ for each node $v$ at $\pi(v,t)\geq\delta$. Then there exists a graph $\mathrm{U}$ and a target node $t$ such that:
  1. (i).
    $\mathrm{U}$ contains $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes;
  2. (ii).
    the algorithm requires $\Omega\left(\min(m,\frac{n}{\delta}\log(n))\right)$ queries to succeed in the worst case.
Benefiting from our flexible proof framework, Theorem 7 significantly strengthens the best-known worst-case lower bound from $\Omega\!\bigl(\min(n,1/\delta)\bigr)$ to $\Omega\!\bigl(\min(m,n/\delta\cdot\log(n))\bigr)$. Crucially, this lower bound matches the known computational upper bound in the prevalent setting $\delta=O(1/n)$, which is standard in PPR query computation [48; 39; 23; 19]. This match implies that the computational complexity of the STPPR problem is essentially optimal, offering strong evidence of a maturity stage in PPR algorithm design. Moreover, the improved lower bound in Theorem 6 further narrows the remaining gap for SSPPR-A. We also observe that with refined hard-instance constructions, our framework can yield even tighter bounds, as developed in the subsequent theorem.
Theorem 8 (Lower bound of SSPPR-A query).
Choose any decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, and any functions $\varepsilon(n)\in(0,1),m(n)\in\Omega(n)\cap\operatorname{\mathcal{O}}(n^{2-\beta})$, where $\beta\in(0,1)$ is an arbitrary small constant. Consider any oracle (possibly randomized) algorithm within the arc-centric model that, for an arbitrary graph $G$ with $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes and any source node $s$, with probability at least 1p1-p gives an estimation $\widehat{\pi}(s,\cdot)$ for SSPPR vector $\pi(s,\cdot)$, satisfying $|\widehat{\pi}(s,v)-\pi(s,v)|\leq\varepsilon$ for each node $v$. Then there exists a graph $\mathrm{U}$ and a source node $s$ such that:
  1. (i).
    $\mathrm{U}$ contains $\operatorname{\mathcal{O}}(m)$ edges and $\operatorname{\mathcal{O}}(n)$ nodes;
  2. (ii).
    the algorithm requires $\Omega\left(\min(m,\frac{1}{\varepsilon^{2}})\right)$ queries to succeed in expectation.
Theorem 8 strengthens the best-known lower bound from $\Omega\!\left(\min(n,1/\varepsilon)\right)$ to $\Omega\!\left(\min(m,1/\varepsilon^{2})\right)$. This improvement is substantial and, strikingly, the $\tfrac{1}{\varepsilon^{2}}$ term partially matches our improved computational upper bound for the SSPPR-A query. This alignment indicates that our new algorithm (Algorithm 1) is essentially optimal for SSPPR-A when the error requirement is relatively coarse, i.e., $\varepsilon=\Omega(n^{-1/2})$.
Remark.
As summarized in Table 1, we substantially narrow the gap between known upper and lower bounds for PPR query computation. In particular, we obtain matching upper and lower bounds in most practical regimes—when $\delta=\Theta(n^{-1})$ for SSPPR-R and STPPR, or when $\varepsilon=\Theta(n^{-1/2})$ for SSPPR-A. These results strongly suggest the optimality of our bounds and provide valuable guidance for future research on the complexity of PPR computations.
2 Other Related Works
Over the past decades, Personalized PageRank (PPR) computation has been extensively studied, with most research focusing on accelerating and scaling single-source [11; 48; 38; 31; 50; 3; 25; 24; 27], single-target [22; 20; 43; 45; 44], and top-$k$ [10; 20; 21; 48; 41] queries on general graphs. Early methods such as [3; 11; 27; 10] employed large-scale Monte Carlo (MC) simulations to estimate PPR values, whereas [22; 2] applied deterministic graph traversal to compute exact vectors. Subsequent algorithms [48; 38; 31; 41; 20; 43] improved scalability by combining deterministic traversal and MC in separate phases. More recent approaches [45; 44; 13] integrate deterministic traversal and MC in a unified Randomized Push (RS) framework; while RS can answer SSPPR-R queries, it requires $\operatorname{\mathcal{O}}(\log^{3}n/\delta)$ time [23], worse than basic MC, and is therefore omitted in many practical comparisons. Other lines of work study PPR in dynamic graphs [34; 32; 16; 30; 52] and its coupling with graph neural networks [28; 29; 26].
Despite this rich algorithmic progress, the theoretical landscape—in particular, tight upper and lower bounds for single-source absolute (SSPPR-A), single-source relative (SSPPR-R), and single-target (STPPR) queries—remains underdeveloped. The main obstacle is the lack of hard instances specifically tailored to single-source computation, which makes proving lower bounds much more challenging. Our work addresses this gap by providing new, significantly tighter lower bounds and a novel proof framework, thereby deepening the theoretical understanding of PPR query complexity and offering fresh directions for future research.
3 Notations and Tools
Notations.
Given an underlying directed graph $G=(V_{G},E_{G})$, let $n=|V_{G}|$ and $m=|E_{G}|$ denote the number of nodes and edges, respectively. Each edge $e=(u,v)\in E_{G}$ indicates that node $u$ is a parent of node $v$, and correspondingly, $v$ is a child of $u$. For a node set $V$, we denote its elements either as $v\in V$ or using index notation $V[i]$ to refer to the node at index $i$. The out-degree and in-degree of a node $u$ are denoted by $d_{\text{out}}(u)$ and $d_{\text{in}}(u)$, respectively. In particular, we denote by $\deg_{V}(v)$ the number of edges between node $v$ and all potential nodes in the set $V$. A permutation of size $n$ is defined as a bijection $\mathrm{p}^{n}\in\mathrm{S}^{n}\mathrel{\mathop{\ordinarycolon}}\{1,2,\ldots,n\}\to\{1,2,\ldots,n\}$; equivalently, $\mathrm{p^{n}}$ is an element of the symmetric group $\mathrm{S}^{n}$. For any index $0\leq i<n$, we use $\mathrm{p}[i]$ to represent the permuted index. We consider random walks on $G$ where, at each step, the walk proceeds to a uniformly random child of the current node. Additionally, we study $\alpha$-decay random walks, where at each step, the walk continues with probability $\alpha\in(0,1)$. Unless otherwise specified, we assume $\alpha$ is a constant. Finally, we allow the graph to contain parallel edges (i.e., multiple edges between the same pair of nodes), which are formally defined in Definition 9 below.
Definition 9 (Multi-graph and edge multiplicity).
A multi-graph is a graph that may contain parallel edges (multiple edges between the same pair of nodes); a simple graph contains no such parallel edges. For a multi-graph $G=(V,E)$, the edge multiplicity is defined as
$\mathrm{MUL}(G)\;=\;\max_{u,v\in V}\bigl|\{\,e\in E\mid e=(u,v)\}\bigr|,$
representing the maximum number of parallel edges between any node pair $u,v\in V$.
Definition 10 (Maximum posterior estimation).
Let $\Theta$ be a prior distribution over a finite set of possible values, and let $\mathbf{x}$ denote the observed data. Then, there exists a maximum a posteriori (MAP) estimate $\widehat{\theta}_{\mathrm{MAP}}\in\Theta$ such that the success probability of any algorithm that predicts $\theta\in\Theta$ based on $\mathbf{x}$ is at most $\operatorname{{\bf Pr}}[\widehat{\theta}_{\mathrm{MAP}}\mid\mathbf{x}]$. Furthermore, replacing $\mathbf{x}$ with any function of $\mathbf{x}$ (i.e., using only partial or derived information) cannot increase the prediction success probability.
Arc–Centric Graph Query Model [12; 14] and Query Complexity.
For our results, we focus on the computational complexity under the standard RAM model. To formalize graph interactions, we allow algorithms to access the underlying graph $G$ through a graph oracle. Specifically, we adopt the standard arc-centric graph-access model [12; 14], where the oracle supports the following queries in unit time:
  • JUMP(): returns a uniformly random node $v\in V$ and marks $v$ as covered;
  • INDEG(v)(v)/OUTDEG(v)(v): for any covered node $v\in V$, returns its in-degree $d_{in}(v)$ or out-degree $d_{out}(v)$, respectively;
  • ADJout(v,i)(v,i)/ADJin(v,i)(v,i): for any covered node $v\in V$ and integer $1\leq i\leq d_{out}(v)/d_{in}(v)$, returns the $i^{th}$ out-/in-neighbor of $v$.
This model has been widely used in prior work [21; 19; 46] to establish lower bound complexities. In our analysis, we strengthen the adjacency query ADJout(v,i)(v,i) (analogously for ADJin) to not only return the out-neighbor node $w$ of $v$ but also provide the label of the in-edge $(v,w)$ for node $w$ denoted as $k$, such that edge $(v,w)$ is the $i$-th out edge for $v$ and $k$-th in-edge for $w$. Note that this stronger model reveals more information than the original, ensuring that any lower bounds we derive remain valid for the standard arc-centric model. For simplicity and clarity in subsequent sections, we adopt this enhanced model by default. In the context of graph algorithms, query complexity refers to the number of queries the algorithm invokes to the graph oracle. Hence, query complexity serves as a lower bound to computational complexity.
Definition 11 (Graph with edge and node indices).
Following the notation in [12], we define a graph with edge and node indices as a graph $G=(V,E)$ together with the following labeling:
  • Node labels. Each node $v\in V$ is assigned a unique label from the set $\{1,\ldots,|V|\}$;
  • Edge labels. For each node $v\in V$, its in-edges and out-edges are distinctly labeled with integers from $\{1,2,\ldots,d_{\mathrm{in}}(v)\}$ and $\{1,2,\ldots,d_{\mathrm{out}}(v)\}$, respectively.
4 Proofs for Improved Upper Bounds
In this section, we revisit the Monte Carlo (MC) method for SSPPR computation and propose a new two-phase estimator that achieves the upper bounds required for Theorem 4. Throughout, we measure query complexity in the standard arc-centric model, and note that an $\alpha$-decay random walk has constant expected length, so the number of queries is proportional to the number of walks.
4.1 Concentration Inequalities
We use the following standard Chernoff bounds.
Theorem 12 (Chernoff Bound).
Let $X_{1},\ldots,X_{n}$ be independent random variables taking values in $[0,1]$. Let $X\mathrel{\mathop{\ordinarycolon}}=\sum_{i=1}^{n}X_{i}$ and $\mu\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[X]$. Then, for any $c>0$,
  1. 1.
    if $c\in[0,1]$, then $\operatorname{{\bf Pr}}[X>(1+c)\mu]\leq\exp\!\big(-\tfrac{c^{2}}{3}\mu\big)$;
  2. 2.
    if $c\in[0,1]$, then $\operatorname{{\bf Pr}}[X<(1-c)\mu]\leq\exp\!\big(-\tfrac{c^{2}}{2}\mu\big)$;
  3. 3.
    if $c\geq 1$, then $\operatorname{{\bf Pr}}[X>(1+c)\mu]\leq\exp\!\big(-\tfrac{c}{3}\mu\big)$.
4.2 Revisiting the Monte Carlo Estimator
The classical MC estimator [11] enforces a uniform accuracy guarantee over all $n$ nodes, which introduces an unavoidable $\log n$ factor via a union bound. This is well matched to regimes with extreme precision (e.g., $\delta\approx 1/n$) but overly conservative when the target precision is moderate. Our lower bound $\Omega\!\big(\min\{m,\log(1/\delta)/\delta\}\big)$ for SSPPR-R highlights the correct dependence on the threshold $\delta$. Motivated by this, we decouple discovery from estimation:
  • Discovery phase. Use a modest number of walks to find a candidate set $R_{\mathrm{out}}$ that contains all nodes of interest $V_{\geq\delta}\mathrel{\mathop{\ordinarycolon}}=\{v\mathrel{\mathop{\ordinarycolon}}\pi(s,v)\geq\delta\}$, while excluding nodes much smaller than $\delta$ by a constant factor.
  • Estimation phase. Restrict high-precision estimation to $R_{\mathrm{out}}$; this replaces the global $\log n$ factor by a $\log\!\bigl(\tfrac{1}{\delta}\bigr)$ dependence governed by the candidate set size.
We now specify the discovery phase precisely. The algorithm performs two filtering rounds.
Input: graph $G=(V,E)$, source $s\in V$, decay factor $\alpha\in(0,1)$, threshold $\delta\in(0,\tfrac{1}{2}]$, failure probability $p\in(0,1)$.
Output: $R_{\text{out}}\subseteq V$.
$T_{1}\leftarrow\lceil\tfrac{1}{\delta}\rceil\left\lceil\!\left(\ln\tfrac{1}{\delta}+\ln\tfrac{2}{p}\right)\right\rceil$; $T_{2}\leftarrow 20\lceil\tfrac{1}{\delta}\rceil\left\lceil\!\left(\ln\tfrac{1}{\delta}+\ln\tfrac{2}{p}\right)\right\rceil$;
$R^{\text{raw}}_{\text{out}}\leftarrow\varnothing$; $H[v]\leftarrow 0$ for all $v\in V$;
for $i\leftarrow 1$ to $T_{1}$ do
 $v\leftarrow$ endpoint of an $\alpha$-decay random walk from $s$;
 $R^{\text{raw}}_{\text{out}}\leftarrow R^{\text{raw}}_{\text{out}}\cup\{v\}$;
end for
for $i\leftarrow 1$ to $T_{2}$ do
 $v\leftarrow$ endpoint of an $\alpha$-decay random walk from $s$;
 $H[v]\leftarrow H[v]+1$;
end for
$\theta\leftarrow 10\left\lceil\ln\tfrac{1}{\delta}+\ln\tfrac{2}{p}\right\rceil$;
$R_{\text{out}}\leftarrow\{v\in R^{\text{raw}}_{\text{out}}\mid H[v]\geq\theta\}$;
return $R_{\text{out}}$
Algorithm 1 Two-Round Discovery
Guarantees for discovery.
The goal is to output a subset $R_{\text{out}}$ such that
  1. (i)
    No false negatives: $V_{\geq\delta}\subseteq R_{\text{out}}$.
  2. (ii)
    No heavy false positives: $R_{\text{out}}\subseteq V_{\geq\delta/(5e)}$.
Round 1 runs a pool of $\alpha$-decay walks and collects all endpoints, yielding $R^{\rm raw}_{\text{out}}$. Round 2 runs additional walks and keeps only nodes with endpoint counts above a threshold. For convenience, recall $V_{\geq\tau}\mathrel{\mathop{\ordinarycolon}}=\{v\in V\mathrel{\mathop{\ordinarycolon}}\pi(s,v)\geq\tau\}$. Since $\sum_{v}\pi(s,v)=1$, we always have $|V_{\geq\tau}|\leq 1/\tau$.
Lemma 13 (Discovering nodes of interest).
Algorithm 1 uses $\operatorname{\mathcal{O}}(\tfrac{\log(1/\delta)}{\delta})$ queries in the arc-centric model and, with probability at least 1p1-p, returns $R_{\mathrm{out}}$ satisfying $V_{\geq\delta}\subseteq R_{\mathrm{out}}\subseteq V_{\geq\delta/(5e)}$.
Proof.
Let $L\mathrel{\mathop{\ordinarycolon}}=\left\lceil\ln\tfrac{1}{\delta}+\ln\tfrac{2}{p}\right\rceil$, and set $T_{1}=L\lceil 1/\delta\rceil$, $T_{2}=20L\lceil 1/\delta\rceil$, and $\theta=10L$. We bound three failure events and allocate their budgets as $p/2$, $p/4$, and $p/4$.
(i) No false negatives in Round 1. Fix $u\in V_{\geq\delta}$. In $T_{1}$ independent walks, the miss probability is
$\operatorname{{\bf Pr}}[u\notin R^{\rm raw}_{\text{out}}]=(1-\pi(s,u))^{T_{1}}\leq(1-\delta)^{T_{1}}\leq e^{-\delta T_{1}}\leq e^{-L}=\delta\tfrac{p}{2}.$
A union bound over at most $|V_{\geq\delta}|\leq 1/\delta$ nodes gives $\operatorname{{\bf Pr}}[V_{\geq\delta}\not\subseteq R^{\rm raw}_{\text{out}}]\leq\tfrac{1}{\delta}\cdot\delta\tfrac{p}{2}=\tfrac{p}{2}$.
(ii) No false negatives in Round 2. For $u\in V_{\geq\delta}$, the count $H[u]$ from $T_{2}$ walks has mean $\mu_{u}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[H[u]]=T_{2}\pi(s,u)\geq 20L$. We threshold at $\theta=10L$, i.e., at most $\mu_{u}/2$. By Chernoff ($c=1/2$):
$\operatorname{{\bf Pr}}[H[u]<\theta]\leq\exp(-\mu_{u}/8)\leq e^{-20L/8}=e^{-2.5L}=\left(\delta\tfrac{p}{2}\right)^{2.5}.$
A union bound over $V_{\geq\delta}$ yields $\operatorname{{\bf Pr}}[\exists u\in V_{\geq\delta}\mathrel{\mathop{\ordinarycolon}}H[u]<\theta]\leq\tfrac{1}{\delta}\cdot\left(\delta\tfrac{p}{2}\right)^{2.5}=\delta^{1.5}\!\left(\tfrac{p}{2}\right)^{2.5}<p/4$ for $\delta\leq 1/2$ and sufficiently small $p$.
(iii) No heavy false positives in Round 2. Fix $u$ with $\pi(s,u)<\delta/(5e)$. Then
$\mu_{u}=T_{2}\pi(s,u)<20L\Big\lceil\tfrac{1}{\delta}\Big\rceil\cdot\tfrac{\delta}{5e}=\tfrac{4L}{e}\cdot\delta\Big\lceil\tfrac{1}{\delta}\Big\rceil<\tfrac{4L}{e}(1+\delta)\leq\tfrac{6L}{e}\qquad(\delta\leq\tfrac{1}{2}).$
With $\theta=10L$, we have $c\mathrel{\mathop{\ordinarycolon}}=\theta/\mu_{u}-1>\tfrac{10L}{6L/e}-1=\tfrac{5e}{3}-1>3.5$. Using Chernoff ($c\geq 1$),
$\operatorname{{\bf Pr}}[H[u]\geq\theta]\leq\exp\!\big(-\tfrac{\theta-\mu_{u}}{3}\big)\leq\exp\!\big(-\tfrac{L(10-6/e)}{3}\big)<e^{-2.5L}=\left(\delta\tfrac{p}{2}\right)^{2.5}.$
Round 1 produces at most $T_{1}$ raw candidates (one endpoint per walk), so by a union bound
$\operatorname{{\bf Pr}}\big[\exists u\notin V_{\geq\delta/(5e)}\mathrel{\mathop{\ordinarycolon}}\ u\in R_{\mathrm{out}}\big]\leq T_{1}\cdot e^{-2.5L}\leq L\Big(\tfrac{1}{\delta}+1\Big)\!\left(\delta\tfrac{p}{2}\right)^{2.5}<\tfrac{p}{4}$
for $\delta\leq 1/2$ and sufficiently small $p$. Summing the three failure budgets yields a total failure probability at most $p$. ∎
Estimation phase: relative error.
Having $V_{\geq\delta}\subseteq R_{\mathrm{out}}\subseteq V_{\geq\delta/(5e)}$ with probability $\geq 1-p/2$, we estimate only on $R_{\mathrm{out}}$.
Lemma 14.
The single-source personalized PageRank problem with a relative-error guarantee can be solved within expected $\operatorname{\mathcal{O}}\!\big(\tfrac{\log(1/\delta)}{\delta}\big)$ queries in the standard arc-centric query model.
Proof.
Run
$T\;=\;256\,\frac{\log\!\big(\tfrac{1}{\delta/(5e)}\big)+\log\!\big(\tfrac{2}{p}\big)}{c^{2}\,(\delta/(5e))}\;=\;\operatorname{\mathcal{O}}\!\Big(\tfrac{\log(1/\delta)}{\delta}\Big)$
independent $\alpha$-decay walks from $s$, and for each $u\in R_{\mathrm{out}}$ estimate $\pi(s,u)$ by empirical frequency. By Chernoff and a union bound over $|R_{\mathrm{out}}|\leq|V_{\geq\delta/(5e)}|\leq 5e/\delta$, the estimation failure probability is at most $p/2$. Combining with Lemma 13 yields overall failure probability $\leq p$. ∎
Estimation phase: absolute error.
For the absolute-error regime we similarly restrict estimation to nodes whose mass is $\Omega(\varepsilon)$, which allows Chernoff to be applied with an absolute tolerance.
Lemma 15.
The single-source personalized PageRank problem with an absolute-error guarantee can be solved within expected $\operatorname{\mathcal{O}}(1/\varepsilon^{2})$ queries in the standard arc-centric query model.
Proof.
By Lemma 13 (with $\delta$ replaced by $\varepsilon$), with probability $\geq 1-p/2$ we obtain $V_{\geq\varepsilon}\subset R_{\mathrm{out}}\subset V_{\geq\varepsilon/(5e)}.$ Execute
$T\;=\;1024\,\frac{\log\!\big(\tfrac{2}{p}\big)}{\big(\tfrac{\varepsilon}{5e}\big)^{2}}\;=\;\operatorname{\mathcal{O}}\!\big(\tfrac{1}{\varepsilon^{2}}\big)$
independent $\alpha$-decay walks from $s$, and estimate $\pi(s,u)$ for $u\in R_{\mathrm{out}}$ by the empirical frequency with absolute tolerance $\varepsilon/(5e)$. Because $\pi(s,u)\geq\varepsilon/(5e)$ for all $u\in R_{\mathrm{out}}$, the tolerance is at most the mean, and Chernoff yields
$\operatorname{{\bf Pr}}[\text{fail at }u]\;\leq\;2\exp\!\Big(-\tfrac{1}{3}\,\pi(s,u)\Big(\tfrac{\varepsilon/(5e)}{\pi(s,u)}\Big)^{\!2}\,T\Big)\;\leq\;2\exp\!\Big(-\ln\tfrac{2}{p}-\tfrac{1}{\pi(s,u)}\Big)\;\leq\;\tfrac{p}{2}\,\pi(s,u),$
where the last inequality uses $e^{-1/x}\leq x/2$ for $x\in(0,1]$, absorbed into the constant in $T$. Summing over $u$ and using $\sum_{u}\pi(s,u)=1$ gives total estimation failure at most $p/2$. Together with the probability in discovery phase, the overall failure probability is at most $p$. ∎
On the choice of constants.
The constants $20,10$ in $T_{2}$ and $\theta$ ensure (a) a factor-2 gap between the mean and the threshold for nodes in $V_{\geq\delta}$ (enabling a strong lower-tail bound), and (b) a large multiplicative overshoot $c>3.5$ for nodes with mass below $\delta/(5e)$ (enabling a strong upper-tail bound). The margin $\delta/(5e)$ is chosen so that (i) $\ln(1/(\delta/(5e)))=\ln(1/\delta)+\operatorname{\mathcal{O}}(1)$, preserving asymptotics, and (ii) the Round 2 thresholds make the Chernoff exponents at least $2.5L$, which yields clean union bounds. The factors $256$ and $1024$ in $T$ are convenient absolute constants chosen to absorb slack in Chernoff and union bounds; any sufficiently large universal constants would work without changing the stated asymptotics.
Combining the discovery and estimation phases yields the claimed $\operatorname{\mathcal{O}}(\log(1/\delta)/\delta)$ and $\operatorname{\mathcal{O}}(1/\varepsilon^{2})$ query upper bounds in the arc-centric model, as required by Theorem 4.
5 Proof Framework: Lower Bound of SSPPR-R
In this section, we present the proof framework for deriving Theorem 5. We couple the proof details of Theorem 6 as they share similar hard instance structure. We begin bt first introducing the high-level ideas.
High-Level Ideas.
We begin by constructing a carefully designed graph instance family, denoted by $\mathcal{U}(n,D,d)$, parameterized by integers $n$, $D$, and $d$ to be specified later. Each instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ contains a distinguished source node $s$ and $2n$ target nodes, evenly partitioned into two disjoint sets: $X_{1}$ and $X_{2}$, as illustrated in Figure 1(a). By carefully engineering the edge structure, we ensure that every node $x_{1}\in X_{1}$ has a higher PPR value from $s$ than any node $x_{2}\in X_{2}$, satisfying $\pi(s,x_{1})-\pi(s,x_{2})\geq 2\varepsilon$ for SSPPR-A, or $\pi(s,x_{1})/\pi(s,x_{2})\geq 2/(1-c)$ for SSPPR-R, as shown in Lemmas 18 and 19, respectively. That is, nodes in $X_{1}$ and $X_{2}$ differ in their PPR values by at least the required additive or relative error threshold. To succeed in this setting, any algorithm $\operatorname{\mathcal{A}}$ must detect structural differences between nodes in $X_{1}$ and $X_{2}$, which are intentionally subtle and primarily reflected in degree statistics. To model this formally, we define a uniform distribution $\Sigma(n,D,d)$ over the legal graph instances, as described in Lemma 20. This distribution is constructed by considering all valid permutations of node and edge labels and all balanced splits of the $X$ node set. Under this formulation, we can quantify the success probability of any algorithm $\operatorname{\mathcal{A}}$ operating under a query budget $T$. A subtle technical challenge arises from the potential presence of parallel edges, which may occur in our multi-graph construction. Lemma 22 guarantees that the maximum multiplicity of parallel edges is bounded by a constant $L$ with high probability. Combined with Theorem 23, this ensures that any algorithm $\operatorname{\mathcal{A}}$ designed for simple graphs can be adapted to work on multi-graphs without incurring significant additional query complexity.
Bringing these components together, we establish Theorem 24, which asserts that if the query budget falls below a certain threshold—specifically, $T=o\left(\min(m,\frac{\log(1/\delta)}{\delta})\right)$ for the SSPPR-R problem—then no algorithm $\operatorname{\mathcal{A}}$ can correctly recover the $X_{1}$$X_{2}$ split with constant probability. Indeed, Lemma 25 shows that under such constraints, the expected success probability is upper-bounded by $1/n$. As a result, any algorithm that satisfies the SSPPR query guarantees must make at least $\Omega\left(\min(m,\frac{\log(1/\delta)}{\delta})\right)$ queries. This completes the proof of the lower bounds stated in Theorems 5 and 6.
(a) Instance $\mathcal{U}$
(b) $r$-padded instance $\mathcal{U}(r)$
Figure 1: Graph instance family $\mathcal{U}(n,D,d)=\mathcal{U}(3,2,1)$. (a) denotes our base graph instance family $\mathcal{U}(n,D,d)$ and (b) represents the $r$-padded instance $\mathcal{U}(r)$ based on $\mathcal{U}$.
5.1 Hard Instance Family — $\mathcal{U}(n,D,d)$
We define the graph family $\mathcal{U}(n,D,d)$ such that, for every sufficiently large $n$, it serves as a collection of hard instances for estimating SSPPR queries. This family is constructed based on a simple yet insightful core structure $\mathbf{U}$, illustrated in Figure 1(a). The graph consists of five disjoint node sets: $\{s\}\cup Y_{1}\cup Y_{2}\cup X_{1}\cup X_{2}$. For brevity, we let $Y=Y_{1}\cup Y_{2}$ and $X=X_{1}\cup X_{2}$. The node $s$ serves as the designated source node, while all nodes in $X$ act as target nodes whose PPR values with respect to $s$ are of interest. We ensure that each of the sets $X_{1}$, $X_{2}$, $Y_{1}$, and $Y_{2}$ contains $\Theta(n)$ nodes. Regarding edge construction, each node $y_{1}\in Y_{1}$ receives an incoming edge from the source node $s$. For each node $x_{1}\in X_{1}$, we construct $D$ incoming edges from nodes in $Y_{1}$ (i.e., $\deg_{Y_{1}}(x_{1})=D$) and DdD-d edges from $Y_{2}$ (i.e., $\deg_{Y_{2}}(x_{1})=D-d$). Similarly, each $x_{2}\in X_{2}$ receives $D$ edges from $Y_{2}$ and DdD-d from $Y_{1}$. The critical difference is that nodes in $X_{1}$ are indirectly reachable from $s$ via a useful path through $Y_{1}$, whereas nodes in $X_{2}$ are only connected through useless paths via $Y_{2}$ that do not connect to $s$. This design ensures a measurable PPR gap between $X_{1}$ and $X_{2}$. Additionally, we impose that every node $y\in Y$ has out-degree $d_{\text{out}}(y)=2D-d$ to ensure uniformity and support our lower-bound construction. With appropriately chosen parameters, we show that nodes in $X_{1}$ and $X_{2}$ exhibit a clear and quantifiable difference in their PPR values with respect to $s$. This distinction is formally captured in Lemma 18 for the SSPPR-A query and in Lemma 19 for the SSPPR-R query. To facilitate analysis in the SSPPR-R setting, we next introduce a modified structure, referred to as the $r$-padded instance, which better accommodates the relative-error guarantees.
Definition 16 ($r$-padded instance).
$\forall\mathrm{U}\in\mathcal{U}(n,D,d)$, we define the $r$-padded instance of $\mathrm{U}$, denoted as $\mathrm{U}(r)=(V_{U(r)},E_{U(r)})$ in the below by involving two independent node sets $Z_{X}$ and $Z_{Y}$:
  1. 1.
    $V_{U(r)}=V_{U}\cup Z_{X}\cup Z_{Y}$, where $Z_{X},Z_{Y}$ are two additional node sets containing exactly $r$ nodes;
  2. 2.
    $E_{U(r)}=E_{U}\cup\{(z_{x},x)\;\mid\forall x\in X,z_{x}\in Z_{X}\}\cup\{(y,z_{y})\;\mid\forall y\in Y,z_{y}\in Z_{Y}\}$, which incorporates extra edges coming from node set $Z_{x}$ to $X$ and $Y$ to $Z_{Y}$.
An illustrative example is provided in Figure 1(b). The $r$-padded instance $\mathrm{U}(r)$ is specifically designed to facilitate the analysis of SSPPR-R queries. Intuitively, when $r$ is sufficiently large relative to $D$, each ADJ query originating from a node in $X$ or $Y$ may be misdirected to the auxiliary node sets $Z_{X}$ or $Z_{Y}$ with non-negligible probability. This increases the expected query cost for algorithms attempting to explore useful edges. With this construction in place, we now proceed to compute the PPR values from the source node $s$ to nodes in the sets $X_{1}$ and $X_{2}$, as formalized in the following lemma:
Lemma 17 (PPR Calculation).
For any instance $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following properties hold:
  1. (i).
    The graph $\mathrm{U}$ contains $4n+1$ nodes and 2n(2Dd)+n2n(2D-d)+n edges;
  2. (ii).
    For each node $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$, the Personalized PageRank values with respect to the source node $s$ are $\pi(s,x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d)n}$ and $\pi(s,x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d)n}$.
Moreover, for the $r$-padded instance $\mathrm{U}(r)$ with $r\geq D$, we have:
  1. (i).
    The graph $\mathrm{U}(r)$ contains $4n+2r+1$ nodes and 2n(2Dd+2r)+n2n(2D-d+2r)+n edges;
  2. (ii).
    For each node $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$, the PPR values are $\pi(s,x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d+r)n}$ and $\pi(s,x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d+r)n}$, and in particular, $\pi(s,x_{2})=\frac{D-d}{D}\cdot\pi(s,x_{1})$.
Proof.
The number of nodes and edges in both instances $\mathrm{U}$ and $\mathrm{U}(r)$ follows directly from their definitions. We now compute the PPR values from the source node $s$ on the instance $\mathrm{U}$. First, since each node $y\in Y_{1}$ receives one incoming edge from $s$, and there are $n$ such nodes, it holds that $\pi(s,y)=\frac{1-\alpha}{n}$. Next, consider nodes $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$. Each node in $X_{1}$ receives $D$ in-edges from $Y_{1}$, and each node in $X_{2}$ receives DdD-d in-edges from $Y_{1}$. Since every $y\in Y_{1}$ has out-degree 2Dd2D-d, the probability of transitioning from a given $y$ to any of its children is $\frac{1}{2D-d}$. By aggregating over all paths from $s$ to $x_{1}$ via $Y_{1}$, we obtain $\pi(s,x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d)n}$ and $\pi(s,x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d)n}$. The difference is then $\pi(s,x_{1})-\pi(s,x_{2})=(1-\alpha)^{2}\cdot\frac{d}{(2D-d)n}$. For the padded instance $\mathrm{U}(r)$, the analysis is similar. Each node in $Y$ now has out-degree 2Dd+r2D-d+r due to the extra outgoing edges to $Z_{Y}$. Thus, the transition probabilities are updated accordingly, and we obtain $\pi(s,x_{1})=(1-\alpha)^{2}\cdot\frac{D}{(2D-d+r)n}$ and $\pi(s,x_{2})=(1-\alpha)^{2}\cdot\frac{D-d}{(2D-d+r)n}$. In particular, this yields the ratio $\pi(s,x_{2})=\frac{D-d}{D}\cdot\pi(s,x_{1})$. This completes the proof. ∎
Equipped with the above lemma, we now carefully select the graph parameters $n$, $D$, and $d$, along with the error tolerances $\varepsilon$ and $\delta$, to ensure that the PPR values of nodes in $X_{1}$ and $X_{2}$ exhibit a controlled but noticeable difference.
Lemma 18 (SSPPR-A).
Choose any small constant $\beta\in(0,1)$, decay factor $\alpha$ and any functions $\varepsilon_{0}(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n^{2-\beta})$. For sufficiently large $n_{0}$, there exists graph parameters $n,D,d$ (as function of $n_{0}$) s.t. for all graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following conditions hold:
  1. (i).
    The node count of $\mathrm{U}$ is in $\operatorname{\mathcal{O}}(n_{0})$, and the edge count $\mathrm{U}$ is in $\operatorname{\mathcal{O}}(m_{0})$;
  2. (ii).
    $(n)^{1-\beta}\geq D\geq 2d$ and $d\leq\log(n)$;
  3. (iii).
[…]
  4. (iv).
    For all node $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$, their PPR values satisfy
    $\pi(s,x_{1})-\pi(s,x_{2})>2\varepsilon_{0}.$
Proof.
We manipulate the parameters $n$ and $D$ to fulfill the conclusion. Let $C$ be a constant to be determined later. We set $C\beta^{2}\min\left(n_{0}^{2-\beta},m_{0},\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}}\right), nDn\geq D$, and $d=\log(n)$. Note that $d\geq\frac{1}{2}\log(nD)$. From Lemma 17, we have:
$\pi(s,x_{1})-\pi(s,x_{2})$ $=\frac{d}{2D-d}\cdot\frac{(1-\alpha)^{2}}{n}\geq\frac{\log(nD)}{8Dn}\cdot(1-\alpha)^{2}\geq\frac{\log(C\beta^{2}\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})}{8C\beta^{2}\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}}}\cdot(1-\alpha)^{2}$
$\geq\frac{(1-\alpha)^{2}}{8C\beta^{2}}\frac{\log(\frac{1}{\varepsilon_{0}})+log(C\beta^{2})+\log\log(\frac{1}{\varepsilon_{0}})}{\log(\frac{1}{\varepsilon_{0}})}\varepsilon_{0}\geq\frac{(1-\alpha)^{2}}{16C\beta^{2}}\varepsilon_{0}$
The last inequality requires $\log(C\beta)+\log(\frac{1}{\varepsilon_{0}})>\frac{1}{2}\log(\frac{1}{\varepsilon_{0}})$, which further needs $\varepsilon_{0}<(C\beta)^{2}$. Setting $C=\Theta((1-\alpha)^{2})$ directly accomplishes the property $(iv)$. Next, we need to fit the conditions $(i),(ii)$ and $(iii)$ to accomplish the proof. We give an overall setting of
$n=\left(C\beta^{2}\min(n_{0}^{2-\beta},m_{0},\log(\frac{1}{\varepsilon_{0}})/{\varepsilon_{0}})\right)^{\frac{1}{2-\beta}}\ \text{and}\ D=\left(C\beta^{2}\min(n_{0}^{2-\beta},m_{0},\log(\frac{1}{\varepsilon_{0}})/{\varepsilon_{0}})\right)^{\frac{1-\beta}{2-\beta}}.$
Condition $(i)$ holds since the total number of edges is at most $2n(2D-d)+n\leq 4nD=\Theta(\beta m_{0})$, and the number of nodes is $4n+1=\Theta(\beta n_{0})$. For condition (ii), it is clear that $n^{1-\beta}\geq D$, and since $2d=2\log n\leq\log\left((C\beta^{2})^{2/(2-\beta)}n_{0}\right)$, we have $2d\leq(C\beta^{2}n_{0})^{(1-\beta)/(2-\beta)}\leq D$ for sufficiently large $n$. Finally, condition $(iii)$ holds directly by the definition of $nD$. ∎
Lemma 19 (SSPPR-R).
Choose any constant $c\in(0,\frac{1}{2}]$ and any functions $\varepsilon_{0}(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n^{2})$. For sufficiently large $n_{0}$, there exists graph parameters $n,D,d,r$ (as function of $n_{0}$), such that for $r$-padded instance $\mathrm{U}(r)$ based on all graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following conditions hold:
  1. (i).
    The node count of $\mathrm{U}(r)$ is in $\operatorname{\mathcal{O}}(n_{0})$, and the edge count $\mathrm{U}(r)$ is in $\operatorname{\mathcal{O}}(m_{0})$;
  2. (ii).
    $(n)^{\frac{1}{2}}>D\geq\frac{3}{2}d$ and $d\leq\log(n)$;
  3. (iii).
    $n(r+D)=\Omega\left(\min(m,\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$;
  4. (iv).
    For all $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ with PPR values satisfying $\pi(s,x_{1})\geq\delta_{0}$ and $\pi(s,x_{2})\geq\delta_{0}$, we have
    $\pi(s,x_{1})\geq\frac{1}{(1-c)^{2}}\cdot\pi(s,x_{2}).$
Proof.
We consider two cases with an overall setting of $\delta$ such that $\frac{\log\frac{1}{\delta}}{\delta}=\min(cm_{0},\frac{\log\frac{1}{\delta_{0}}}{\delta_{0}}).$
Case 1:
$n_{0}\leq\frac{1}{32}(1-\alpha)^{2}\frac{1}{\delta}.$ And we here set $n=n_{0},\quad d=\log(n_{0}),\quad D=\left(1+\frac{1}{4c}\right)d$ and $r=\frac{(1-\alpha)^{2}d}{16cn_{0}\delta}>0.$ Under these parameters, conditions (i), (ii), and (iii) naturally hold when $n$ is sufficiently large. Moreover for condition (iv), we have
$nr=\Omega\left((1-\alpha)^{2}d/\delta c\right)=\Omega\left(\min(m_{0},\frac{\log\frac{1}{\delta_{0}}}{\delta_{0}})\right).$
By Lemma 17, the PPR value for any $x_{2}\in X_{2}$ satisfies
$\pi(s,x_{2})=(1-\alpha)^{2}\frac{D-d}{(2D-d+r)n_{0}}=\frac{(1-\alpha)^{2}}{(4c+2)n_{0}+\frac{(1-\alpha)^{2}}{4\delta}}\geq\delta\geq\delta_{0}.$
Case 2:
$n_{0}>\frac{1}{32}(1-\alpha)^{2}\frac{1}{\delta}.$ Under this setting, we set $r=0$ since no padded instance is required here. Further, we give $n=\frac{1}{32}(1-\alpha)^{2}\frac{1}{\delta}$ and $D=\left(1+\frac{1}{4c}\right)d.$ Similarly, for sufficiently large $n$, conditions (i), (ii), and (iii) hold. Additionally, it holds that $nD=nd\left(1+\frac{1}{4c}\right)=\Omega\left(\frac{\log(\frac{1}{\delta})}{\delta}\right)=\Omega\left(\min(m_{0},\frac{\log\frac{1}{\delta_{0}}}{\delta_{0}})\right)$ and $\pi(s,x_{2})=\frac{(1-\alpha)^{2}}{n(4c+2)}\geq\delta\geq\delta_{0}$, which ensures condition (iv) and thus we complete the proof. ∎
Lemma 18 shows that the node sets $X_{1}$ and $X_{2}$ differ in their PPR values by more than twice the error parameter $\varepsilon_{0}$. Consequently, any algorithm $\operatorname{\mathcal{A}}$ must distinguish between these two sets with constant success probability, otherwise we argue that it is impossible to distinguish the two node sets with expected query cost at $o(\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}})$ in later sections.
Notice that under the assumption $\min(m_{0},\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}})\leq n_{0}^{2-\beta}$, the instance family we construct contains $nD=\Theta\left(\min(m_{0},\frac{\log(1/\varepsilon_{0})}{\varepsilon_{0}})\right)$ edges, which is proportional to the total query complexity stated in Theorem 6. To prove this main theorem, it suffices to show that there exists an absolute constant $C_{A}$ such that any algorithm attempting to solve the SSPPR-A problem with expected query cost at most $C_{A}\cdot(1-p)nD$ must fail on our constructed graph family with probability at least $p$. By Markov’s inequality, we can enforce that the algorithm halts and raises an error if it performs more than $C_{A}\cdot\frac{nD}{\beta}$ queries, and subsequently demonstrate the existence of a graph instance on which the failure probability is at least $\frac{p}{2}$.
A similar argument applies in the SSPPR-R setting, where the expected query budget is bounded by $C_{R}\cdot nD$ for some constant $C_{R}$. To derive these conclusions, we first formalize the execution of a specific query $Q$ and its corresponding response $R$ as an event on an instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ with assigned node and edge labels.
Node and Edge Labeling in $\mathcal{U}(n,D,d)$.
In accordance with the model specified in Definition 11, our graph instances are equipped with both node and edge labels. While our overarching goal is to preserve in-distinguishability among instances within the family, we partially fix the labeling scheme to ensure clarity and consistency throughout the construction process described below:
  1. (i).
    The source node $s$ is assigned label $1$;
  2. (ii).
    Nodes in $X$ are labeled consecutively from $2$ through $2n+1$;
  3. (iii).
    Nodes in $Y_{1}$ are labeled from $2n+2$ to $3n+1$, and nodes in $Y_{2}$ from $3n+2$ to $4n+1$;
  4. (iv).
    The $i^{\text{th}}$ out-edge of the source node $s$ points to the node labeled $2n+1+i$.
For ease of reference, we adopt the following conventions in place of explicit numeric labels:
  1. (i).
    $s$ denotes the node labeled $1$, which has in-degree zero and out-degree $n$;
  2. (ii).
    $x_{1},x_{2},\ldots,x_{2n}$ refer to the nodes labeled $2$ through $2n+1$, each with in-degree 2Dd2D-d and out-degree zero;
  3. (iii).
    $y_{1},y_{2},\ldots,y_{2n}$ correspond to the nodes labeled $2n+2$ through $4n+1$, each with in-degree one and out-degree 2Dd2D-d;
  4. (iv).
    We use $y^{(1)}_{1},\ldots,y^{(1)}_{n}$ to refer to nodes $y_{1},\ldots,y_{n}\in Y_{1}$, and $y^{(2)}_{1},\ldots,y^{(2)}_{n}$ for $y_{n+1},\ldots,y_{2n}\in Y_{2}$. For edge references, we denote $(y^{(1)}_{i},k)$ as the $k^{\text{th}}$ out-edge from node $y^{(1)}_{i}\in Y_{1}$, and analogously for nodes in $Y_{2}$.
The above notations serve interchangeably to denote both the node label and the corresponding node in the graph. Note that the labeling scheme clearly identifies whether a node belongs to $Y_{1}$, $Y_{2}$, or $X$, but it does not distinguish between the subsets $X_{1}$ and $X_{2}$. Additionally, once a node’s label is known, its in-degree and out-degree are assumed to be fully determined.
We now illustrate how queries operate on a graph $\mathrm{U}\in\mathcal{U}(n,D,d)$ under this labeling convention. Each query-response pair $(Q,R)$ corresponds to an observable property of $\mathrm{U}$. For example, consider the node labeled $x_{1}\in X$ and the node labeled $y^{(1)}_{1}\in Y_{1}$. Suppose the $3^{\text{rd}}$ in-edge of $x_{1}$ corresponds to the $2^{\text{nd}}$ out-edge of $y^{(1)}_{1}$. Then the query $Q={\textsf{ADJ}}_{\text{in}}(x_{1},3)$ yields the response $R=(y^{(1)}_{1},2)$. Conversely, for any graph $\mathrm{U}_{0}\in\mathcal{U}(n,D,d)$, if executing the query $Q={\textsf{ADJ}}_{\text{in}}(x_{1},3)$ returns the response $R=(y^{(1)}_{1},2)$, then $\mathrm{U}_{0}$ must exhibit exactly this structural property. Under this procedure, any query-response pair $(Q,R)$ is referred to as an event with respect to a graph instance distribution, i.e., $\Sigma(n,D,d)$ (defined later in Section 5.2). For any $\sigma\sim\Sigma(n,D,d)$, it corresponds to a graph $U_{\sigma}$ with fixed edges and node labels. For brevity, we write such an event as $Q=R$ when the context is clear.
5.2 Distribution $\Sigma(n,D,d)$
Figure 2: An example for generating graph distribution $\Sigma(n,D,d)=\Sigma(2,2,1)$.
We now describe the procedure for generating an instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ with assigned node and edge labels, which together define the distribution $\Sigma(n,D,d)$ over all possible label permutations. An illustrative example is provided in Figure 2. Intuitively, $\Sigma(n,D,d)$ is constructed by randomly splitting the node set $X$, thereby determining which node labels in $X$ belong to $X_{1}$ or $X_{2}$, and then independently sampling every possible edge permutation that is consistent with the hard-instance template. The complete construction procedure is detailed as follows:
  1. (i).
    Split of $X$. We first choose a split $\mathrm{SP}$ of the index set $\{1,\dots,2n\}$ into two disjoint subsets of size $n$, which determines the split of the node set $X$ into $X_{1}$ and $X_{2}$. Specifically, the nodes labeled $x_{1},\dots,x_{2n}$ are assigned to $X_{1}$ or $X_{2}$ according to the selected split. We denote the set of all possible such splits by $\mathrm{SP}\in\binom{X}{n}$. For example, in Figure 2, the last column lists nodes $x_{1},\dots,x_{4}$ (from top to bottom), and the chosen split $\mathrm{SP}=\{1,3\}$ assigns nodes labeled $x_{1}$ and $x_{3}$ to $X_{1}$, while nodes $x_{2}$ and $x_{4}$ belong to $X_{2}$;
  2. (ii).
    Permuting in-edges of nodes in $X$. For each node $x_{i}\in X$, where $1\leq i\leq 2n$, we independently apply a permutation $\mathrm{P}_{i}^{2D-d}\in\mathrm{S}^{2D-d}\mathrel{\mathop{\ordinarycolon}}\{1,\dots,2D-d\}\rightarrow\{1,\dots,2D-d\}$, determining the order of its 2Dd2D-d in-edges. These in-edges are then divided into two groups: those coming from $Y_{1}$ and those from $Y_{2}$. Given the fixed split $\mathrm{SP}$, if $x_{i}\in X_{1}$, the first $D$ edges in the permuted order are assigned to originate from nodes in $Y_{1}$, and the remaining DdD-d edges from $Y_{2}$; the assignment is reversed if $x_{i}\in X_{2}$. We denote the space of all such independent permutations over the node set $X$ as $\mathrm{S}^{2D-d}(X)$. As illustrated in Figure 2, this process is concretely depicted using virtual nodes in the second to fourth columns. Each virtual node in the fourth column represents an in-edge of a corresponding node in $X$. For example, the fourth virtual node in the fourth column corresponds to the first in-edge of $x_{2}$. Applying the permutation $\mathrm{P}_{3}=(1,3,2)$, the first and third in-edges of $x_{2}$ originate from $Y_{2}$ (shown as white-filled nodes), while the second originates from $Y_{1}$ (gray-filled);
  3. (iii).
    Permuting out-edges of nodes in $Y_{1}$ and $Y_{2}$. We apply two independent permutations $\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}\in\mathrm{S}^{n(2D-d)}$ over the set $\{1,\dots,n(2D-d)\}$ to determine the ordering of the out-edges from nodes in $Y_{1}$ and $Y_{2}$, respectively. These permutations govern whether the $k^{\text{th}}$ out-edge of a node $y^{(1)}_{i}\in Y_{1}$ (or $y^{(2)}_{i}\in Y_{2}$) connects to a node in $X_{1}$ or $X_{2}$. Specifically, the $k^{\text{th}}$ out-edge of $y^{(1)}_{i}$ points to a node in $X_{1}$ if $\mathrm{P}_{Y_{1}}^{n(2D-d)}[(2D-d)(i-1)+k]\leq nD$; otherwise, it connects to a node in $X_{2}$. An analogous rule applies for nodes in $Y_{2}$ via $\mathrm{P}_{Y_{2}}^{n(2D-d)}$. We denote the collections of such permutations as $\mathrm{S}^{n(2D-d)}(Y_{1})$ and $\mathrm{S}^{n(2D-d)}(Y_{2})$, respectively. This step corresponds to determining connections between virtual nodes in the second and third columns of Figure 2. For example, if we choose $\mathrm{P}_{Y_{1}}^{n(2D-d)}=(1,3,2,4,5,6)$, then the first four virtual nodes in the second column represent out-edges connecting to nodes in $X_{1}$, as determined by the corresponding permutation values;
  4. (iv).
    Establishing connections between nodes in $X$ and $Y$. Finally, we determine the actual connections between nodes in $X$ and $Y$. Without loss of generality, we describe the connection process for the pair $X_{1}$ and $Y_{1}$; all other pairs follow analogously by swapping the roles of $D$ and DdD-d and adjusting indices accordingly. We begin by selecting a random bijection $\mathrm{S}(Y_{1},X_{1})\mathrel{\mathop{\ordinarycolon}}\{1,\ldots,nD\}\rightarrow\{1,\ldots,nD\}$ to match out-edges of nodes in $Y_{1}$ with in-edges of nodes in $X_{1}$. (For other pairings, such as $Y_{2}$ to $X_{1}$, we use a corresponding bijection $\mathrm{S}(Y_{2},X_{1})$ over $\{1,\ldots,n(D-d)\}$.)
    From previous steps, the $nD$ selected out-edges from $Y_{1}$ are encoded as a set of tuples $\{(y^{(1)}_{i_{j}},k_{j})\}_{j=1}^{nD}$, where each tuple represents the $k_{j}^{\text{th}}$ out-edge of node $y^{(1)}_{i_{j}}$. Likewise, the $nD$ selected in-edges to $X_{1}$ are represented as $\{(x^{(1)}_{i_{j}^{\prime}},k_{j}^{\prime})\}_{j=1}^{nD}$, with each tuple representing the $k_{j}^{\prime\text{th}}$ in-edge of node $x^{(1)}_{i_{j}^{\prime}}$. Both edge sets are naturally ordered; for instance, each out-edge $(y^{(1)}_{i_{j}},k_{j})$ can be uniquely indexed by the expression $(2D-d)i_{j}+k_{j}$. The bijection $\mathrm{S}(Y_{1},X_{1})$ then defines the matching: the $k_{j}^{\text{th}}$ out-edge of node $y^{(1)}_{i_{j}}$ is connected to the $(k_{j}^{\prime})^{\text{th}}$ in-edge of node $x^{(1)}_{i_{j}^{\prime}}$. In Figure 2, the mapping $\mathrm{S}(Y_{1},X_{1})=(4,3,2,1)$ is visually indicated by edges between the virtual nodes (shown as black-filled) in the third and fourth columns.
By the above construction, we can then derive the following lemmas and definitions, which recover a graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$ accordingly.
Lemma 20.
For all possible $\mathrm{SP},\mathrm{P}_{i}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)},\mathrm{P}\in\binom{X}{n}\times\mathrm{S}^{2D-d}(X)\times\mathrm{S}^{n(2D-d)}(Y_{1})\times\mathrm{S}^{n(2D-d)}(Y_{2})\times\Pi_{Y_{a}\in\{Y_{1},Y_{2}\},X_{b}\in\{X_{1},X_{2}\}}S(Y_{a},X_{b})$: the above procedure links a legal graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$.
Definition 21.
$\Sigma(n,D,d)$ is defined as the uniform distribution over $\binom{X}{n}\times\mathrm{S}^{2D-d}(X)\times\mathrm{S}^{n(2D-d)}(Y_{1})\times\mathrm{S}^{n(2D-d)}(Y_{2})\times\Pi_{Y_{i}\in\{Y_{1},Y_{2}\},X_{i}\in\{X_{1},X_{2}\}}S(Y_{i},X_{i})$. For $\sigma\sim\Sigma(n,D,d)$, we use $\mathrm{U}_{\sigma}$ to denote the corresponding labeled graph instance $\mathrm{U}\in\mathcal{U}(n,D,d)$. For ease of presentation, we write $\mathrm{U}\sim\Sigma(n,D,d)$ to denote a graph instance sampled from this distribution.
Based on the above construction and definitions, we now formulate the probability that a graph instance contains parallel edges in the following lemma:
Lemma 22.
(proof in Appendix A.2) $\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\bigl[\text{edge multiplicity of }\mathrm{U}\geq L\bigr]\leq 4n^{2}(\frac{4D}{n})^{L}$, for $1\leq L\leq D$.
Lemma 22 demonstrates that, with high probability, the edge multiplicity in our constructed graph instances is bounded by a small constant. In other words, high multiplicity edges rarely appear under our construction. Leveraging this result, we show in the next section that any algorithm designed for simple graphs can be adapted to handle multi-graphs with bounded edge multiplicity. From Lemmas 19 and 18, since $D$ is upper bounded by a sublinear polynomial in $n$ (see condition (ii) in both lemmas), we can choose a constant $L$ such that
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\bigl[\text{edge multiplicity of }\mathrm{U}>L\bigr]\leq\frac{1}{n}.$
In particular, we set $L=\frac{8}{\beta}$ for the SSPPR-A setting and $L=16$ for SSPPR-R.
5.3 Establishing Main Results
To prove our main results, such as Theorem 5, we assume the existence of an algorithm $\operatorname{\mathcal{A}}$ that successfully answers the SSPPR-R problem with lower query complexity than our target lower bound. We then derive a contradiction by showing that the overall success probability over the graph distribution becomes $\operatorname{\mathcal{O}}(1/n)$, which violates the requirement of constant success probability. Note that $\operatorname{\mathcal{A}}$ is designed for simple graphs, whereas instances $\mathrm{U}\sim\Sigma(n,D,d)$ may admit parallel edges. Fortunately, our earlier analysis guarantees that the edge multiplicity is bounded by a constant with high probability. The following theorem provides a mechanism for executing $\operatorname{\mathcal{A}}$ on such multi-graphs.
Lemma 23.
(proof in Appendix A.3) [Reduction from simple to multi-graphs] Choose any constant $L$, decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, constant $\beta\in(0,1)$, and any function $\varepsilon(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2-\beta}),T(n_{0})>0$ (Or, error bound $c\in(0,\frac{1}{2}]$, and any function $\delta(n_{0})\in(0,1),m(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2})$ for SSPPR-R setting). Suppose there exists an algorithm $\operatorname{\mathcal{A}}$ such that, for any simple directed graph with $\Theta(Ln_{0})$ nodes and $\Theta(Ln_{0}+Lm_{0})$ edges, $\operatorname{\mathcal{A}}$ answers the SSPPR-A problem with decay factor $\alpha$, error $\varepsilon$ and failure probability bounded by $p$ (or SSPPR-R problem with threshold $\delta$, error bound $c$), and expected query complexity $T$ in arc-centric model.
Then there exists an algorithm $\operatorname{\mathcal{A}}_{L}$ in arc-centric model such that, for any integers $n\leq n_{0},\;D,\;d,\;r,\;n(D+r)\leq m$, and any $\mathrm{U}\in\mathcal{U}(n,D,d)$ with edge multiplicity bounded by $L$, $\operatorname{\mathcal{A}}_{L}$ solves the SSPPR-A problem with decay factor $1-(1-\alpha)^{3}$ and error parameter $\varepsilon$ (or SSPPR-R problem with threshold $\delta$, error requirement $c$) on $\mathrm{U}(r)$ for target nodes in $X$, with expected query complexity bounded by a function of $n_{0}$ in $\operatorname{\mathcal{O}}(T)$ and with failure probability bounded by $p$. Specifically, $\mathcal{A}_{L}$ executes each query ${\textsf{ADJ}}_{in/out}(v,k)$ by a uniformly randomly selection of an uncovered label $k$ among $v$’s in or out edges.
Combining all the preceding results, we arrive at the following theorem, which encapsulates the current milestone of our analysis:
Theorem 24.
Choose any constant $L\geq\frac{16}{\beta}$ , decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, constant $\beta\in(0,1)$, and any function $\varepsilon_{0}(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2-\beta}),T(n_{0})=o\left(\min(m_{0},\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})\right)$ (Or, $L\geq 32$, error bound $c\in(0,\frac{1}{2}]$, and any function $\delta_{0}(n_{0})\in(0,1),m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2}),T(n_{0})=o\left(\min(m_{0},\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$ for SSPPR-R setting). Assume there is an algorithm $\operatorname{\mathcal{A}}$ with failure probability bounded by $p$, and expected query complexity at $T$ solves the SSPPR-A problem (or SSPPR-R problem), Then given arbitrary small constant $\gamma$, when $n_{0}$ is sufficiently large, there exists integers $n$ (sufficiently large),D,d,r,\;D,\;d,\;r such that:
  1. (i).
    For any $\mathrm{U}\in\mathcal{U}(n,D,d)$, $\mathrm{U}(r)$ contains no more than $\beta n_{0}$ nodes and no more than $\beta m_{0}$ edges;
  2. (ii).
    $d\leq\log(n),n\geq 2D,D\geq\frac{3}{2}d$;
And there exists an algorithm $\mathcal{A}^{{}^{\prime}}$ holding that:
  1. (iii).
    $\mathcal{A}^{{}^{\prime}}$ does at most $T(n_{0})\leq\gamma nD$ queries;
  2. (iv).
    For $\mathrm{U}\sim\Sigma(n,D,d)$, the algorithm $\mathcal{A}^{{}^{\prime}}$ can be applied on any multi-graph instance $\mathrm{U}$, with expected probability at least $\frac{1-p}{4}$, $\mathcal{A}^{{}^{\prime}}$ can output the correct split $X_{1},X_{2}$ of $X$.
Proof.
We begin with the SSPPR-R case. For the graph related parameters, Lemma 19 gives the integers n,D,d,rn,\;D,\;d,\;r, directly satisfying the condition $(i)$ and $(ii)$ with $n(D+r)=\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\delta})}{\delta_{0}})\right)$. Next, we’d construct algorithm $\operatorname{\mathcal{A}}^{\prime}$ to fit the condition $(iii)$ and $(iv)$. The steps are described as:
  1. (i).
    $\mathcal{A}$ to $\mathcal{A}_{L}$ for executable on multi graph $\mathrm{U}(r)$. As established in our earlier analysis, for $L=16$, a graph sampled from the distribution $\Sigma(n,D,d)$ admits edge multiplicity at most $L$ with high probability, specifically at least $1-\frac{1}{n}$. By invoking Theorem 23, this guarantees the existence of an algorithm $\mathcal{A}_{L}$ that can solve the SSPPR-R problem on such multi-graphs with expected success probability at least 1p1-p, with expected query complexity of $o(n(D+r))$.
  2. (ii).
    From PPR estimation to split of $X$. To ensure that $\mathcal{A}_{L}$ produces PPR estimates of nodes in $X$ within the specified relative error tolerance $c$, it must be capable of distinguishing the two subsets $X_{1}$ and $X_{2}$, whose PPR values differ by a clear, quantifiable gap. We therefore append a simple post-processing step to $\mathcal{A}_{L}$ that outputs a split of $X$ based on the estimated PPR values. This split is correct if the estimates produced by $\mathcal{A}_{L}$ satisfy the SSPPR-R accuracy requirements. Specifically, by Lemma 19, for any $x_{i_{1}}\in X_{1}$ and $x_{i_{2}}\in X_{2}$, we have $\pi(s,x_{i_{1}})\geq\frac{1}{(1-c)^{2}}\pi(s,x_{i_{2}})$. This implies the existence of a threshold $\pi_{0}$ such that for all $x_{i_{1}}\in X_{1}$, $\pi(s,x_{i_{1}})(1-c)>\pi_{0}$, and for all $x_{i_{2}}\in X_{2}$, $\pi(s,x_{i_{2}})(1+c)<\pi_{0}$. Furthermore, Lemma 19 ensures that $\pi(s,x_{i})\geq\delta$ for all $x_{i}\in X$. If $\mathcal{A}_{L}$ succeeds on $\mathrm{U}(r)$, its estimate $\widehat{\pi}(s,x_{i})$ must satisfy $|\widehat{\pi}(s,x_{i})-\pi(s,x_{i})|\leq c\pi(s,x_{i})$. As a result, for all $x_{i_{1}}\in X_{1}$, we have $\widehat{\pi}(s,x_{i_{1}})\geq(1-c)\pi(s,x_{i_{1}})>\pi_{0}$, and for all $x_{i_{2}}\in X_{2}$, we have $\widehat{\pi}(s,x_{i_{2}})\leq(1+c)\pi(s,x_{i_{2}})<\pi_{0}$. Thus, a simple thresholding rule at $\pi_{0}$ suffices to correctly separate $X$ into $X_{1}$ and $X_{2}$.
  3. (iii).
    From queries on $\mathrm{U}(r)$ to queries on $\mathrm{U}$. In the previous step, we constructed an algorithm that successfully distinguishes between $X_{1}$ and $X_{2}$ using an expected $o(n(D+r))$ queries on the $r$-padded graph $\mathrm{U}(r)$, under the assumption that the edge multiplicity of the underlying graph $\mathrm{U}$ is bounded by $L=16$ with high probability $1-\frac{1}{n}$. We now demonstrate that queries involving the auxiliary padding nodes in $Z_{X}$ and $Z_{Y}$ are not essential. Consequently, the algorithm can be adapted to the original graph $\mathrm{U}$ with an expected query complexity of $o(nD)$.
    Recall that we have not fixed a specific labeling scheme for $\mathrm{U}(r)$ in earlier steps; instead, we simply assigned labels to the padding nodes $Z_{X}$ and $Z_{Y}$ as $4n+2,\dots,4n+2r+1$. The out-edges from each node $y_{i}\in Y$ to $Z_{Y}$ are assigned labels $2D-d+1,\dots,2D-d+r$, and the in-edges from $Z_{X}$ to each $x_{i}\in X$ are labeled analogously. Given this fixed labeling, all information regarding $Z_{X}$ and $Z_{Y}$ is known in advance from the perspective of $\mathcal{A}_{L}$. Since $\mathcal{A}_{L}$ operates only on $\mathrm{U}(r)$, these edges can be hard-coded into the algorithm. We initialize a subgraph $G_{0}$ internal to $\mathcal{A}_{L}$ that stores all such predetermined edges in $\mathrm{U}(r)$—specifically, all the nodes $s$, $X$, $Y$, $Z_{X}$, and $Z_{Y}$, along with the edges from $Z_{X}$ to $X$ and from $Y$ to $Z_{Y}$.
    We then define the modified algorithm $\mathcal{A}^{\prime}_{L}$ as follows. The core idea is to simulate $\mathcal{A}_{L}$ on the original graph $\mathrm{U}$ by internally emulating its behavior on the padded graph $\mathrm{U}(r)$. Whenever $\mathcal{A}_{L}$ issues a degree query or an $\mathsf{Adj}$ query involving a node in the padding sets $Z_{X}\cup Z_{Y}$, $\mathcal{A}^{\prime}_{L}$ returns the predetermined response encoded in $G_{0}$, which is consistent across all $\mathrm{U}(r)$ constructed from some $\mathrm{U}\in\mathcal{U}(n,D,d)$. In the case where $\mathcal{A}_{L}$ issues a query of the form $\mathsf{Adj}^{\mathrm{in}}(x_{i},k)$, we proceed as follows: if k>2Ddk>2D-d, the query refers to a synthetic edge from $Z_{X}$, so we return the (k2D+d+1)(k-2D+d+1)-st node in $Z_{X}$; otherwise, we forward the actual query $\mathsf{Adj}^{\mathrm{in}}(x_{i},k)$ to the graph $\mathrm{U}$, and add the resulting edge to the internal subgraph $G_{0}$. In this way, $\mathcal{A}^{\prime}_{L}$ simulates the interaction between $\mathcal{A}_{L}$ and $\mathrm{U}(r)$ using queries exclusively on $\mathrm{U}$. Therefore, with success probability at least $p$, $\mathcal{A}^{\prime}_{L}$ correctly outputs the split $(X_{1},X_{2})$.
  4. (iv).
    Bounding query complexity of $\mathcal{A}_{L}^{{}^{\prime}}$. We now bound the query complexity of $\mathcal{A}^{\prime}_{L}$. As described in Theorem 23, the algorithm $\mathcal{A}_{L}$ issues queries of the form ${\textsf{ADJ}}_{\mathrm{out}/\mathrm{in}}(v,k)$ by selecting $k$ uniformly at random from the un-queried incident edge labels. For any node $v\in\mathrm{U}$ where fewer than $(D+r)/2$ incident labels have been examined, the probability that a new query returns a real (non-padding) edge is at most (2Dd)/(2Dd+r(D+r)/2)(2D-d)/(2D-d+r-(D+r)/2). Consequently, the expected number of genuine adjacency queries issued by $\mathcal{A}^{\prime}_{L}$ can be bounded by the sum of two terms: the first is $(2D-d)\cdot\mathbb{E}[|\{v\in\mathrm{U}\mid\text{at least }(D+r)/2\text{ incident labels examined}\}|]$; the second is $(2D-d)/(2D-d+r-(D+r)/2)\cdot o((D+r)n)$. Since both terms are asymptotically $o(nD)$, the overall expected query complexity of $\mathcal{A}^{\prime}_{L}$ is $o(nD)$.
For notational simplicity, we refer to the final algorithm constructed in the above theorem as $\operatorname{\mathcal{A}}$. Naturally, we decompose $\operatorname{\mathcal{A}}$ into two components: the first part, denoted $\mathcal{A}_{1}$, is an exploration procedure that performs queries on the graph; the second part, $\mathcal{A}_{2}$, processes the information collected and outputs a split of $X$ into subsets $X_{1}$ and $X_{2}$. The algorithm $\operatorname{\mathcal{A}}$ succeeds if and only if $\mathcal{A}_{2}$ outputs a correct split. Since we are working under the RAM model, we can assume that $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ operate independently—that is, $\mathcal{A}_{2}$ is a separate algorithm whose input is the data written into memory by $\mathcal{A}_{1}$, and whose output is the inferred split of $X$. Given that RAM is finite (though arbitrarily large) and that the only inputs to $\mathcal{A}_{1}$ are the query responses $R_{1},\ldots,R_{T}$, we may treat the final RAM state as a function of the query-response sequence $Q_{1},\ldots,Q_{T},R_{1},\ldots,R_{T}$, along with any internal randomness. According to the definition in 10, this tuple $(Q_{1},\ldots,Q_{T},R_{1},\ldots,R_{T})$ constitutes the observation, while the correct split of $X$ represents the prior distribution. The expected success probability of $\operatorname{\mathcal{A}}$ can thus be formulated as:
$\mathbb{E}_{\mathrm{U}\sim\Sigma(n,D,d)}\bigl[\operatorname{{\bf Pr}}[\operatorname{\mathcal{A}}\text{ succeeds on }U]\bigr]\leq\sum_{Q_{1}=R_{1},\cdots,Q_{T}=R_{T}}\operatorname{{\bf Pr}}_{\operatorname{\mathcal{A}},{\mathrm{U}\sim\Sigma(n,D,d)}}[Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}]$
$\times\max_{\mathrm{\mathrm{SP}}\in\binom{X}{n}}\operatorname{{\bf Pr}}_{\mathcal{U}\sim\Sigma(n,D,d)}[\mathrm{SP}\text{ is the correct split of }\mathrm{U}\mid Q_{1}=R_{1},\cdots,Q_{T}=R_{T}].$ (1)
Note that the notion of an event $Q=R$ was defined in the final part of Section 5.1. Our goal now is to evaluate the conditional probability
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\left[\mathrm{SP}\text{ is the correct split of }\mathrm{U}\,\middle|\,Q_{1},\ldots,Q_{T};\,R_{1},\ldots,R_{T}\right],$
and to show that it is bounded by $\frac{1}{n}$. This result is formally captured in the following Lemma 25.
Lemma 25.
(proof in Appendix A.4) There is a constant $\gamma$, such that. For sufficiently large $n$, $d\leq log(n)$, $D>\frac{3d}{2}$, $T\leq\gamma nD$. Given any query and returning sequence $Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}$, and a split of $X$, $\mathrm{SP}$ we have:
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}[\mathrm{SP}\text{ is the correct split of }\mathrm{U}\mid Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}]\leq\frac{1}{n}.$
The intuition for why this lemma holds is that, considering two naive ways of exploring $\mathrm{U}$, both require $\Theta(nD)$ queries to get the correct split: The first way is a traversal of the graph. For instance in a breadth-first search, one need expected $\Omega(nD)$ queries to get all the information. The second approach is a random walk. For each node $x\in X$, there are at least DdD-d edges from $Y_{1}$ to $X$ and at least DdD-d edges from $Y_{2}$ to $x$, the key for determining whether $x$ is in $X_{1}$ or $X_{2}$ lies in the left $d$ edges. Notice that each random walk of length $1$ from $X$ has probability at most $\frac{d}{D}$ to find one these key edges. Assume we have done $o(D)$ such walks at each point, the probability of finding no key-edges at some node is at least $n(1-\frac{d}{D})^{o(D)}$. Since $d\leq log(n),n(1-\frac{d}{D})^{o(D)}=\Omega(1)$. Details can be found in Appendix A.4.
By Lemma 25, we then conclude that the expected success probability of $\operatorname{\mathcal{A}}$ is less than $\frac{1}{n}$ under the total query limits $T\in o\left(\min(m_{0},\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$ for SSPPR-R problem and $T\in o\left(\min(m_{0},\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})\right)$ for SSPPR-A query. This contradiction illustrates that the assumption in Theorem 24 does not hold. Thus, it requires $\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\varepsilon_{0}})}{\varepsilon_{0}})\right)$ queries for SSPPR-R and $\Omega\left(\min(m_{0},\frac{\log(\frac{1}{\delta_{0}})}{\delta_{0}})\right)$ for SSPPR-A problems in expectation, which finally leads to our main lower bound results in Theorem 5 and Theorem 6.
6 Lower Bound of STPPR
In this section, we present the proof of Theorem 7. The overall argument follows the same structure as the proof of Theorem 5 in Section 5. The key difference lies in the construction of the hard-instance graph family: to create the required discrepancy in target PPR values, we reverse all edge directions in the base graph $\mathrm{U}$, obtaining its inverse family $\mathrm{U}^{-1}$ as depicted in Figure 3. The details of this construction and its properties are formalized in Lemma 26 below.
Figure 3: Instance $\mathcal{U}^{-1}(r)$
Lemma 26 (STPPR-R).
Let $\mathrm{U}^{-1}$ denote the inverse graph of a graph $\mathrm{U}$ obtained by reversing the direction of each edge. Choose any decay factor $\alpha\in(0,1)$, error parameter $c\leq\tfrac{1}{2}$, and functions $\delta_{0}(n_{0})\in\bigl(0,\tfrac{(1-\alpha)^{2}}{16}\bigr)$ and $m_{0}(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2})$. For sufficiently large $n_{0}$, there exist graph parameters $n,D,d$ (as functions of $n_{0}$) such that for all graph instances $\mathrm{U}\in\mathcal{U}(n,D,d)$, the following conditions hold for $\mathrm{U}^{-1}$:
  1. (i).
    The node count of $\mathrm{U}$ is in $\operatorname{\mathcal{O}}(n_{0})$, and the edge count of $\mathrm{U}$ is in $\operatorname{\mathcal{O}}(m_{0})$.
  2. (ii).
    $n^{1/2}\geq D\geq 2d$ and $d\leq\log n$.
  3. (iii).
    $Dn=\Omega\!\left(\min\!\left(m_{0},n_{0}\log n_{0}\cdot\tfrac{1}{\delta_{0}}\right)\right)$.
  4. (iv).
    For all $x_{1}\in X_{1}$ and $x_{2}\in X_{2}$ with PPR values satisfying $\pi(x_{1},s)\geq\delta_{0}$ and $\pi(x_{2},s)\geq\delta_{0}$, we have
    $\pi(x_{1},s)\;\geq\;\frac{1}{(1-c)^{2}}\,\pi(x_{2},s).$
Proof.
As in the original hard instance, in $\mathrm{U}^{-1}$ we have for $x_{i}\in X_{1}$:
$\pi(x_{i},s)=(1-\alpha)^{2}\frac{D}{2D-d+r},$
and for $x_{i^{\prime}}\in X_{2}$:
$\pi(x_{i^{\prime}},s)=(1-\alpha)^{2}\frac{D-d}{2D-d+r},$
so that their ratio remains $\frac{D-d}{D}$.
Consider first the case $m_{0}>n_{0}\log n_{0}\cdot\tfrac{1}{\delta_{0}}$. Set $n=n_{0}$, $d=\log n_{0}$, $D=\left(1+\frac{1}{4c}\right)d$, and
$r=d\left(\frac{(1-\alpha)^{2}}{4c\delta_{0}}-1-\frac{1}{4c}\right).$
The edge count satisfies $O(n(D+r))=O(n_{0}\log n_{0}\tfrac{1}{\delta_{0}})\subset O(m_{0})$; the node count is $O(\max(n,r))=O(\max(n,m/n))=O(n_{0})$. For each $x_{2}\in X_{2}$, $\pi(x_{2},s)=\delta_{0}$, and the ratio condition holds as in Lemma 19.
For the second case, where $\tfrac{1}{\delta_{0}}\leq m_{0}<n_{0}\log n_{0}\cdot\tfrac{1}{\delta_{0}}$, by continuity we can find positive $n,\delta$ such that
$n\log n\cdot\tfrac{1}{\delta}=m_{0},$
with $n\leq n_{0}$, $\delta\geq\delta_{0}$, and $\binom{n}{2}\geq m_{0}$. Setting $d=\log n$, $D=\left(1+\frac{1}{4c}\right)d$, and
$r=d\left(\frac{(1-\alpha)^{2}}{4c\delta}-1-\frac{1}{4c}\right),$
we similarly ensure that $\pi(x_{2},s)=\delta\geq\delta_{0}$, and the ratio condition again holds. Moreover,
$n(r+D)\in\Omega\!\left(\tfrac{nd}{\delta}\right)=\Omega(m_{0}),$
the edge count is $O(n(r+D))=O(\tfrac{nd}{\delta})=O(m_{0})$, and the node count satisfies $O(\max(n,r))=O(n,m/n)\subset O(n_{0})$. ∎
As in the proof of Theorem 5, to estimate $\pi(x_{i},s)$ it is necessary to know the exact partition of $X$ into $X_{1}$ and $X_{2}$. Because in the arc-centric query model a graph and its inverse are indistinguishable, we can apply Lemma 25 to $\mathrm{U}\in\mathcal{U}(n,D,d)^{-1}$. This yields that any algorithm solving STPPR requires
$\Omega\!\left(\min\!\left(m_{0},n_{0}\log n_{0}\cdot\tfrac{1}{\delta_{0}}\right)\right)$
oracle queries, completing the proof of Theorem 7.
7 Lower Bound of SSPPR-A
In this section, we present the proof of Theorem 8.
High-Level Ideas.
In contrast to the relative-error case, the hardness in the absolute-error setting stems not from uniform probability requirements but from fundamental sampling variance when estimating a large Personalized PageRank (PPR) value $\pi(s,t)$. An estimator of $\pi(s,t)$ based on $N$ random walks behaves approximately as $\mathcal{N}(\pi(s,t),\sigma^{2})$ with standard deviation $\sigma=\Theta(\pi(s,t)N^{-1/2})$. Thus, when $\pi(s,t)=\Theta(1)$, achieving an absolute error of $\varepsilon$ with constant success probability inherently requires $N=\Omega(1/\varepsilon^{2})$ walks—an unavoidable statistical barrier.
Our proof embeds a hidden binary matrix into the graph and links $\pi(s,t)$ to the number of $1$’s in this matrix. Since each bit is Bernoulli, the total count of ones follows a binomial law and is well approximated by a normal distribution, paralleling the distribution of the random-walk estimator. Any algorithm must therefore query a large fraction of the graph to recover this count; otherwise, its estimate of $\pi(s,t)$ incurs statistical uncertainty exceeding $\varepsilon$, establishing the desired lower bound.
Graph Family and Distribution
We define a graph family $\mathcal{G}(D,\mathbf{b})$, where $D$ is an integer and $\mathbf{b}$ is a $D\times D$ binary matrix with entries $b_{ij}$. Each graph has $4D+2$ vertices partitioned into six disjoint sets: $\{s\}\cup\{t\}\cup X\cup Y\cup X^{\prime}\cup Y^{\prime},$ where $s$ and $t$ are the designated source and target nodes, and each of $X,Y,X^{\prime},Y^{\prime}$ contains exactly $D$ nodes. Edges are constructed as follows: - The source $s$ connects to every node in $Y$. - Every node in $X$ connects to the target $t$. - For every pair $(i,j)$, - if $b_{ij}=1$, add edges $(y_{i},x_{j})$ and $(y_{i}^{\prime},x_{j}^{\prime})$; - if $b_{ij}=0$, add edges $(y_{i},x_{j}^{\prime})$ and $(y_{i}^{\prime},x_{j})$. This $\mathbf{b}$-controlled wiring introduces structured randomness essential for the hard-instance arguments in our lower-bound proofs. An intuitive example is shown in Figure 4 with $\mathcal{G}(D,\mathbf{b})$ of
$D=2,\mathbf{b}=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$.
Figure 4: An example for generating graph distribution $\mathcal{G}(D,\mathbf{b})$.
Lemma 27 (Properties of $\mathcal{G}(D,\mathbf{b})$).
For any graph instance $\mathcal{G}(D,\mathbf{b})$, the following properties hold:
  1. (i).
    The graph contains $4D+2$ nodes and $2D(D+1)$ edges.
  2. (ii).
    The out-degree of each node in $Y$ or $Y^{\prime}$ is $D$. The in-degree of each node in $X$ or $X^{\prime}$ is $D$.
  3. (iii).
    $\pi(s,t)=\alpha\left(\sum_{i,j=1}^{D}b_{ij}\right)\frac{(1-\alpha)^{3}}{D^{2}}$.
Proof.
The number of nodes follows directly from the construction: $2$ special nodes ($s$ and $t$) plus $4D$ nodes in the sets $X,Y,X^{\prime},Y^{\prime}$, for a total of $2+4D$ nodes. For the edges, there are $D$ edges from $s$ to $Y$, $D$ edges from $X$ to $t$, and $2D^{2}$ edges between $\{Y,Y^{\prime}\}$ and $\{X,X^{\prime}\}$, giving a total of $2D^{2}+2D\;=\;2D(D+1).$ For property (ii), take any node $y_{i}\in Y$ (or $y_{i}^{\prime}\in Y^{\prime}$). For every $j\in\{1,\dots,D\}$ there is exactly one outgoing edge—either to $x_{j}$ or to $x_{j}^{\prime}$. Hence each $y_{i}$ (and each $y_{i}^{\prime}$) has out-degree exactly $D$. Similarly, for any node $x_{j}\in X$ (or $x_{j}^{\prime}\in X^{\prime}$), each $i\in\{1,\dots,D\}$ contributes exactly one incoming edge (from either $y_{i}$ or $y_{i}^{\prime}$), so every $x_{j}$ (and $x_{j}^{\prime}$) has in-degree exactly $D$. Now we compute the PPR value $\pi(s,t)$ by tracing the probability mass of an $\alpha$-decay random walk starting at $s$. Initially the mass at $s$ is $\alpha$.
  1. 1.
    First step. With probability $(1-\alpha)$ the walk leaves $s$ and spreads uniformly to its $D$ out-neighbors in $Y$. Hence each $y_{i}\in Y$ receives
    $\frac{\alpha(1-\alpha)}{D}.$
  2. 2.
    Second step. Each $y_{i}$ has out-degree $D$ and forwards its mass to nodes in $X$ or $X^{\prime}$ according to $b_{ij}$. The total contribution to a node $x_{j}\in X$ is
    $\sum_{i=1}^{D}b_{ij}\cdot\frac{\alpha(1-\alpha)}{D}\cdot\frac{1-\alpha}{D}=\left(\sum_{i=1}^{D}b_{ij}\right)\frac{\alpha(1-\alpha)^{2}}{D^{2}}.$
  3. 3.
    Third step. Each $x_{j}$ has a single outgoing edge to $t$, passing on all of its mass with an additional $(1-\alpha)$ factor. Summing over $j$ gives
    $\pi(s,t)=\sum_{j=1}^{D}\left[\left(\sum_{i=1}^{D}b_{ij}\right)\frac{\alpha(1-\alpha)^{2}}{D^{2}}\right](1-\alpha)=\alpha\left(\sum_{i,j=1}^{D}b_{ij}\right)\frac{(1-\alpha)^{3}}{D^{2}}.$
This establishes the desired formula for $\pi(s,t)$.
We now define $\Sigma(D)$, a probability distribution over the graph family $\mathcal{G}(D,\mathbf{b})$, which serves as our hard distribution for the lower bound. For each fixed $D$, every entry $b_{ij}$ of the binary matrix $\mathbf{b}$ is chosen independently from a Bernoulli distribution $\operatorname{Bern}(1/2)$. Under this distribution, the total number of ones $S\mathrel{\mathop{\ordinarycolon}}=\sum_{i,j=1}^{D}b_{ij}$ follows a binomial law $\operatorname{Binomial}(D^{2},1/2)$. For node and edge indexing, we adopt the natural scheme: the $i$th outgoing edge of $s$ connects to $y_{i}$, the $j$th incoming edge of $t$ originates from $x_{j}$, and for each $y_{i}$, its $j$th outgoing edge goes to $x_{j}$ or $x_{j}^{\prime}$ depending on the value of $b_{ij}$. Combining all the above preparations, we now proceed to prove the lower bound. As in the relative-error case, the proof naturally decomposes into two key lemmas: (i) establishing the existence of an appropriate hard distribution and its parameters, and (ii) analyzing the conditional distribution of outcomes under a bounded number of oracle queries.
Lemma 28.
Given constants $\alpha\in(0,1)$, $p\in(0,1)$, and $m_{0}\in\Omega(n_{0}),m_{0}\leq\binom{n_{0}}{2},\varepsilon_{0}\in(0,1))$ as function of $n_{0}$ and assuming $\varepsilon_{0}$ is sufficiently small when $n_{0}$ is sufficiently large. there exist constants $\gamma>0$ such that for any sufficiently small absolute error tolerance $\varepsilon>0$, we can choose an integer $D(n_{0})$ satisfying:
  1. (i).
    The graph $\mathcal{G}(D,\mathbf{b})$ has no more than $n_{0}$ nodes and no more then $m_{0}$ edges.
  2. (ii).
    If two matrices $\mathbf{b}$ and $\mathbf{b}^{\prime}$ have sums $S=\sum b_{ij}$ and $S^{\prime}=\sum b^{\prime}_{ij}$ such that $|S-S^{\prime}|\geq\gamma D$, then the corresponding PPR values satisfy $|\pi(s,t)-\pi^{\prime}(s,t)|>2\varepsilon$.
  3. (iii).
    $D^{2}=\Omega(\min(1/\varepsilon_{0}^{2},m_{0}))$.
Proof.
Choose $D=\sqrt{min(1/\varepsilon^{2},m_{0})}\frac{(1-\alpha)^{3}}{8}$. Condition $(i)$ follows directly from Lemma 27, since $n_{0}\geq\sqrt{m_{0}}/2\geq D$. Condition $(iii)$ follows by definition. Now our goal is to show condition $(ii)$. From the properties of our graph family, we have the change in PPR value:
$|\pi(s,t)-\pi^{\prime}(s,t)|=\alpha\frac{(1-\alpha)^{3}}{D^{2}}|S-S^{\prime}|$
Given that $|S-S^{\prime}|\geq\gamma D$ for some constant $\gamma$ we will choose, this implies:
$|\pi(s,t)-\pi^{\prime}(s,t)|\geq\alpha(1-\alpha)^{3}\gamma/D>2\varepsilon$
The next lemma formalizes the statistical difficulty of estimating the sum $S$ with the precision required for our lower bound.
Lemma 29.
For any constant $p\in(0,1)$, there exists a constant $\gamma^{\prime}>0$ such that for a sufficiently large $D$, any algorithm making at most $T\leq D^{2}/2$ queries to a graph drawn from $\Sigma(D)$ will fail to estimate $S=\sum_{i,j=1}^{D^{2}}b_{ij}$ with an error less than $\gamma^{\prime}D$. Specifically, if $\widehat{S}$ is the algorithm’s estimate, then:
$\operatorname{{\bf Pr}}_{\mathbf{b}\sim\Sigma(D)}[|S-\widehat{S}|>\gamma^{\prime}D]>p$
Proof.
Let $\mathcal{A}$ be an arbitrary algorithm operating within our query model. We first analyze the information gained from its queries. The graph structure is highly regular. The node set is fixed, as are the degrees of all nodes. Queries like DEG are uninformative about the matrix $\mathbf{b}$. Furthermore, the neighbors of $s$ (the set $Y$) and the nodes pointing to $t$ (the set $X$) are also fixed. Thus, ADJ queries on $s$ and $t$ yield no information about $\mathbf{b}$.
The only queries that reveal the structure of $\mathbf{b}$ are ADJ queries on nodes in $\{Y,Y^{\prime},X,X^{\prime}\}$. For instance, querying the $j$-th neighbor of $y_{i}$ reveals whether the edge is $(y_{i},x_{j})$ or $(y_{i},x_{j}^{\prime})$, which is equivalent to learning the value of $b_{ij}$. Each such query reveals exactly one entry of the matrix $\mathbf{b}$.
The algorithm $\mathcal{A}$ makes at most $T\leq D^{2}/2$ queries. Let $K$ be the set of $T$ indices $(i,j)$ corresponding to the entries of $\mathbf{b}$ that the algorithm learns. The algorithm knows the values $\{b_{k}\}_{k\in K}$. The remaining $M=D^{2}-T$ entries, where $k\notin K$, remain unknown. Since $T\leq D^{2}/2$, the number of unknown entries is $M\geq D^{2}/2$.
Let $S$ be the true sum of all entries in $\mathbf{b}$. We can write $S=S_{obs}+S_{unobs}$, where $S_{obs}=\sum_{k\in K}b_{k}$ is the sum of observed entries and $S_{unobs}=\sum_{k\notin K}b_{k}$ is the sum of unobserved entries. The algorithm knows $S_{obs}$ exactly.
The unobserved sum, $S_{unobs}$, is a sum of $M$ independent and identically distributed $Bern(1/2)$ random variables. Thus, $S_{unobs}$ follows a binomial distribution, $Binomial(M,1/2)$, with mean $\mu=M/2$ and standard deviation $\sigma=\sqrt{M}/2$.
An optimal algorithm’s best estimate for this unknown sum is its mean, $\widehat{S}_{unobs}=M/2$. The algorithm’s estimation error is therefore $|S_{unobs}-M/2|$. We want to show that $\operatorname{{\bf Pr}}[|S_{unobs}-M/2|>\gamma^{\prime}D]>p$.
Let’s standardize the variable: $Z_{M}=(S_{unobs}-\mu)/\sigma$. The failure probability becomes $\operatorname{{\bf Pr}}[|Z_{M}|>\gamma^{\prime}D/\sigma]$. Since $M\geq D^{2}/2$, the standard deviation $\sigma\geq D/(2\sqrt{2})$, which means the normalized error threshold is bounded by a constant: $\gamma^{\prime}D/\sigma\leq 2\sqrt{2}\gamma^{\prime}=\mathrel{\mathop{\ordinarycolon}}c_{\gamma}$. Thus, the failure probability is at least $\operatorname{{\bf Pr}}[|Z_{M}|>c_{\gamma}]$.
By the Berry-Esseen theorem, the cumulative distribution function $F_{M}(z)$ of $Z_{M}$ is pointwise close to the standard normal CDF $\Phi(z)$, with $|F_{M}(z)-\Phi(z)|\leq C/\sqrt{M}$ for a constant $C$. This allows us to lower-bound the failure probability:
$\operatorname{{\bf Pr}}[|Z_{M}|>c_{\gamma}]\geq 2(1-\Phi(c_{\gamma}))-\frac{2C}{\sqrt{M}}$
For any target probability $p\in(0,1)$, we can choose the constant $\gamma^{\prime}>0$ to be small enough such that $2(1-\Phi(c_{\gamma}))>p$ (since $\Phi(c_{\gamma})\to 0.5$ as $\gamma^{\prime}\to 0$). For a sufficiently large $D$, the error term $2C/\sqrt{M}$ becomes negligible, guaranteeing that the failure probability exceeds $p$. This completes the proof. ∎
Using the constructed $D$ in Lemma 28, we can conclude that using only $o(\min(1/\varepsilon_{0}^{2},m_{0}))$ queries will lead to a more than $p$ failure probability in evaluation of $\pi(s,t)$ with absolute error guarantee.∎
8 Conclusion
In this paper, we studied the computational complexity of Personalized PageRank (PPR) queries under both additive and relative error settings, focusing on three central problems: single-source PPR (SSPPR) with absolute error (SSPPR-A), single-source PPR with relative error (SSPPR-R), and single-target PPR (STPPR). Compared with the best-known lower bounds, we establish substantially tighter results:
  • For SSPPR-R, we prove an improved lower bound of $\Omega\!\left(\min\!\bigl(m,\tfrac{\log(1/\delta)}{\delta}\bigr)\right)$ for any $\delta\in(0,1)$, which matches the best known upper bounds when $\delta=\Theta(1/n)$—the most widely used regime in practice on power-law graphs.
  • For SSPPR-A, we obtain a lower bound of $\Omega\!\left(\min\!\bigl(m,\tfrac{\log(1/\varepsilon)}{\varepsilon^{2}}\bigr)\right)$ and further refine it to $\Omega\!\left(\min\!\bigl(m,\tfrac{1}{\varepsilon^{2}}\bigr)\right)$ under mild sparsity conditions $m\in O(n^{2-\beta})$ for any constant $\beta\in(0,1)$, partially matching our new $O(1/\varepsilon^{2})$ upper bound.
  • For STPPR, by adapting our hard-instance construction we prove a lower bound of $\Omega\!\left(\min\!\bigl(m,\tfrac{n}{\delta}\log(n)\bigr)\right)$, which also matches the best known upper bound when $\delta=\Theta(1/n)$.
These results substantially strengthen the theoretical foundation of PPR computation. They not only close the long-standing logarithmic gaps in SSPPR lower bounds but also achieve matching upper and lower bounds in key parameter regimes— notably $\delta=\Theta(1/n)$ for SSPPR-R and STPPR, and $\varepsilon=\Theta(n^{-1/2})$ for SSPPR-A. This demonstrates that the classical Monte Carlo (MC) method and its refined combinations such as FORA are, in fact, optimal in these practical settings.
More broadly, our new proof framework is flexible and of independent interest. Its ability to be reconfigured for different PPR query types (and potentially for other graph query problems) suggests new directions for future research in both designing optimal algorithms and deepening the complexity theory of graph similarity and ranking computations.
Appendix A Appendix
A.1 Upper Bounds Discussion
Since there lacks a unified summarization regarding the computational upper bounds for SSPPR-A and SSPPR-R queries, we first review the three basic PPR computation components: Monte Carlo (MC), Forward Push (FP), and Backward Push, and formulate the best-known upper bounds of later state-of-the-art (sota) methods.
A.1.1 Basic Components
MC.
The Monte Carlo method (MC) is a widely used technique for estimating desired values via sample-based computations. Leveraging the interpretation of Personalized PageRank (PPR) as $\alpha$-decay random walks, MC effectively approximates the SSPPR vector. Despite its apparent simplicity, MC provides unbiased and accurate PPR estimates with high probability, potentially surpassing theoretical lower bounds established for deterministic algorithms. The core idea of MC is to simulate $\alpha$-decay random walks starting from node $s$ and use the empirical distribution of their termination nodes as an approximation for $\pi(s,\cdot)$. This approach offers a straightforward and intuitive solution to the SSPPR query. Assuming $W$ independent random walks, the expected computational time is $\Theta(W/\alpha)$. By applying Chernoff bounds [8] and union bounds over all potential nodes, setting the number of random walks to $W=\Theta\left(\frac{\log(n/p_{f})}{c^{2}\delta}\right)$ suffices to answer the SSPPR-R query with relative error guarantees. Note that the $\log(n)$ term arises from the union bound over all $n$ nodes. Since only nodes with PPR value larger than $\delta$ are of interest, their number is at most $\frac{1}{\delta}$. This results in an expected computation time of $\operatorname{\mathcal{O}}\left(\frac{\log\left(1/\delta p_{f}\right)}{\alpha c^{2}\delta}\right)=\operatorname{\mathcal{O}}\left(\frac{\log\left(1/\delta\right)}{\delta}\right)$. Similarly, MC can be adapted to answer the SSPPR-A query with high probability in time $\operatorname{\mathcal{O}}\left(\frac{\log\left(1/\varepsilon\right)}{\varepsilon^{2}}\right)$.
Forward Push (FP).
FP [1] is a celebrated “local-push” algorithm for the SSPPR query, which takes as input a parameter $r_{\max}$ and runs in $O(1/r_{\max})$ time. As detailed in Algorithm 2, it simulates random walks deterministically by transport graph traversal probability and obtains an estimation $\widehat{\pi}(s,\cdot)$ which satisfies an overall approximation quality with $\widehat{\pi}(s,\cdot)-\pi(s,\cdot)||_{1}=r_{sum}=\sum_{u\in V}r_{s}(u)\leq m\cdot r_{max}$. To meet the error requirement of SSPPR-A, $r_{max}$ is required to set to $\varepsilon/m$ and results a pessimistic bound of $\operatorname{\mathcal{O}}(m/\varepsilon)$.
Backward Push (BP).
BP [21] is also a “local-push” algorithm, while for approximating $\pi(v,t)$ from all $v\in V$ to a given target node $t\in V$, known as the Single-Target PPR (STPPR) query. It works similarly as FP while from a inverse direction as described in Algorithm 3 and BP can also utilized to answer the SSPPR-A query by performing it with $r_{\max}=\varepsilon$ for each $t\in V$, resulting in a complexity of $O(m/\varepsilon)$, too. This bound is inferior but latter methods will modify it to achieve better results.
Input: graph $G$, source node $s$, decay factor $\alpha$ and error threshold $r_{max}$.
Output: estimation $\widehat{\pi}(s,\cdot)$ and resulted residue $r_{s}(\cdot)$.
$r_{s}(s)\leftarrow 1;r_{s}(u)\leftarrow 0$ for all $u\neq s$;
$\widehat{\pi}(s,u)\leftarrow 0$ for all $u\in V$;
while $\exists u\in V\ s.t.\ r_{s}(u)\geq r_{max}/d_{out}(u)$ do
 for each $v\in N_{out}(u)$ do
    $r_{s}(v)\leftarrow r_{s}(v)+(1-\alpha)\cdot r_{s}(u)/d_{out}(u)$;
   end for
 $\widehat{\pi}(s,u)\leftarrow\widehat{\pi}(s,u)+\alpha\cdot r_{s}(u);$
 $r_{s}(u)\leftarrow 0;$
end while
return $\widehat{\pi}(s,\cdot)$ and $r_{s}(\cdot)$
Algorithm 2 Forward Push [1] (FP)
Input: graph $G$, target node $t$, decay factor $\alpha$ and error threshold $r_{max}$.
Output: estimation $\widehat{\pi}(t,\cdot)$ and resulted residue $r_{t}(\cdot)$.
$r_{t}(t)\leftarrow 1;r_{t}(u)\leftarrow 0$ for all $u\neq t$;
$\widehat{\pi}(t,u)\leftarrow 0$ for all $u\in V$;
while $\exists u\in V\ s.t.\ r_{t}(u)\geq r_{max}$ do
 for each $v\in N_{in}(u)$ do
    $r_{t}(v)\leftarrow r_{t}(v)+(1-\alpha)\cdot r_{t}(u)/d_{out}(v)$;
   end for
 $\widehat{\pi}(t,u)\leftarrow\widehat{\pi}(t,u)+\alpha\cdot r_{t}(u);$
 $r_{t}(u)\leftarrow 0;$
end while
return $\widehat{\pi}(t,\cdot)$ and $r_{t}(\cdot)$
Algorithm 3 Backward Push [21] (BP)
A.1.2 SOTA Methods
FORA [48].
The FORA method is a representative algorithm for answering the SSPPR-R query with sub-linear time complexity. It adopts a two-phase framework by combining the FP and MC components. Generally speaking, it conducts random walks deterministically through the FP method before doing so randomly through MC, aiming to reduce the number of samples needed in MC. In detail, it first uses the FP method obtain an initial estimation. To further eliminating the error for each specific node, it performs MC on nodes with non-zero residues. Specifically, for each node $u$ with $r_{s}(u)>0$, it generates $W_{u}=\lceil r_{s}(u)\cdot W\rceil$ $\alpha$-decay random walks. Denoted $w(u,v)$ the number of such walks stops at $v$, it will add on the estimator $\widehat{\pi}(s,v)$ by $\frac{r_{s}(u)\cdot w(u,v)}{W_{v}}$ and the final estimation $\widehat{\pi}^{\prime}_{s}(v)$ is calculated as
$\widehat{\pi}^{\prime}_{s}(v)=\widehat{\pi}(s,v)+\sum_{u\in V\mathrel{\mathop{\ordinarycolon}}\ r_{s}(u)>0}\frac{r_{s}(u)\cdot w(u,v)}{W_{v}}.$
The FP part requires $\operatorname{\mathcal{O}}(\frac{1}{r_{max}})$ running costs and the MC phase needs to generate at most $\operatorname{\mathcal{O}}(m\cdot W\cdot r_{max})$ random walks. Thus the overall expected time complexity can be minimized by setting $r_{max}=\frac{1}{\sqrt{m\cdot W}}$, which results in the $\operatorname{\mathcal{O}}\left(\frac{1}{c}\sqrt{\frac{m\log\left(n/p_{f}\right)}{\delta}}\right)=\operatorname{\mathcal{O}}\left(\sqrt{\frac{m\log\left(n\right)}{\delta}}\right)$ bound.
SpeedPPR [38].
The SpeedPPR method is another sota method which can answer the SSPPR-A query efficiently. Similar to the concept of FORA, it also incorporates the phases of push and random walk. While it discovers that the push order in FP will greatly influence the decreasing rate of remaining residues. The arbitrary push operation ordering of FP makes the analysis of the error bound during the algorithm very challenging. Instead, it conducts the the push operations and makes the entire process more structured to bound the decrease rate of the residues’ summation $r_{sum}$ in an exponential manner feasibly. This results in a different complexity of $\operatorname{\mathcal{O}}(m\cdot\log(\frac{1}{m\cdot r_{max}})+m)$ in the push phase. By setting $r_{max}=m/\varepsilon$, a superior $\operatorname{\mathcal{O}}\left(m\log(\frac{1}{\varepsilon})\right)$ time bound is achieved for SSPPR-A query. Furthermore by combination with MC where for each node $u$ with non-zero residue $r_{s}(u)>0$, perform $\lceil r_{s}(u)\cdot W\rceil\leq\lceil d_{u}\cdot r_{max}\cdot W\rceil$ random walks, the overall complexity
$\operatorname{\mathcal{O}}\left(m\cdot\log\left(\frac{1}{m\cdot r_{max}}+m+m\cdot r_{max}\cdot W\right)\right)$
is minimized by setting $r_{max}=1/W$, resulting in $\operatorname{\mathcal{O}}\left(m\log(\frac{1}{\varepsilon})\right)$ time cost for SSPPR-R query.
ABP [47].
The ABP method is a recent sota method derived to answer the SSPPR-A query in sublinear time. Notably that it combines MC with an adaptive version of Backward Push by conducting push operation through a candidate node sets. This candidate node set is constructed by first running a $\operatorname{\mathcal{O}}(\log(n)/\varepsilon)$ random walks to obtain a rough estimation of $\pi_{s}^{\prime}(\cdot)$. Then, the adaptive backward push will iteratively halves it $r_{max}$ parameter until the cost exceeding a pre-defined threshold. Finally, it runs extra random walks to refine its estimation to meet the additive error requirement and achieves an overall expected time cost of $\operatorname{\mathcal{O}}\left(\frac{\sqrt{m\log(n)}}{\varepsilon}\right)$ for SSPPR-A query.
RBS [45].
RBS is a recent state-of-the-art method tailored to the STPPR query. It injects randomness into the Backward Push (BP) procedure to obtain an unbiased estimator with controlled variance. Specifically, each in-neighbor $v$ of a node $u$ is sampled with probability $\frac{(1-\alpha)r(u)}{\delta d_{\mathrm{out}}(v)}$, and $r_{t}(u)$ is increased by $\operatorname{\mathcal{O}}(\delta)$ only for the sampled $v\in\mathcal{N}_{\mathrm{in}}(u)$. This ensures that the expected increment of $r(v)$ remains $\operatorname{\mathcal{O}}\!\bigl(\frac{(1-\alpha)r(v)}{d_{\mathrm{out}}(u)}\bigr)$ for each $u\in\mathcal{N}_{\mathrm{in}}(v)$, while greatly reducing the number of actual updates to $r(u)$. RBS repeats this randomized push for every $v$ with nonzero $r(v)$ and achieves an expected time complexity of $\widetilde{\operatorname{\mathcal{O}}}\!\bigl(\sum_{v\in V}\frac{\pi(v,t)}{\delta}\bigr)$ for estimating $\pi(s,t)$ for all sources $s$. In the worst case, this yields a total running time of $\operatorname{\mathcal{O}}(n\log n/\delta)$.
A.2 Proof of Lemma 22
Lemma.
$\operatorname{{\bf Pr}}_{\mathrm{U}(r)\sim\Sigma(n,D,d)}[\mathrm{U}(r)\text{'s edge multiplicity is at least $L$}]\leq 4n^{2}(\frac{4D}{n})^{L}$, for $L\leq D$.
Proof.
Fix an arbitrary realization of $\left(\mathrm{SP},\,\mathrm{P}^{2D-d}_{1},\dots,\mathrm{P}^{2D-d}_{2n},\,\mathrm{P}^{n(2D-d)}_{Y_{1}},\,\mathrm{P}^{n(2D-d)}_{Y_{2}}\right)$, i.e., the split of $X$ and all in-/out-edge permutations chosen in Steps (i)–(iii) as described in Section 5.2. These choices determine the following graph information:
  • The set of labels for out-edges starting from nodes in $Y_{1}$ and pointing to nodes in $X_{1}$;
  • For every $x_{i}\in X_{1}$, the set of labels of its in-edges originating from $Y_{1}$.
The only remaining randomness that may affect parallel edge multiplicity arises from the uniform bijection $\mathrm{S}(Y_{1},X_{1})$. As described in Step (iv), this bijection maps the set of all out-edge labels from nodes in $Y_{1}$ directed toward $X_{1}$ to the set of all in-edge labels of nodes in $X_{1}$ originating from $Y_{1}$, thereby determining the actual edge connections. Fix a pair of nodes $x_{i}\in X_{1}$ and $y_{j}\in Y_{1}$ (an analogous argument applies to the cases $(X_{1},Y_{2})$, $(X_{2},Y_{1})$, and $(X_{2},Y_{2})$). Let $\mathrm{LABEL}_{\text{out}}(y_{j},X_{1})$ denote the set of at most 2Dd2D-d out-edge labels of $y_{j}$ that are directed toward $X_{1}$, and let $\mathrm{LABEL}_{\text{in}}(Y_{1},x_{i})$ denote the set of $D$ in-edge labels of $x_{i}$ originating from nodes in $Y_{1}$. We also define $\mathrm{LABEL}_{\text{out}}(Y_{1},X_{1})$ as the complete set of out-edge labels from all nodes in $Y_{1}$ targeting $X_{1}$, where each label includes both the source node $y_{j^{\prime}}^{(1)}$ and its local out-edge label. Since $\mathrm{S}(Y_{1},X_{1})$ is a uniformly random bijection over two sets of size $nD$, for any given $L$-subset of labels in $\mathrm{LABEL}_{\text{out}}(y_{j},X_{1})$ (i.e., any collection of $L$ out-edges from node $y_{j}$), its image under $\mathrm{S}^{-1}(Y_{1},X_{1})$ is distributed uniformly over all possible $L$-subsets of $\mathrm{LABEL}_{\text{out}}(Y_{1},X_{1})$. Therefore, for any particular $L$-subset $F\subset\mathrm{LABEL}_{in}(Y_{1},x_{i})$, we have that
$\operatorname{{\bf Pr}}\bigl[\mathrm{S}^{-1}(Y_{1},X_{1})(F)\subseteq\mathrm{LABEL}_{out}(y_{j},X_{1})\bigr]\;=\;\frac{\binom{D}{L}}{\binom{nD}{L}}$.
There are at most $\binom{2D-d}{L}$ ways to choose such an $F$. A union bound therefore gives
$\operatorname{{\bf Pr}}\bigl[|\{\text{edges between }x_{i},y_{j}\}|\geq L\,\bigm|\,\mathrm{P}^{2D-d}_{1},\dots,\mathrm{P}^{2D-d}_{2n},\,\mathrm{P}^{n(2D-d)}_{Y_{1}},\,\mathrm{P}^{n(2D-d)}_{Y_{2}}\bigr]\leq\binom{2D}{L}\cdot\frac{\binom{D}{L}}{\binom{nD}{L}}$.
When $n\geq 4$ (so that $nD\geq 4D\geq 2L$), we have the bounds $\binom{D}{L}\leq\frac{D^{L}}{L!}$, $\binom{2D}{L}\leq\frac{(2D)^{L}}{L!}$, and $\binom{nD}{L}\geq\frac{(nD/2)^{L}}{L!}$. These imply $\binom{2D}{L}\cdot\frac{\binom{D}{L}}{\binom{nD}{L}}\leq\left(\frac{2D}{n}\right)^{L}$. The same inequality holds for all other combinations of $(x_{i},y_{j})$, including $x_{i}\in X_{2}$ or $y_{j}\in Y_{2}$. Since there are only $2n$ nodes in $X$ and $2n$ nodes in $Y$, we conclude that
$\operatorname{{\bf Pr}}_{\mathrm{U}(r)\sim\Sigma(n,D,d)}\left[\text{edge multiplicity of }\mathrm{U}\text{ is at least }L\right]\leq 4n^{2}\left(\frac{4D}{n}\right)^{L},$
which completes the proof.
A.3 Proof of Lemma 23
Lemma (Reduction from simple to multi-graphs).
Choose any constant $L$, decay factor $\alpha\in(0,1)$, failure probability $p\in(0,1)$, error bound $c\in(0,\frac{1}{2}]$, and any function $\delta(n_{0})\in(0,1),m(n_{0})\in\Omega(n_{0})\cap\operatorname{\mathcal{O}}(n_{0}^{2})$. Suppose there exists an algorithm $\operatorname{\mathcal{A}}$ such that, for any simple directed graph with $\Theta(Ln_{0})$ nodes and $\Theta(Ln_{0}+Lm_{0})$ edges, $\operatorname{\mathcal{A}}$ answers the SSPPR-R problem with threshold $\delta$, error bound $c$, and expected query complexity $T$ in arc-centric model.
Then, there exists an algorithm $\mathcal{A}_{L}$ in arc-centric model such that for any integers $n\leq n_{0},\;D,\;d,\;r,\;n(D+r)\leq m$, and any $r$-padded instance of $\mathrm{U}\in\mathcal{U}(n,D,d)$ with edge multiplicity bounded by $L$:
  1. (i).
    Expected query complexity of $\mathcal{A}_{L}$ is bounded by a function of $n_{0}$ in $\operatorname{\mathcal{O}}(T)$ , and the failure probability of $\mathcal{A}_{L}$ is also bounded by $p$;
  2. (ii).
    $\mathcal{A}_{L}$ solves the SSPPR-R problem with decay factor $1-(1-\alpha)^{3}$, error requirement $c$, threshold $\delta$ for target nodes in set $X$ on $\mathrm{U}(r)$;
  3. (iii).
    $\mathcal{A}_{L}$ executes each query ${\textsf{ADJ}}_{in/out}(v,k)$ by a uniformly random selection of an uncovered label $k$ among $v$’s in or out edges.
An analogous conclusion holds for SSPPR-A case.
Our proof begins by considering how an algorithm $\mathcal{A}$ can operate on a multi-graph $\mathrm{U}(r)$ whose edge multiplicity is bounded by $L$. Intuitively, we construct an $L$-simple lift of $\mathrm{U}(r)$ by creating $L$ parallel copies of every node and edge, thereby transforming $\mathrm{U}(r)$ into a simple graph. This lifted graph allows $\mathcal{A}$ to run as if it were operating on a simple graph, while its queries can be simulated through queries on the original multi-graph $\mathrm{U}(r)$. However, this construction presents subtle challenges: since we are restricted to local queries on $\mathrm{U}(r)$, we might not even know the number of parallel edges between two already explored nodes. To address this, we introduce independent randomness at each local structure to probabilistically simulate a consistent global lift.
Lift of graph.
Figure 5: Illustration of a 22-Lift.
We now formalize this idea by describing the $L$-lift of a multi-graph, which encompasses the family of lifted simple graphs on which $\mathcal{A}$ will be executed, and which can be generated based on local information. Let $\mathrm{U}(r)=(V_{\mathrm{U}(r)},E_{\mathrm{U}(r)})$ be a multi-graph with $|V_{\mathrm{U}(r)}|\leq n$, $|E_{\mathrm{U}(r)}|\leq m$, and edge multiplicity $\mathrm{MUL}(V_{\mathrm{U}(r)})\leq L$, We then describe a $L$-lift of $\mathrm{U}(r),\mathcal{L}_{\rho}(\mathrm{U}(r))=(\widehat{V}_{U},\widehat{E}_{U})$ with new node set $\widehat{V}_{U}$, edge set $\widehat{E}_{U})$ and $\rho=(\rho_{(w,v)})_{(w,v)}$ is a sequence of attributes, which we will define later. Concerning $\widehat{V}_{U}$, we incorporate $2L$ extra nodes for every $v\in V_{\mathrm{U}(r)}$, : $v^{\mathrm{in}}_{1},\dots,v^{\mathrm{in}}_{L}$ , $v^{\mathrm{out}}_{1},\dots,v^{\mathrm{out}}_{L}$, which results in $\widehat{V}_{\mathrm{U}(r)}\;=\;V\;\cup\;\{v^{\mathrm{in}}_{1},\dots,v^{\mathrm{in}}_{L},v^{\mathrm{out}}_{1},\dots,v^{\mathrm{out}}_{L}\mid\forall v\in V_{\mathrm{U}(r)}\}.$ For $\widehat{E}_{U}$, We further construct it by two parts:
  • New edges. Add the edge $(v,\,v^{\mathrm{out}}_{i})$ (if any $v$’s out-degree is 0, we do not insert such edges) and $(v^{\mathrm{in}}_{i},\,v)$ (if any $v$’s in-degree is 0, we do not insert such edges) and $(v^{\mathrm{in}}_{i},\,v)$ for every $v\in V$ and $i\in\{1,\dots,L\}$.
  • Copy of original edges. For any node pair $(w,v)$, and choose an arbitrary injection $\rho_{(w,v)}\mathrel{\mathop{\ordinarycolon}}\{\text{edges from $w$ to $v$}\}\to\{1,\dots,L\}$, for each edge $e$ add below edges to $\widehat{E}_{U}$ for $i=1,\dots,L$:
    $\bigl(w^{\mathrm{out}}_{i},\;v^{\mathrm{in}}_{(i+\rho_{(w,v)}[e]-1)\bmod L}\bigr).$
Later, we use $\mathrm{INJ}_{L}(w,v)$ to denote all possible $\rho_{(v,w)}$, and define $\mathrm{INJ}_{L}(\mathrm{U}(r))\mathrel{\mathop{\ordinarycolon}}=\prod_{w,v\in V_{\mathrm{U}(r)}}\mathrm{INJ}_{L}(w,v)$ as the product over all node pairs in $\mathrm{U}(r)$. Since by construction there are at most $L$ edges from $w$ to $v$, the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$ is guaranteed to be a simple graph—i.e., one without parallel edges.
Construction of the algorithm.
Now that we have defined $L$-lift of a graph, then we can dive into the details of $\mathcal{A}_{L}$. Our goal is to construct an algorithm that queries on the original graph $\mathrm{U}(r)$ as if the original algorithm $\mathcal{A}$ were querying on a $L$-lift of $\mathrm{U}(r)$. Because $\mathcal{A}_{L}$ learns only from the original multi-graph $\mathrm{U}(r)$, it must maintain both a recording of all queries that it have done, and an auxiliary table about the lift parameter $\rho$. Specifically, $\mathcal{A}_{L}$ has an inner simulated algorithm $\mathcal{A}$ and internal state $\mathrm{U}_{0},\rho^{0}>$. We next describe the two components $\mathrm{U}_{0}$ and $\rho^{0}$.
  • $\mathrm{U}_{0}$ is the covered subgraph of $\mathrm{U}(r)$ by the queries it has executed on $\mathrm{U}(r)$; It consists of all nodes in $\mathrm{U}(r)$, all covered edges with original labels. For example, if the first query ${\textsf{Adj}}_{in}(x_{i},k)$ returns $(y_{j},k^{{}^{\prime}})$, we put an edge in $U_{0}$ between $x_{i},y_{j}$, and label it with the $k^{th}$ in-edge of $x_{i}$ as well as the $k^{\prime th}$ out edge of $y_{j}$, though there is only one edge to $x_{i}$ in $\mathrm{U}_{0}$.
  • $\rho^{0}=(\rho^{0}_{(w,v)})_{(w,v)}\in\mathrm{INJ}_{L}(\mathrm{U}_{0})$. Specifically for each pair of nodes $v,w$ in $\mathrm{U}(r)$, an injection that maps from $\{\text{covered edges from $w$ to $v$}\}$ to $\{1,\cdots,L\}$.
To simplify notation, we claim that an $\mathrm{U},\rho\in\mathrm{INJ}(\mathrm{U}(r))$ is over $\mathrm{U}_{0},\rho^{0}>$ if and only if $\mathcal{L}_{\rho_{0}}(\mathrm{U_{0}})\subset\mathcal{L}_{\rho}(\mathrm{U}(r))$; Equivalently, this means $\mathrm{U}_{0}\subset\mathrm{U}(r)$ up to their labeling , and $\rho_{(w,v)}(e)=\rho^{0}_{(w,v)}(e)$, for arbitrary node pairs $w,v$ in $\mathrm{U}(r)$ and $e$ from $w$ to $v$ in $\mathrm{U}_{0}$. Intuitively, this means $\mathcal{A}_{L}$ knows that it is possibly executed on $\mathrm{U}(r)$, and the inner $\mathcal{A}$ knows it is possible on $\mathcal{L}_{\rho}(\mathrm{U}(r))$. With this information stored, $\mathcal{A}_{L}$ is now able to simulate $\mathcal{A}$. Specifically, when $\mathcal{A}$ tries to execute a query $Q$, $\mathcal{A}_{L}$ processes it according to the strategies outlined below:
  1. (i).
    $\mathcal{A}_{L}$ needs not to execute any query, and can gives $\mathcal{A}$ with an inferred return, because the return value of $Q$ is identical on $\mathcal{L}_{\sigma}(\mathcal{U}(r))$ for all $\mathrm{U},\rho$ over $\mathrm{U}_{0},\rho^{0}>$. This means no matter how we choose $\mathrm{U}$ and how we modify $\rho^{0}$ later, the returning value of that query is the same, and then $\mathcal{A}_{L}$ need not do actual queries. The following queries fall into this case:
    • $Q={\textsf{INDEG}}$ or OUTDEG. This is because whatever $\mathrm{U}\in\mathrm{U}(r)$, the degree is fixed for each node, and in every $L$-lift of $\mathrm{U}(r)$, the copied nodes inherits those degrees as in definition;
    • $Q$ is an ${\textsf{ADJ}}_{in/out}$ query with an index $k$ exceeding the corresponding degree; its answer is likewise predetermined;
    • $Q={\textsf{ADJ}}_{in/out}(v,k)$. The answer is simply the node $v^{in/out}_{k}$ or ”None”;
    • $Q={\textsf{ADJ}}_{out}(w_{i}^{out},k)$ (the case $Q={\textsf{ADJ}}_{in}(w_{i}^{in},k)$ is analogous), where the label $k^{th}$ out-edge” of $w$ (or $k^{th}$ in-edge of $w$ ) corresponds to an edge $e$ in $\mathrm{U}_{0}$. Denote the other node of that edge $e$ is $v$ and $e$ is $v^{\prime}$s $k^{{}^{\prime}th}$ in-edge. Following the construction of simple lift, we have the target label $j=((i-1+\rho^{0}_{w,v}(e))\bmod L)+1$, then the inferred answer is $v^{in}_{j}$.
  2. (ii).
    $Q={\textsf{JUMP}}$(): just execute JUMP() on $\mathrm{U}(r)$ and return $w$. Then we choose an answer for $\mathcal{A}$ uniformly random in $\{w,w^{in}_{1},\cdots,w^{in}_{L},w^{out}_{1},w^{out}_{L}\}$.
  3. (iii).
    $Q={\textsf{ADJ}}_{out}(w_{i}^{out},k)$ or ${\textsf{ADJ}}_{in}(w_{i}^{in},k)$, where the label $k^{th}$ out-edge” of $w$ (or $k^{th}$ in-edge of $w$ ) has not appear in $\mathrm{U}_{0}$. In this case, we can execute query ${\textsf{ADJ}}_{out}(w,k)$ and then it turns to Case $(i)$ in the above.
Up to now, we finish the construction of $\mathcal{A}_{L}$ and then we will reveal next its properties.
Properties of the algorithm
Now it is time to verify each condition in the lemma for $\mathcal{A}_{L}$. Let us begin with the Condition $(iii)$ in the lemma. We first pre-process on $\mathcal{A}$ to make it satisfy this condition. Since $\mathcal{A}$ works on general graphs (without random or arbitrary labeling), and a uniform bound for complexity or failure probability, we can assume that whenever $\mathcal{A}$ tries to execute a ADJ query at a node $v$, it selects an uncovered in (or out) edge of $v$ uniformly at random. Switching to these random selection strategy will make $\mathcal{A}$’s complexity and failure probability to be average on each labeling case, for all possible edge labeling, while earlier works [7; 46], it’s by default that the algorithm executes this random strategy on unlabeled graphs.
Now we address Conditions $(i)$ and $(iii)$. Before proceeding, we compute the PPR values on the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$. We demonstrate that for any $\mathrm{U}(r)$, where $\mathrm{U}\in\mathcal{U}(n,D,d)$, the PPR values of the target nodes in $X$ remain unchanged under the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$. As shown in Lemma 17, when the decay factor is $1-(1-\alpha)^{3}$, the PPR value satisfies $\pi_{U(r)}(s,x_{i})=\frac{D}{n(2D-d+r)}(1-\alpha)^{6}$. For the lifted graph $\mathcal{L}_{\rho}(\mathrm{U}(r))$, we compute the PPR values step by step. For any $y_{j}\in Y_{1}$, we have $\pi_{\mathcal{L}_{\rho}(U(r))}(s,y_{j})=\frac{1}{n}\alpha(1-\alpha)$. Since each $x_{i}\in X_{1}$ has no out-edges, we have
$\pi_{\mathcal{L}_{\rho}(\mathrm{U}(r))}(s,x_{i})=\frac{\alpha(1-\alpha)^{6}D}{n(2D-d+r)}\cdot\sum_{t=0}^{\infty}(1-\alpha)^{t}=\frac{(1-\alpha)^{6}D}{n(2D-d+r)}=\pi_{\mathrm{U}(r)}(s,x_{i}).$
An analogous argument holds for any $x_{i^{\prime}}\in X_{2}$, for which $\pi_{\mathcal{L}_{\rho}(U(r))}(s,x_{i^{\prime}})=\frac{D-d}{n(2D-d+r)}(1-\alpha)^{6}=\pi_{U(r)}(s,x_{i^{\prime}})$. Thus, Condition $(ii)$ is satisfied when running $\mathcal{A}$ on $\mathcal{L}_{\rho}(\mathrm{U}(r))$. In conclusion, for any choice of $\rho$, the PPR values on the node set $X$ remain identical between $\mathrm{U}(r)$ and its lift $\mathcal{L}_{\rho}(\mathrm{U}(r))$. This validates the construction of $\mathcal{A}_{L}$, since it simulates the behavior of $\mathcal{A}$ on some lifted version $\mathcal{L}_{\rho}(\mathrm{U}(r))$. By definition, $\mathcal{A}$ is capable of solving the SSPPR problem on any lifted graph $\mathcal{L}_{\sigma}(\mathrm{U}(r))$, for arbitrary $\sigma$, with failure probability at most $p$ and query complexity $\mathcal{O}(T)$. Therefore, if we sample $\rho\sim\mathrm{Uniform}(\mathrm{INJ}(\mathrm{U}(r)))$, and execute $\mathcal{A}$ on $\mathcal{L}_{\rho}(\mathrm{U}(r))$, we obtain expected query complexity $\mathcal{O}(T)$ and failure probability bounded by $p$, successfully solving the SSPPR problem on set $X$.
Then for Condition $(ii)$ and $(iii)$, we need to show that the query-and-answer sequence which $\mathcal{A}_{L}$ receives and sends to its inner simulated $\mathcal{A}$, has the same distribution as the distribution of query-and-answer sequence which $\mathcal{A}$ generates when executed on $\mathcal{L}_{\rho}(\mathrm{U})$ for a uniformly random sample $\rho\sim Uniform(\mathrm{INJ}_{L}(\mathrm{U}(r)))$. That is, for all $Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}$
$|\mathrm{INJ}(\mathrm{U}(r))|^{-1}\sum_{\rho}\operatorname{{\bf Pr}}[\mathcal{A}\text{ generates }Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}\text{ on }\mathcal{L}_{\rho}(\mathrm{U}(r))]$
$=\operatorname{{\bf Pr}}[\mathcal{A}_{L}\text{ generates }Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}].$
Before we use induction on $T$ to prove this result, we will first evaluate each term of
$\operatorname{{\bf Pr}}[\mathcal{A}\text{ generates }Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}\text{ on }\mathcal{L}_{\rho}(\mathrm{U}(r))],$
and then showing that it is almost a constant by another induction.
Now we begin this induction with the recursive formula, by showing it is almost a constant. Since $\mathcal{A}$ can only learn information through the sequence $Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}$, the distribution of next query: $P(Q_{t+1}|Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t})$ is independent of $\rho$. Also notice that the distribution of returning value is independent of $\mathcal{A}$, then we have
[…]
$=\operatorname{{\bf Pr}}_{\mathcal{A}}[Q_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}]\operatorname{{\bf Pr}}_{\rho}[Q_{t+1}=R_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}].$
In addition, $\operatorname{{\bf Pr}}_{\rho}[Q_{t+1}=R_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}]$ is either $1$ or 0 (If the query has a deterministic answer on each $\mathcal{L}_{\rho}(\mathrm{U}(r))$), or a constant not related to $\rho$ (JUMP), then the induction is completed, and we have a deterministic function $f$ such that:
$\operatorname{{\bf Pr}}_{\rho}[Q_{t+1}=R_{t+1}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}]$
$=\begin{cases}f(Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}),\text{ event }Q_{1}=R_{1},\cdots,Q_{t}=R_{t}\text{ happens on }{L}_{\rho}(\mathrm{U})\\ 0\end{cases}$
Then we can go back to our first induction that we need to proof
[…]
$=\operatorname{{\bf Pr}}[Q_{t+1}=R_{t+1}\text{ for }\mathcal{A}_{L}\mid Q_{1},\cdots,Q_{t};R_{1},\cdots,R_{t}].$
When $\mathcal{A}_{L}$ can infer the result of $Q_{t+1}$(Case (i)(i)which means for all possible $\rho,Q_{1}=R_{1},\cdots,Q_{t+1}=R_{t+1}\text{ happens on }\mathcal{L}_{\rho}(\mathrm{U}(r))$ they share the same answer), both sides of the equation are either $1$ or 0, and the JUMP case is similar. So the only non-trivial case is Case (iii) in the construction of $\mathcal{A}_{L}$. Without loss of generality, we assume $Q_{t+1}={\textsf{ADJ}}_{in}(v^{in}_{i},k),R_{t+1}={\textsf{ADJ}}_{out}(w^{out}_{j},k^{{}^{\prime}})$. In these cases we randomly choose a new edge’s embedding, therefore both side of the equation are
[…]
A.4 Proof of Lemma 25
Lemma.
There is a constant $\gamma$, such that. For $n$ sufficiently large, $d\leq log(n)$, $D>\frac{3d}{2}$, $T\leq\gamma nD$. Given any query and returning sequence $Q_{1},\cdots,Q_{T};R_{1},\cdots,R_{T}$, and a split of $X$, $\mathrm{SP}$ we have:
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}[\mathrm{SP}\text{ is the correct split of }\mathrm{U}\mid Q_{1},..,Q_{T};R_{1},\cdots,R_{T}]\leq\frac{1}{n}.$
Proof.
Our goal is to bound
$\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)}\!\Bigl[\,\mathrm{SP}\text{ is the true split}\;\Bigm|\;Q_{1}=R_{1},\dots,Q_{T}=R_{T}\Bigr].$
The distribution $\Sigma(n,D,d)$ is generated in two independent steps:
  1. (i).
    Choose the split.  Pick $\mathrm{SP}\in\binom{X}{n}$ uniformly at random.
  2. (ii).
    Choose all permutations. We choose all $\bigl(\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\,\mathrm{P}_{Y_{1}}^{n(2D-d)},\,\mathrm{P}_{Y_{2}}^{n(2D-d)},\,\mathrm{S}(Y_{a},X_{b})\bigr)$, where $\mathrm{S}(Y_{a},X_{b})\in\{\mathrm{S}(Y_{1},X_{1}),\mathrm{S}(Y_{1},X_{2}),\mathrm{S}(Y_{2},X_{1}),\mathrm{S}(Y_{2},X_{2})\}$, randomly and independently.
By Bayes’ rule, the posterior probability mass on the event: $\mathrm{SP}\text{ is the true split}\,$, equals to
$\frac{\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]}{\displaystyle\sum_{\mathrm{SP}^{{}^{\prime}}\in\binom{X}{n}}\operatorname{{\bf Pr}}_{\mathrm{U}\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]},$ (2)
where the sum runs over any $\sigma=(\mathrm{SP},\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)},\mathrm{S}(Y_{a},X_{b})),\forall a,b\in\{1,2\}$. Because all splits are a-prior equiprobable, the posterior on any particular $\mathrm{SP}$ is proportional to the likelihood of the observed transcript. Now we will give a systematic evaluation of each term in this fraction. Since queries from $s$ or to $s$ is the same on all the $\mathrm{U}(r)\in\mathcal{U}(n,D,d)$, we assume all the queries $Q_{t},R_{t}$ is only between $X$ and $Y$. We first utilize a toy case for illustration of our proof.
Warm-up.
Denote an event as $Q=R\mathrel{\mathop{\ordinarycolon}}Q\;=\;{\textsf{ADJ}}_{in}(x_{i},k_{Q}),R\;=\;(y_{j},k_{R}),$ i.e. query the $k_{Q}^{th}$ in-edge of $x_{i}$ and find it is the $k_{R}^{th}$ out-edge of $y_{j}$. Assume (just for exposition) that the split $\mathrm{SP}$ places $x_{i}\in X_{1}$ and $y_{j}\in Y_{1}$. Four independent choices determine whether this exact edge appears:
  1. (i).
    The split $\mathrm{SP}$ determines whether $x_{i}$ is in $X_{1}$ or $X_{2}$. In our assumption, it is in $X_{1}$.
  2. (ii).
    The permutation $\mathrm{P}^{2D-d}_{i}$ decides whether the $k_{Q}^{th}$ edge of $x_{i}$ is sourced from $Y_{1}$. This occurs with probability $\frac{D}{2D-d}$.
  3. (iii).
    The permutation $\mathrm{P}_{Y_{1}}^{n(2D-d)}$ determines whether the $k_{R}^{th}$ edge of $x_{i}$ points to $X_{1}$. This occurs with probability $\frac{nD}{n(2D-d)}$.
  4. (iv).
    The permutation $\mathrm{S}(Y_{1},X_{1})$ of $nD$ decides the particular out-edge index from the reverse of the $k_{Q}^{th}$ in-edge of $x_{i}$ among $nD$ choices. The probability of finding it is exactly the $k_{R}^{th}$ out-edge of $y_{j}$ is $\frac{1}{nD}$.
  5. (v).
    Other components, such as $\mathrm{P}_{Y_{2}}^{n(2D-d)}$ and $\mathrm{P}_{i^{\prime}}^{2D-d}$ for $i^{\prime}\neq i$, are irrelevant to the query–response pair $(Q,R)$ due to the independence across different parts of the construction.
Therefore, we have $\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)\,|_{SP}}[Q=R]=\frac{D}{D-d}\times\frac{nD}{n(2D-d)}\times\frac{1}{nD}$. Bearing these steps in mind, we now derive Lemma 30 to compute the overall probability. For ease of representation, we include the following additional notations:
  • Query-returning pairs are represented by: $Q_{t}={\textsf{ADJ}}_{in}(x_{i_{t}},k_{Q_{t}}),R_{t}=(y_{j_{t}},k_{R_{t}})$, or $Q_{t}={\textsf{ADJ}}_{out}(y_{j_{t}},k_{Q_{t}}),R_{t}=(x_{i_{t}},k_{R_{t}}),1\leq t\leq T_{0}$.
  • We use $e_{t}$ to represent the edge in $\mathrm{U}(r)$ with the query-returning pair $Q_{t},R_{t}$ goes through. It is an edge from $y_{j_{t}}$ to $x_{i_{t}}$. For node sets (or single node labels) $V_{1},V_{2}$, we use $E_{t}(V_{1}\rightarrow V_{2})$ for the count of $\{1\leq t^{\prime}<t,e_{t^{\prime}}\text{ originates from }V_{1},\text{pointing to }V_{2}\}$, e.g., $E_{t}(Y_{1},x_{i_{t}})$ is the count of edges from $Y_{1}$ to $x_{i_{t}}$ which on of the query-returning pairs $(Q_{1},R_{1}),\cdots,(Q_{t-1},R_{t-1})$ has passed through.
We now formulate the calculation in the following lemma.
Lemma 30.
(Conditional probability of queries) For any $x_{i_{T_{0}+1}}\in X_{b},\;y_{j_{T_{0}+1}}\in Y_{a},\;a,b\in\{1,2\},\mathbf{1}_{a=b}=\begin{cases}1,a=b\\ 0,a\neq b\end{cases}$,
$\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\bigl[Q_{T_{0}+1}=R_{T_{0}+1}\mid Q_{1}=R_{1},\dots,Q_{T_{0}}=R_{T_{0}}\bigr]$
$=\frac{D-(1-\mathbf{1}_{a=b})d-E_{T_{0}+1}(Y_{a}\rightarrow{x_{T_{0}+1}})}{\left(2D-d-E_{T_{0}+1}({Y\rightarrow{x_{T_{0}+1}}})\right)\left(n(2D-d)-E_{T_{0}+1}({Y_{a}\rightarrow X}\right)}).$
We proceed by induction; our induction hypothesis is: In the distribution $\Sigma(n,D,d)|_{SP}$ conditioned on $Q_{1}=R_{1},\dots Q_{t}=R_{t}$, $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ are still independent of each other, and that $\mathrm{S}(Y_{1},X_{1}),\mathrm{S}(Y_{1},X_{2}),\mathrm{S}(Y_{2},X_{1}),\mathrm{S}(Y_{2},X_{2})$ are also independent of each other. Without loss of generality, we prove the case $x_{i_{T_{0}+1}}\in X_{1},y_{j_{T_{0}+1}}\in Y_{1}$. As in the single query:
  1. (i).
    By definition, $\mathrm{SP}$ determines whether $x_{i_{T0+1}}$ is in $X_{1}$ or $X_{2}$. Here we assume that it lies in $X_{1}$.
  2. (ii).
    $\mathrm{P}^{2D-d}_{i_{T_{0}+1}}$ decides whether the $k_{Q_{T_{0}+1}}^{th}$ edge of $x_{i_{T_{0}+1}}$ is sourced from $Y_{1}$. This occurs with probability $\frac{D-E_{T_{0}+1}({Y_{1}\rightarrow{x_{i_{T_{0}+1}}}})}{2D-d-E_{T_{0}+1}({Y\rightarrow{x_{i_{T_{0}+1}}}})}$.
  3. (iii).
    $\mathrm{P}_{Y_{1}}^{n(2D-d)}$ decides whether the $k_{R_{T_{0}+1}}^{th}$ edge of $y_{j_{T_{0}+1}}$ targets to $X_{1}$. This occurs with probability $\frac{nD-E_{T_{0}+1}{(Y_{1}\rightarrow X_{1}})}{n(2D-d)-E_{T_{0}+1}({Y_{1}\rightarrow X})}$.
  4. (iv).
    Since in the composition, the $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ terms are irrelevant to each other by induction hypothesis: For simplification of notations, we call the event for $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ suitable for $R_{t},Q_{t}$: When $x_{i_{t}}\in X_{1},y_{j_{t}}\in Y_{2}$, $Q_{t}={\textsf{ADJ}}_{in}(x_{i_{t}},k_{Q_{t}}),R_{t}=(y_{j_{t}},k_{R_{t}})$, the event holds if and only if the $k_{Q_{t}}$ in-edge of $x_{i_{t}}$, i.e. $(x_{i_{t}},k_{Q_{t}})$ originates from nodes in $Y_{2}$; and the out-edge $(y_{j_{t}},k_{R_{t}})$ points to nodes in $X_{1}$. We analogously define the same event for the other seven cases $x_{i_{t}}\in X_{2}$ or $y_{j_{t}}\in Y_{1}$ or $Q_{t}={\textsf{ADJ}}_{out}(y_{j_{t}},k_{Q_{t}}),R_{t}=(x_{i_{t}},k_{R_{t}})$. Combined with induction hypothesis, the probability of
    $\operatorname{{\bf Pr}}_{\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}}\bigl[\text{suitable for }\mathrm{R}_{T_{0}+1},Q_{T_{0}+1}\mid\text{suitable for }R_{t},Q_{t},t=1,\cdots,T_{0}\bigr]$
    $=\frac{D-E_{T_{0}+1}({Y_{1}\rightarrow{x_{i_{T_{0}+1}}}})}{2D-d-E_{T_{0}+1}({Y\rightarrow{x_{i_{T_{0}+1}}}})}\times\frac{nD-E_{T_{0}+1}({Y_{1}\rightarrow X_{1}})}{n(2D-d)-E_{T_{0}+1}({Y_{1}\rightarrow X})}.$
  5. (v).
    Conditioned on arbitrary possible $\mathrm{P}^{n(2D-d)}_{i},\mathrm{P}^{n(2D-d)}_{Y_{1}},\mathrm{P}^{n(2D-d)}_{Y_{2}}$, the bijection $\mathrm{S}(Y_{1},X_{1})$ of $nD$ elements decides the particular out-edge index from the reverse of the $k_{Q_{T_{0}+1}}^{th}$ in-edge of $x_{i_{T_{0}+1}}$ among $nD-E_{Y_{1}\rightarrow X_{1}}$ choices, The probability of exactly finding it is the $k_{R_{T_{0}+1}}^{th}$ out-edge of $y_{j_{T_{0}+1}}$ is $\frac{1}{nD-E_{T_{0}+1}({Y_{1}\rightarrow X_{1}})}$.
Consequently, the resulting probability matches the expression in Lemma 30. Since we only set some restriction independently for $\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}^{2D-d}_{i_{T_{0}+1}}$, the $\mathrm{P}_{1}^{2D-d},\cdots,\mathrm{P}_{2n}^{2D-d},\mathrm{P}_{Y_{1}}^{n(2D-d)},\mathrm{P}_{Y_{2}}^{n(2D-d)}$ are still independent conditioned on $Q_{t+1}=R_{t+1}$, and also for the independence of
$\mathrm{S}(Y_{1},X_{1}),\;\mathrm{S}(Y_{1},\;X_{2}),\;\mathrm{S}(Y_{2},X_{1}),\;\mathrm{S}(Y_{2},X_{2}).$
Thus the induction hypothesis remains valid for $T_{0}+1$.
Combining the analysis above, it remains to show the left hand side expression is smaller than $\frac{1}{n}$ for Lemma 25. When $T\leq\gamma nD$, there are at most $\frac{\gamma}{\gamma_{0}}n$ nodes $i$ such that $E_{T}({Y\rightarrow x_{i}})\geq\gamma_{0}D$. Then under the split $\mathrm{SP}$, there are at least $(1-\frac{\gamma}{\gamma_{0}})\cdot n$ nodes in $X_{1}$, and at least $(1-\frac{\gamma}{\gamma_{0}})\cdot n$ nodes in $X_{2}$. For each pair of $x_{i},x_{i^{{}^{\prime}}}$ such that $E_{T}({Y\rightarrow x_{i}}),E_{T}({Y\rightarrow x_{i^{{}^{\prime}}})}<\gamma_{0}D$ and in $\mathrm{SP}$ $x_{i}\in X_{1},x_{i^{{}^{\prime}}}\in X_{2}$, we have another split $\mathrm{SP}^{{}^{\prime}}$, that $\mathrm{SP}^{{}^{\prime}}$ sends $x_{i}$ to $X_{2}$, $x_{i^{{}^{\prime}}}$ to $X_{1}$, while other nodes are the same as $\mathrm{SP}$; Further, there are at least $(1-\frac{\gamma}{\gamma_{0}})^{2}n^{2}$ such pairs. Intuitively, $\mathrm{SP}^{{}^{\prime}}$ is very close to $\mathrm{SP}$, and the success probability should be very close to that of $\mathrm{SP}$ as well if we have only known insufficient data. Specifically, we calculate the proportion of their success probability as follows:
$\frac{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]}{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP^{{}^{\prime}}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]}$
$=\Pi_{t=0,\cdots,T-1}\frac{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP}}\!\bigl[Q_{t+1}=R_{t+1}\mid Q_{1}=R_{1},\dots,Q_{t}=R_{t}\bigr]}{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{SP^{{}^{\prime}}}}\!\bigl[Q_{t+1}=R_{t+1}\mid Q_{1}=R_{1},\dots,Q_{t}=R_{t}\bigr]}.$
Denote $Q_{t},R_{t}$ by either ${\textsf{ADJ}}_{in}(x_{i_{t}},k_{Q_{t}}),(y_{j_{t}},k_{R_{t}})$ or ${\textsf{ADJ}}_{out}(y_{j_{t}},k_{Q_{t}}),(x_{i_{t}},k_{R_{t}})\ \text{for }t=1,..,T$. By Lemma 30, we have that for any $x_{i_{t+1}}\in X_{b},y_{j_{t+1}}\in Y_{a},a,b\in\{1,2\}$,
$\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\!\bigl[Q_{t+1}=R_{t+1}\mid Q_{1}=R_{1},\dots,Q_{t}=R_{t}\bigr]$
$=\frac{D-(1-\mathbf{1}_{a=b})d-E_{t+1}(Y_{a}\rightarrow{x_{i_{t+1}}})}{\left(2D-d-E_{t+1}({Y\rightarrow{x_{i_{t+1}}}})\right)\left(n(2D-d)-E_{t+1}({Y_{a}\rightarrow X}\right)}.$
[…]
$1+\frac{d}{2D-d-E_{Y\rightarrow x_{i_{t+1}}}}\leq 1+\frac{d}{2D-d-\gamma_{0}D}\leq 1+\frac{2}{4-3\gamma_{0}}\frac{d}{D}.$
Thus, it holds that
$\frac{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]}{\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]}\leq\left(1+\frac{2}{4-3\gamma_{0}}\frac{d}{D}\right)^{\gamma_{0}D}.$
Since $d\leq\log(n)$, this expression is at most $n^{\frac{2\gamma_{0}}{3-4\gamma_{0}}}$. Notice that there are at least $(n-\frac{\gamma}{\gamma_{0}}n)^{2}$ such pairs, which creates $(n-\frac{\gamma}{\gamma_{0}}n)^{2}$ terms of $\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]$ in the lower part of Equation 2, and they are at least $n^{-\frac{2\gamma_{0}}{3-4\gamma_{0}}}\cdot\operatorname{{\bf Pr}}_{U\sim\Sigma(n,D,d)|_{\mathrm{SP}^{{}^{\prime}}}}\!\bigl[Q_{1}=R_{1},\dots,Q_{T}=R_{T}\bigr]$. Thereby the expression is no more than $(1-\frac{\gamma}{\gamma_{0}})n^{-2+\frac{2\gamma_{0}}{3-4\gamma_{0}}}$. When $\gamma$ is a small constant and $n$ is sufficiently large, the probability is smaller than $\frac{1}{n}$. ∎